gnu Core-utils

A set of command line utilities for version 4.5.4, 2 December 2002

David MacKenzie et al.

This manual documents version 4.5.4 of the gnu command line utilities. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994, 1995, 1996, 2000, 2001, 2002 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

1 Introduction

This manual is a work in progress: many sections make no attempt to explain basic concepts in a way suitable for novices. Thus, if you are interested, please get involved in improving this manual. The entire gnu community will benefit.

The gnu utilities documented here are mostly compatible with the posix standard. Please report bugs to bug-coreutils@gnu.org. Remember to include the version number, machine architecture, input files, and any other information needed to reproduce the bug: your input, what you expected, what you got, and why it is wrong. Diffs are welcome, but please include a description of the problem as well, since this is sometimes difficult to infer. See section “Bugs” in Using and Porting GNU CC.

This manual was originally derived from the Unix man pages in the distributions, which were written by David MacKenzie and updated by Jim Meyering. What you are reading now is the authoritative documentation for these utilities; the man pages are no longer being maintained. The original fmt man page was written by Ross Paterson. François Pinard did the initial conversion to Texinfo format. Karl Berry did the indexing, some reorganization, and editing of the results. Brian Youmans of the Free Software Foundation office staff combined the manuals for textutils, fileutils, and sh-utils to produce the present omnibus manual. Richard Stallman contributed his usual invaluable insights to the overall process.

2 Common options

Certain options are available in all of these programs. Rather than writing identical descriptions for each of the programs, they are described here. (In fact, every gnu program accepts (or should accept) these options.)

Normally options and operands can appear in any order, and programs act as if all the options appear before any operands. For example, ‘sort -r passwd -t :’ acts like ‘sort -r -t : passwd’, since ‘:’ is an option-argument of ‘-t’. However, if the POSIXLY_CORRECT environment variable is set, options must appear before operands, unless otherwise specified for a particular command.

Some of these programs recognize the ‘--help’ and ‘--version’ options only when one of them is the sole command line argument.

‘--help’ Print a usage message listing all available options, then exit successfully.

‘--version’

Print the version number, then exit successfully.

‘--’ Delimit the option list. Later arguments, if any, are treated as operands even if they begin with ‘-’. For example, ‘sort -- -r’ reads from the file named ‘-r’.

A single ‘-’ is not really an option, though it looks like one. It stands for standard input, or for standard output if that is clear from the context, and it can be used either as an operand or as an option-argument. For example, ‘sort -o - -’ outputs to standard output and reads from standard input, and is equivalent to plain ‘sort’. Unless otherwise specified, ‘-’ can appear in any context that requires a file name.

2.1 Backup options

Some gnu programs (at least cp, install, ln, and mv) optionally make backups of files before writing new versions. These options control the details of these backups. The options are also briefly mentioned in the descriptions of the particular programs.

‘-b’
‘--backup[=method]’

Make a backup of each file that would otherwise be overwritten or removed. Without this option, the original versions are destroyed. Use method to determine the type of backups to make. When this option is used but method is not specified, then the value of the VERSION_CONTROL environment variable is used. And if VERSION_CONTROL is not set, the default backup type is ‘existing’. Note that the short form of this option, ‘-b’ does not accept any argument. Using ‘-b’ is equivalent to using ‘--backup=existing’. This option corresponds to the Emacs variable ‘version-control’; the values for method are the same as those used in Emacs. This option also accepts more descriptive names. The valid methods are (unique abbreviations are accepted):

          ‘none’

‘off’ Never make backups.

          ‘numbered’

‘t’ Always make numbered backups.

          ‘existing’

‘nil’ Make numbered backups of files that already have them, simple backups of the others.

          ‘simple’

‘never’ Always make simple backups. Please note ‘never’ is not to be confused with ‘none’.

‘-S suffix’
‘--suffix=suffix’

Append suffix to each backup file made with ‘-b’. If this option is not specified, the value of the SIMPLE_BACKUP_SUFFIX environment variable is used. And if SIMPLE_BACKUP_SUFFIX is not set, the default is ‘~’, just as in Emacs.

‘--version-control=method’

This option is obsolete and will be removed in a future release. It has been replaced with --backup.

2.2 Block size

Some gnu programs (at least df, du, and ls) display sizes in “blocks”. You can adjust the block size and method of display to make sizes easier to read. The block size used for display is independent of any filesystem block size. Fractional block counts are rounded up to the nearest integer.

The default block size is chosen by examining the following environment variables in turn; the first one that is set determines the block size.

DF_BLOCK_SIZE

This specifies the default block size for the df command. Similarly, DU_BLOCK_ SIZE specifies the default for du and LS_BLOCK_SIZE for ls.

BLOCK_SIZE

This specifies the default block size for all three commands, if the above command-specific environment variables are not set.

POSIXLY_CORRECT

If neither the command_BLOCK_SIZE nor the BLOCK_SIZE variables are set, but this variable is set, the block size defaults to 512.

If none of the above environment variables are set, the block size currently defaults to 1024 bytes in most contexts, but this number may change in the future. For ls file sizes, the block size defaults to 1 byte.

A block size specification can be a positive integer specifying the number of bytes per block, or it can be human-readable or si to select a human-readable format. Integers may be followed by suffixes that are upward compatible

with the SI prefixes (http://www.bipm.fr/enus/3_SI/si-prefixes.html)

for decimal multiples and with the IEC 60027-2 prefixes for binary multiples

(http://physics.nist.gov/cuu/Units/binary.html).

With human-readable formats, output sizes are followed by a size letter such as ‘M’ for megabytes. BLOCK_SIZE=human-readable uses powers of 1024; ‘M’ stands for 1,048,576 bytes. BLOCK_SIZE=si is similar, but uses powers of 1000 and appends ‘B’; ‘MB’ stands for 1,000,000 bytes.

A block size specification preceded by ‘'’ causes output sizes to be displayed with thousands separators. The LC_NUMERIC locale specifies the thousands separator and grouping. For example, in an American English locale, ‘--block-size="'1kB"’ would cause a size of 1234000 bytes to be displayed as ‘1,234’. In the default C locale, there is no thousands separator so a leading ‘'’ has no effect.

An integer block size can be followed by a suffix to specify a multiple of that size. A bare size letter, or one followed by ‘iB’, specifies a multiple using powers of 1024. A size letter followed by ‘B’ specifies powers of 1000 instead. For example, ‘1M’ and ‘1MiB’ are equivalent to ‘1048576’, whereas ‘1MB’ is equivalent to ‘1000000’.

A plain suffix without a preceding integer acts as if ‘1’ were prepended, except that it causes a size indication to be appended to the output. For example, ‘--block-size="kB"’ displays 3000 as ‘3kB’.

The following suffixes are defined. Large sizes like 1Y may be rejected by your computer due to limitations of its arithmetic.

‘kB’ kilobyte: 10^{3} = 1000.

‘k’ ‘K’ ‘KiB’ kibibyte: 2^{1}0 = 1024. ‘K’ is special: the SI prefix is ‘k’ and the IEC 60027-2

prefix is ‘Ki’, but tradition and posix use ‘k’ to mean ‘KiB’.

‘MB’ megabyte: 10^{6} = 1, 000, 000.

‘M’ ‘MiB’ mebibyte: 2^{2}0 = 1, 048, 576.

‘GB’ gigabyte: 10^{9} = 1, 000, 000, 000.

‘G’ ‘GiB’ gibibyte: 2^{3}0 = 1, 073, 741, 824.

‘TB’ terabyte: 10^{1}2 = 1, 000, 000, 000, 000.

‘T’ ‘TiB’ tebibyte: 2^{4}0 = 1, 099, 511, 627, 776.

‘PB’ petabyte: 10^{1}5 = 1, 000, 000, 000, 000, 000.

‘P’ ‘PiB’ pebibyte: 2^{5}0 = 1, 125, 899, 906, 842, 624.

‘EB’ exabyte: 10^{1}8 = 1, 000, 000, 000, 000, 000, 000.

‘E’ ‘EiB’ exbibyte: 2^{6}0 = 1, 152, 921, 504, 606, 846, 976.

‘ZB’ zettabyte: 10^{2}1 = 1, 000, 000, 000, 000, 000, 000, 000 ‘Z’ ‘ZiB’ 2^{7}0 = 1, 180, 591, 620, 717, 411, 303, 424. (‘Zi’ is a GNU extension to IEC 60027-

2.)

‘YB’ yottabyte: 10^{2}4 = 1, 000, 000, 000, 000, 000, 000, 000, 000.

‘Y’ ‘YiB’ 2^{8}0 = 1, 208, 925, 819, 614, 629, 174, 706, 176. (‘Yi’ is a GNU extension to IEC

60027-2.)

Block size defaults can be overridden by an explicit ‘--block-size=size’ option. The ‘-k’ option is equivalent to ‘--block-size=1K’, which is the default unless the POSIXLY_ CORRECT environment variable is set. The ‘-h’ or ‘--human-readable’ option is equivalent to

‘--block-size=human-readable’. The ‘--si’ option is equivalent to ‘--block-size=si’.

2.3 Target directory

Some gnu programs (at least cp, install, ln, and mv) allow you to specify the target directory via this option:

‘--target-directory=directory’

Specify the destination directory. The interface for most programs is that after processing options and a finite (possibly zero) number of fixed-position arguments, the remaining argument list is either expected to be empty, or is a list of items (usually files) that will all be handled identically. The xargs program is designed to work well with this convention. The commands in the mv-family are unusual in that they take a variable number of arguments with a special case at the end (namely, the target directory). This makes it nontrivial to perform some operations, e.g., “move all files from here to ../d/”, because mv * ../d/ might exhaust the argument space, and ls | xargs ... doesn’t have a clean way to specify an extra final argument for each invocation of the subject command. (It can be done by going through a shell command, but that requires more human labor and brain power than it should.) The --target-directory option allows the cp, install, ln, and mv programs to be used conveniently with xargs. For example, you can move the files from the current directory to a sibling directory, d like this: (However, this doesn’t move files whose names begin with ‘.’.)

     ls |xargs mv --target-directory=../d

If you use the gnu find program, you can move all files with this command:

     find . -mindepth 1 -maxdepth 1 \
       | xargs mv --target-directory=../d

But that will fail if there are no files in the current directory or if any file has a name containing a newline character. The following example removes those limitations and requires both gnu find and gnu xargs:

     find . -mindepth 1 -maxdepth 1 -print0 \
       | xargs --null --no-run-if-empty \

                     mv --target-directory=../d

2.4 Trailing slashes

Some gnu programs (at least cp and mv) allow you to remove any trailing slashes from each source argument before operating on it. The --strip-trailing-slashes option enables this behavior.

This is useful when a source argument may have a trailing slash and specify a symbolic link to a directory. This scenario is in fact rather common because some shells can automatically append a trailing slash when performing file name completion on such symbolic links. Without this option, mv, for example, (via the system’s rename function) must interpret a trailing slash as a request to dereference the symbolic link and so must rename the indirectly referenced directory and not the symbolic link. Although it may seem surprising that such behavior be the default, it is required by posix and is consistent with other parts of that standard.

2.5 Standards conformance

In a few cases, the gnu utilities’ default behavior is incompatible with the posix standard. To suppress these incompatibilities, define the POSIXLY_CORRECT environment variable. Unless you are checking for posix conformance, you probably do not need to define

POSIXLY_CORRECT.

Newer versions of posix are occasionally incompatible with older versions. For example, older versions of posix required the command ‘sort +1’ to sort based on the second and succeeding fields in each input line, but starting with posix 1003.1-2001 the same command is required to sort the file named ‘+1’, and you must instead use the command ‘sort -k 2’ to get the field-based sort.

The gnu utilities normally conform to the version of posix that is standard for your system. To cause them to conform to a different version of posix, define the _POSIX2_VERSION environment variable to a value of the form yyyymm specifying the year and month the standard was adopted. Two values are currently supported for _POSIX2_VERSION: ‘199209’ stands for posix 1003.2-1992, and ‘200112’ stands for posix 1003.1-2001. For example, if you are running older software that assumes an older version of posix and uses ‘sort +1’, you can work around the compatibility problems by setting ‘_POSIX2_VERSION=199209’ in your environment.

3 Output of entire files

These commands read and write entire files, possibly transforming them in some way.

3.1 cat: Concatenate and write files

cat copies each file (‘-’ means standard input), or standard input if none are given, to standard output. Synopsis:

cat [option] [file]...

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-A’
‘--show-all’

Equivalent to ‘-vET’.

‘-B’
‘--binary’

On MS-DOS and MS-Windows only, read and write the files in binary mode. By default, cat on MS-DOS/MS-Windows uses binary mode only when standard output is redirected to a file or a pipe; this option overrides that. Binary file I/O is used so that the files retain their format (Unix text as opposed to DOS text and binary), because cat is frequently used as a file-copying program. Some options (see below) cause cat to read and write files in text mode because in those cases the original file contents aren’t important (e.g., when lines are numbered by cat, or when line endings should be marked). This is so these options work as DOS/Windows users would expect; for example, DOS-style text files have their lines end with the CR-LF pair of characters, which won’t be processed as an empty line by ‘-b’ unless the file is read in text mode.

‘-b’
‘--number-nonblank’

Number all nonblank output lines, starting with 1. On MS-DOS and MS-Windows, this option causes cat to read and write files in text mode.

‘-e’ Equivalent to ‘-vE’.

‘-E’
‘--show-ends’

Display a ‘$’ after the end of each line. On MS-DOS and MS-Windows, this option causes cat to read and write files in text mode.

‘-n’
‘--number’

Number all output lines, starting with 1. On MS-DOS and MS-Windows, this option causes cat to read and write files in text mode.

‘-s’
‘--squeeze-blank’

Replace multiple adjacent blank lines with a single blank line. On MS-DOS and MS-Windows, this option causes cat to read and write files in text mode.

‘-t’ Equivalent to ‘-vT’.

‘-T’
‘--show-tabs’

Display TAB characters as ‘^I’.

‘-u’ Ignored; for Unix compatibility.

‘-v’
‘--show-nonprinting’

Display control characters except for LFD and TAB using ‘^’ notation and precede characters that have the high bit set with ‘M-’. On MS-DOS and MS-Windows, this option causes cat to read files and standard input in DOS binary mode, so the CR characters at the end of each line are also visible.

3.2 tac: Concatenate and write files in reverse

tac copies each file (‘-’ means standard input), or standard input if none are given, to standard output, reversing the records (lines by default) in each separately. Synopsis:

     tac [option]... [file]...

Records are separated by instances of a string (newline by default). By default, this separator string is attached to the end of the record that it follows in the file.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--before’

The separator is attached to the beginning of the record that it precedes in the file.

‘-r’

‘--regex’ Treat the separator string as a regular expression. Users of tac on MS-DOS/MS-Windows should note that, since tac reads files in binary mode, each line of a text file might end with a CR/LF pair instead of the Unix-style LF.

‘-s separator’
‘--separator=separator’

Use separator as the record separator, instead of newline.

3.3 nl: Number lines and write files

nl writes each file (‘-’ means standard input), or standard input if none are given, to standard output, with line numbers added to some or all of the lines. Synopsis:

     nl [option]... [file]...

nl decomposes its input into (logical) pages; by default, the line number is reset to 1 at the top of each logical page. nl treats all of the input files as a single document; it does not reset line numbers or logical pages between files.

A logical page consists of three sections: header, body, and footer. Any of the sections can be empty. Each can be numbered in a different style from the others.

The beginnings of the sections of logical pages are indicated in the input file by a line containing exactly one of these delimiter strings:

‘\:\:\:’ start of header;

‘\:\:’ start of body;

‘\:’ start of footer.

The two characters from which these strings are made can be changed from ‘\’ and ‘:’ via options (see below), but the pattern and length of each string cannot be changed.

A section delimiter is replaced by an empty line on output. Any text that comes before the first section delimiter string in the input file is considered to be part of a body section, so nl treats a file that contains no section delimiters as a single body section.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b style’
‘--body-numbering=style’

Select the numbering style for lines in the body section of each logical page. When a line is not numbered, the current line number is not incremented, but the line number separator character is still prepended to the line. The styles are:

‘a’ number all lines,

‘t’ number only nonempty lines (default for body),

‘n’ do not number lines (default for header and footer),

‘pregexp’ number only lines that contain a match for regexp.

‘-d cd’
‘--section-delimiter=cd’

Set the section delimiter characters to cd; default is ‘\:’. If only c is given, the second remains ‘:’. (Remember to protect ‘\’ or other metacharacters from shell expansion with quotes or extra backslashes.)

‘-f style’
‘--footer-numbering=style’
          Analogous to ‘--body-numbering’.
‘-h style’
‘--header-numbering=style’
          Analogous to ‘--body-numbering’.
‘-i number’
‘--page-increment=number’

Increment line numbers by number (default 1).

‘-l number’
‘--join-blank-lines=number’

Consider number (default 1) consecutive empty lines to be one logical line for numbering, and only number the last one. Where fewer than number consecutive empty lines occur, do not number them. An empty line is one that contains no characters, not even spaces or tabs.

‘-n format’
‘--number-format=format’

Select the line numbering format (default is rn):

‘ln’ left justified, no leading zeros;

‘rn’ right justified, no leading zeros;

‘rz’ right justified, leading zeros.

‘-p’
‘--no-renumber’

Do not reset the line number at the start of a logical page.

‘-s string’
‘--number-separator=string’

Separate the line number from the text line in the output with string (default is the TAB character).

‘-v number’
‘--starting-line-number=number’

Set the initial line number on each logical page to number (default 1).

‘-w number’
‘--number-width=number’

Use number characters for line numbers (default 6).

3.4 od: Write files in octal or other formats

od writes an unambiguous representation of each file (‘-’ means standard input), or standard input if none are given. Synopses:

     od [option]... [file]...
     od --traditional [file] [[+]offset [[+]label]]

Each line of output consists of the offset in the input, followed by groups of data from the file. By default, od prints the offset in octal, and each group of file data is two bytes of input printed as a single octal number.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-A radix’
‘--address-radix=radix’

Select the base in which file offsets are printed. radix can be one of the following:

‘d’ decimal; ‘o’ octal;

‘x’ hexadecimal;

‘n’ none (do not print offsets).

The default is octal.

‘-j bytes’
‘--skip-bytes=bytes’

Skip bytes input bytes before formatting and writing. If bytes begins with ‘0x’ or ‘0X’, it is interpreted in hexadecimal; otherwise, if it begins with ‘0’, in octal; otherwise, in decimal. Appending ‘b’ multiplies bytes by 512, ‘k’ by 1024, and ‘m’ by 1048576.

‘-N bytes’
‘--read-bytes=bytes’

Output at most bytes bytes of the input. Prefixes and suffixes on bytes are interpreted as for the ‘-j’ option.

‘-s n’
‘--strings[=n]’

Instead of the normal output, output only string constants: at least n consecutive ascii graphic characters, followed by a null (zero) byte. If n is omitted with ‘--strings’, the default is 3. On older systems, gnu od instead supports an obsolete option ‘-s[n]’, where n also defaults to 3. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow ‘-s’ without an argument; use ‘--strings’ instead.

‘-t type’
‘--format=type’

Select the format in which to output the file data. type is a string of one or more of the below type indicator characters. If you include more than one type indicator character in a single type string, or use this option more than once, od writes one copy of each output line using each of the data types that you specified, in the order that you specified. Adding a trailing “z” to any type specification appends a display of the ascii character representation of the printable characters to the output line generated by the type specification.

‘a’ named character

‘c’ ascii character or backslash escape,

‘d’ signed decimal

‘f’ floating point

‘o’ octal

‘u’ unsigned decimal

‘x’ hexadecimal The type a outputs things like ‘sp’ for space, ‘nl’ for newline, and ‘nul’ for a null (zero) byte. Type c outputs ‘ ’, ‘\n’, and \0, respectively. Except for types ‘a’ and ‘c’, you can specify the number of bytes to use in interpreting each number in the given data type by following the type indicator character with a decimal integer. Alternately, you can specify the size of one of the C compiler’s built-in data types by following the type indicator character with one of the following characters. For integers (‘d’, ‘o’, ‘u’, ‘x’): ‘C’ char ‘S’ short ‘I’ int ‘L’ long For floating point (f): F float D double L long double

‘-v’
‘--output-duplicates’

Output consecutive lines that are identical. By default, when two or more consecutive output lines would be identical, od outputs only the first line, and puts just an asterisk on the following line to indicate the elision.

‘-w n’
‘--width[=n]’

Dump n input bytes per output line. This must be a multiple of the least common multiple of the sizes associated with the specified output types. If this option is not given at all, the default is 16. If n is omitted with ‘--width’, the default is 32. On older systems, gnu od instead supports an obsolete option ‘-w[n]’, where n also defaults to 32. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow ‘-w’ without an argument; use ‘--width’ instead.

The next several options are shorthands for format specifications. gnu od accepts any combination of shorthands and format specification options. These options accumulate. ‘-a’ Output as named characters. Equivalent to ‘-ta’. ‘-b’ Output as octal bytes. Equivalent to ‘-toC’. ‘-c’ Output as ascii characters or backslash escapes. Equivalent to ‘-tc’. ‘-d’ Output as unsigned decimal shorts. Equivalent to ‘-tu2’. ‘-f’ Output as floats. Equivalent to ‘-tfF’. ‘-h’ Output as hexadecimal shorts. Equivalent to ‘-tx2’. ‘-i’ Output as decimal shorts. Equivalent to ‘-td2’.

‘-l’ Output as decimal longs. Equivalent to ‘-td4’.

‘-o’ Output as octal shorts. Equivalent to ‘-to2’.

‘-x’ Output as hexadecimal shorts. Equivalent to ‘-tx2’.

‘--traditional’

Recognize the non-option arguments that traditional od accepted. The following syntax:

     od --traditional [file] [[+]offset[.][b] [[+]label[.][b]]]

can be used to specify at most one file and optional arguments specifying an offset and a pseudo-start address, label. By default, offset is interpreted as an octal number specifying how many input bytes to skip before formatting and writing. The optional trailing decimal point forces the interpretation of offset as a decimal number. If no decimal is specified and the offset begins with ‘0x’ or ‘0X’ it is interpreted as a hexadecimal number. If there is a trailing ‘b’, the number of bytes skipped will be offset multiplied by 512. The label argument is interpreted just like offset, but it specifies an initial pseudo-address. The pseudo-addresses are displayed in parentheses following any normal address.

4 Formatting file contents

These commands reformat the contents of files.

4.1 fmt: Reformat paragraph text

fmt fills and joins lines to produce output lines of (at most) a given number of characters (75 by default). Synopsis:

     fmt [option]... [file]...

fmt reads from the specified file arguments (or standard input if none are given), and writes to standard output.

By default, blank lines, spaces between words, and indentation are preserved in the output; successive input lines with different indentation are not joined; tabs are expanded on input and introduced on output.

fmt prefers breaking lines at the end of a sentence, and tries to avoid line breaks after the first word of a sentence or before the last word of a sentence. A sentence break is defined as either the end of a paragraph or a word ending in any of ‘.?!’, followed by two spaces or end of line, ignoring any intervening parentheses or quotes. Like TEX, fmt reads entire “paragraphs” before choosing line breaks; the algorithm is a variant of that in “Breaking Paragraphs Into Lines” (Donald E. Knuth and Michael F. Plass, Software—Practice and Experience, 11 (1981), 1119–1184).

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c’
‘--crown-margin’

Crown margin mode: preserve the indentation of the first two lines within a paragraph, and align the left margin of each subsequent line with that of the second line.

‘-t’
‘--tagged-paragraph’

Tagged paragraph mode: like crown margin mode, except that if indentation of the first line of a paragraph is the same as the indentation of the second, the first line is treated as a one-line paragraph.

‘-s’
‘--split-only’

Split lines only. Do not join short lines to form longer ones. This prevents sample lines of code, and other such “formatted” text from being unduly combined.

‘-u’
‘--uniform-spacing’

Uniform spacing. Reduce spacing between words to one space, and spacing between sentences to two spaces.

‘-width’
‘-w width’
‘--width=width’

Fill output lines up to width characters (default 75). fmt initially tries to make lines about 7% shorter than this, to give it room to balance line lengths.

‘-p prefix’
‘--prefix=prefix’

Only lines beginning with prefix (possibly preceded by whitespace) are subject to formatting. The prefix and any preceding whitespace are stripped for the formatting and then re-attached to each formatted output line. One use is to format certain kinds of program comments, while leaving the code unchanged.

4.2 pr: Paginate or columnate files for printing

pr writes each file (‘-’ means standard input), or standard input if none are given, to standard output, paginating and optionally outputting in multicolumn format; optionally merges all files, printing all in parallel, one per column. Synopsis:

     pr [option]... [file]...

By default, a 5-line header is printed at each page: two blank lines; a line with the date, the filename, and the page count; and two more blank lines. A footer of five blank lines is also printed. With the ‘-F’ option, a 3-line header is printed: the leading two blank lines are omitted; no footer is used. The default page length in both cases is 66 lines. The default number of text lines changes from 56 (without ‘-F’) to 63 (with ‘-F’). The text line of the header takes the form ‘date string page’, with spaces inserted around string so that the line takes up the full page width. Here, date is the date (see the ‘-D’ or ‘--date-format’ option for details), string is the centered header string, and page identifies the page number. The LC_MESSAGES locale category affects the spelling of page; in the default C locale, it is ‘Page number’ where number is the decimal page number.

Form feeds in the input cause page breaks in the output. Multiple form feeds produce empty pages.

Columns are of equal width, separated by an optional string (default is ‘space’). For multicolumn output, lines will always be truncated to page width (default 72), unless you use the ‘-J’ option. For single column output no line truncation occurs by default. Use ‘-W’ option to truncate lines in that case.

The following changes were made in version 1.22i and apply to later versions of pr: - Brian

• Some small letter options (‘-s’, ‘-w’) have been redefined for better posix compliance. The output of some further cases has been adapted to other Unix systems. These changes are not compatible with earlier versions of the program.

• Some new capital letter options (‘-J’, ‘-S’, ‘-W’) have been introduced to turn off unexpected interferences of small letter options. The ‘-N’ option and the second argument last page of ‘+FIRST_PAGE’ offer more flexibility. The detailed handling of form feeds set in the input files requires the ‘-T’ option.

• Capital letter options override small letter ones.

• Some of the option-arguments (compare ‘-s’, ‘-e’, ‘-i’, ‘-n’) cannot be specified as separate arguments from the preceding option letter (already stated in the posix specification).

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘+first_page[:last_page]’
‘--pages=first_page[:last_page]’

Begin printing with page first page and stop with last page. Missing ‘:last_page’ implies end of file. While estimating the number of skipped pages each form feed in the input file results in a new page. Page counting with and without ‘+first_page’ is identical. By default, counting starts with the first page of input file (not first page printed). Line numbering may be altered by ‘-N’ option.

‘-column’
‘--columns=column’

With each single file, produce column columns of output (default is 1) and print columns down, unless ‘-a’ is used. The column width is automatically decreased as column increases; unless you use the ‘-W/-w’ option to increase page width as well. This option might well cause some lines to be truncated. The number of lines in the columns on each page are balanced. The options ‘-e’ and ‘-i’ are on for multiple text-column output. Together with ‘-J’ option column alignment and line truncation is turned off. Lines of full length are joined in a free field format and ‘-S’ option may set field separators. ‘-column’ may not be used with ‘-m’ option.

‘-a’
‘--across’

With each single file, print columns across rather than down. The ‘-column’ option must be given with column greater than one. If a line is too long to fit in a column, it is truncated.

‘-c’
‘--show-control-chars’

Print control characters using hat notation (e.g., ‘^G’); print other nonprinting characters in octal backslash notation. By default, nonprinting characters are not changed.

‘-d’
‘--double-space’

Double space the output.

‘-D format’
‘--date-format=format’

Format header dates using format, using the same conventions as for the the command ‘date +format’; See Section 21.1 [date invocation], page 117. Except for directives, which start with ‘%’, characters in format are printed unchanged. You can use this option to specify an arbitrary string in place of the header

date, e.g., ‘--date-format="Monday morning"’.

If the POSIXLY_CORRECT environment variable is not set, the date format defaults to ‘%Y-%m-%d %H:%M’ (for example, ‘2001-12-04 23:59’); otherwise, the format depends on the LC_TIME locale category, with the default being ‘%b %e

%H:%M %Y’ (for example, ‘Dec  4 23:59 2001’.
‘-e[in-tabchar[in-tabwidth]]’
‘--expand-tabs[=in-tabchar[in-tabwidth]]’

Expand tabs to spaces on input. Optional argument in-tabchar is the input tab character (default is the TAB character). Second optional argument intabwidth is the input tab character’s width (default is 8).

‘-f’
‘-F’
‘--form-feed’

Use a form feed instead of newlines to separate output pages. The default page length of 66 lines is not altered. But the number of lines of text per page changes from default 56 to 63 lines.

‘-h HEADER’
‘--header=HEADER’

Replace the filename in the header with the centered string header. When using the shell, header should be quoted and should be separated from ‘-h’ by a space.

‘-i[out-tabchar[out-tabwidth]]’
‘--output-tabs[=out-tabchar[out-tabwidth]]’

Replace spaces with tabs on output. Optional argument out-tabchar is the output tab character (default is the TAB character). Second optional argument out-tabwidth is the output tab character’s width (default is 8).

‘-J’
‘--join-lines’

Merge lines of full length. Used together with the column options ‘-column’, ‘-a -column’ or ‘-m’. Turns off ‘-W/-w’ line truncation; no column alignment used; may be used with ‘--sep-string[=string]’. ‘-J’ has been introduced (together with ‘-W’ and ‘--sep-string’) to disentangle the old (posix-compliant) options ‘-w’ and ‘-s’ along with the three column options.

‘-l page_length’
‘--length=page_length’

Set the page length to page length (default 66) lines, including the lines of the header [and the footer]. If page length is less than or equal to 10 (or <= 3 with ‘-F’), the header and footer are omitted, and all form feeds set in input files are eliminated, as if the ‘-T’ option had been given.

‘-m’

‘--merge’ Merge and print all files in parallel, one in each column. If a line is too long to fit in a column, it is truncated, unless the ‘-J’ option is used. ‘--sep-string[=string]’ may be used. Empty pages in some files (form feeds set) produce empty columns, still marked by string. The result is a continuous line numbering and column marking throughout the whole merged file. Completely empty merged pages show no separators or line numbers. The default header becomes ‘date page’ with spaces inserted in the middle; this may be used with the ‘-h’ or ‘--header’ option to fill up the middle blank part.

‘-n[number-separator[digits]]’
‘--number-lines[=number-separator[digits]]’

Provide digits digit line numbering (default for digits is 5). With multicolumn output the number occupies the first digits column positions of each text column or only each line of ‘-m’ output. With single column output the number precedes each line just as ‘-m’ does. Default counting of the line numbers starts with the first line of the input file (not the first line printed, compare the ‘--page’ option and ‘-N’ option). Optional argument number-separator is the character appended to the line number to separate it from the text followed. The default separator is the TAB character. In a strict sense a TAB is always printed with single column output only. The TAB-width varies with the TAB-position, e.g. with the left margin specified by ‘-o’ option. With multicolumn output priority is given to ‘equal width of output columns’ (a posix specification). The TAB-width is fixed to the value of the first column and does not change with different values of left margin. That means a fixed number of spaces is always printed in the place of the number-separator tab. The tabification depends upon the output position.

‘-N line_number’
‘--first-line-number=line_number’

Start line counting with the number line number at first line of first page printed (in most cases not the first line of the input file).

‘-o margin’
‘--indent=margin’

Indent each line with a margin margin spaces wide (default is zero). The total page width is the size of the margin plus the page width set with the ‘-W/-w’ option. A limited overflow may occur with numbered single column output (compare ‘-n’ option).

‘-r’
‘--no-file-warnings’

Do not print a warning message when an argument file cannot be opened. (The exit status will still be nonzero, however.)

‘-s[char]’
‘--separator[=char]’

Separate columns by a single character char. The default for char is the TAB character without ‘-w’ and ‘no character’ with ‘-w’. Without ‘-s’ the default separator ‘space’ is set. ‘-s[char]’ turns off line truncation of all three column options (‘-COLUMN’|‘-a -COLUMN’|‘-m’) unless ‘-w’ is set. This is a posixcompliant formulation.

‘-S string’
‘--sep-string[=string]’

Use string to separate output columns. The ‘-S’ option doesn’t affect the ‘-W/-w’ option, unlike the ‘-s’ option which does. It does not affect line truncation or column alignment. Without ‘-S’, and with ‘-J’, pr uses the default output separator, TAB. Without ‘-S’ or ‘-J’, pr uses a ‘space’ (same as ‘-S" "’). With ‘-Sstring’, string must be nonempty; ‘--sep-string’ with no string is

equivalent to ‘--sep-string=""’.

On older systems, pr instead supports an obsolete option ‘-S[string]’, where string is optional. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this older usage. To specify an empty string portably,

use ‘--sep-string’.
‘-t’
‘--omit-header’

Do not print the usual header [and footer] on each page, and do not fill out the bottom of pages (with blank lines or a form feed). No page structure is produced, but form feeds set in the input files are retained. The predefined pagination is not changed. ‘-t’ or ‘-T’ may be useful together with other options; e.g.: ‘-t -e4’, expand TAB characters in the input file to 4 spaces but don’t make any other changes. Use of ‘-t’ overrides ‘-h’.

‘-T’
‘--omit-pagination’

Do not print header [and footer]. In addition eliminate all form feeds set in the input files.

‘-v’
‘--show-nonprinting’

Print nonprinting characters in octal backslash notation.

‘-w page_width’
‘--width=page_width’

Set page width to page width characters for multiple text-column output only (default for page width is 72). ‘-s[CHAR]’ turns off the default page width and any line truncation and column alignment. Lines of full length are merged, regardless of the column options set. No page width setting is possible with single column output. A posix-compliant formulation.

‘-W page_width’
‘--page_width=page_width’

Set the page width to page width characters. That’s valid with and without a column option. Text lines are truncated, unless ‘-J’ is used. Together with one of the three column options (‘-column’, ‘-a -column’ or ‘-m’) column alignment is always used. The separator options ‘-S’ or ‘-s’ don’t affect the ‘-W’ option. Default is 72 characters. Without ‘-W page_width’ and without any of the column options NO line truncation is used (defined to keep downward compatibility and to meet most frequent tasks). That’s equivalent to ‘-W 72 -J’. The header line is never truncated.

4.3 fold: Wrap input lines to fit in specified width

fold writes each file (‘-’ means standard input), or standard input if none are given, to standard output, breaking long lines. Synopsis:

     fold [option]... [file]...

By default, fold breaks lines wider than 80 columns. The output is split into as many lines as necessary.

fold counts screen columns by default; thus, a tab may count more than one column, backspace decreases the column count, and carriage return sets the column to zero.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’

‘--bytes’ Count bytes rather than columns, so that tabs, backspaces, and carriage returns are each counted as taking up one column, just like other characters.

‘-s’
‘--spaces’

Break at word boundaries: the line is broken after the last blank before the maximum line length. If the line contains no such blanks, the line is broken at the maximum line length as usual.

‘-w width’
‘--width=width’

Use a maximum line length of width columns instead of 80.

On older systems, fold supports an obsolete option ‘-width’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-w width’ instead.

5 Output of parts of files

These commands output pieces of the input.

5.1 head: Output the first part of files

head prints the first part (10 lines by default) of each file; it reads from standard input if no files are given or when given a file of ‘-’. Synopsis:

     head [option]... [file]...

If more than one file is specified, head prints a one-line header consisting of

     ==> file name <==

before the output for each file.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c bytes’
‘--bytes=bytes’

Print the first bytes bytes, instead of initial lines. Appending ‘b’ multiplies bytes by 512, ‘k’ by 1024, and ‘m’ by 1048576.

‘-n n’
‘--lines=n’

Output the first n lines.

‘-q’
‘--quiet’
‘--silent’

Never print file name headers.

‘-v’
‘--verbose’

Always print file name headers.

On older systems, head supports an obsolete option ‘-countoptions’, which is recognized only if it is specified first. count is a decimal number optionally followed by a size letter (‘b’, ‘k’, ‘m’) as in -c, or ‘l’ to mean count by lines, or other option letters (‘cqv’). posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-c count’ or ‘-n count’ instead.

5.2 tail: Output the last part of files

tail prints the last part (10 lines by default) of each file; it reads from standard input if no files are given or when given a file of ‘-’. Synopsis:

     tail [option]... [file]...

If more than one file is specified, tail prints a one-line header consisting of

==> file name <==

before the output for each file.

gnu tail can output any amount of data (some other versions of tail cannot). It also has no ‘-r’ option (print in reverse), since reversing a file is really a different job from printing the end of a file; BSD tail (which is the one with -r) can only reverse files that are at most as large as its buffer, which is typically 32 KiB. A more reliable and versatile way to reverse files is the gnu tac command.

If any option-argument is a number n starting with a ‘+’, tail begins printing with the nth item from the start of each file, instead of from the end.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c bytes’
‘--bytes=bytes’

Output the last bytes bytes, instead of final lines. Appending ‘b’ multiplies bytes by 512, ‘k’ by 1024, and ‘m’ by 1048576.

‘-f’
‘--follow[=how]’

Loop forever trying to read more characters at the end of the file, presumably because the file is growing. This option is ignored when reading from a pipe. If more than one file is given, tail prints a header whenever it gets output from a different file, to indicate which file that output is from.

There are two ways to specify how you’d like to track files with this option, but that difference is noticeable only when a followed file is removed or renamed. If you’d like to continue to track the end of a growing file even after it has been unlinked, use ‘--follow=descriptor’. This is the default behavior, but it is not useful if you’re tracking a log file that may be rotated (removed or renamed, then reopened). In that case, use ‘--follow=name’ to track the named file by reopening it periodically to see if it has been removed and recreated by some other program.

No matter which method you use, if the tracked file is determined to have shrunk, tail prints a message saying the file has been truncated and resumes tracking the end of the file from the newly-determined endpoint.

When a file is removed, tail’s behavior depends on whether it is following the name or the descriptor. When following by name, tail can detect that a file has been removed and gives a message to that effect, and if ‘--retry’ has been specified it will continue checking periodically to see if the file reappears. When following a descriptor, tail does not detect that the file has been unlinked or renamed and issues no message; even though the file may no longer be accessible via its original name, it may still be growing.

The option values ‘descriptor’ and ‘name’ may be specified only with the long form of the option, not with ‘-f’.

‘-F’ This option is the same as ‘--follow=name --retry’. That is, tail will attempt to reopen a file when it is removed. Should this fail, tail will keep trying until it becomes accessible again.

‘--retry’ This option is meaningful only when following by name. Without this option, when tail encounters a file that doesn’t exist or is otherwise inaccessible, it reports that fact and never checks it again.

‘--sleep-interval=number’

Change the number of seconds to wait between iterations (the default is 1.0). During one iteration, every specified file is checked to see if it has Historical implementations of tail have required that number be an integer. However, GNU tail accepts an arbitrary floating point number.

‘--pid=pid’

When following by name or by descriptor, you may specify the process ID, pid, of the sole writer of all file arguments. Then, shortly after that process terminates, tail will also terminate. This will work properly only if the writer and the tailing process are running on the same machine. For example, to save the output of a build in a file and to watch the file grow, if you invoke make and tail like this then the tail process will stop when your build completes. Without this option, you would have had to kill the tail -f process yourself.

     $ make >& makerr & tail --pid=$! -f makerr

If you specify a pid that is not in use or that does not correspond to the process that is writing to the tailed files, then tail may terminate long before any files stop growing or it may not terminate until long after the real writer has terminated. Note that ‘--pid’ cannot be supported on some systems; tail will print a warning if this is the case.

‘--max-unchanged-stats=n’

When tailing a file by name, if there have been n (default n=5) consecutive iterations for which the size has remained the same, then open/fstat the file to determine if that file name is still associated with the same device/inode-number pair as before. When following a log file that is rotated, this is approximately the number of seconds between when tail prints the last pre-rotation lines and when it prints the lines that have accumulated in the new log file. This option is meaningful only when following by name.

‘-n n’
‘--lines=n’

Output the last n lines.

‘-q’
‘--quiet’
‘--silent’

Never print file name headers.

‘-v’
‘--verbose’

Always print file name headers.

On older systems, tail supports an obsolete option ‘-countoptions’, which is recognized only if it is specified first. count is a decimal number optionally followed by a size letter (‘b’, ‘k’, ‘m’) as in -c, or ‘l’ to mean count by lines, or other option letters (‘cfqv’). Some older tail implementations also support an obsolete option ‘+count’ with the same meaning as ‘-+count’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow these options; use ‘-c count’ or ‘-n count’ instead.

5.3 split: Split a file into fixed-size pieces

split creates output files containing consecutive sections of input (standard input if none is given or input is ‘-’). Synopsis:

     split [option] [input [prefix]]

By default, split puts 1000 lines of input (or whatever is left over for the last section), into each output file.

The output files’ names consist of prefix (‘x’ by default) followed by a group of letters (‘aa’, ‘ab’, … by default), such that concatenating the output files in sorted order by file name produces the original input file. If the output file names are exhausted, split reports an error without deleting the output files that it did create.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a length’
‘--suffix-length=length’

Use suffixes of length length. The default length is 2.

‘-l lines’
‘--lines=lines’

Put lines lines of input into each output file.

On older systems, split supports an obsolete option ‘-lines’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-l lines’ instead.

‘-b bytes’
‘--bytes=bytes’

Put the first bytes bytes of input into each output file. Appending ‘b’ multiplies bytes by 512, ‘k’ by 1024, and ‘m’ by 1048576.

‘-C bytes’
‘--line-bytes=bytes’

Put into each output file as many complete lines of input as possible without exceeding bytes bytes. For lines longer than bytes bytes, put bytes bytes into each output file until less than bytes bytes of the line are left, then continue normally. bytes has the same format as for the ‘--bytes’ option.

‘--verbose’

Write a diagnostic to standard error just before each output file is opened.

5.4 csplit: Split a file into context-determined pieces

csplit creates zero or more output files containing sections of input (standard input if input is ‘-’). Synopsis:

     csplit [option]... input pattern...

The contents of the output files are determined by the pattern arguments, as detailed below. An error occurs if a pattern argument refers to a nonexistent line of the input file (e.g., if no remaining line matches a given regular expression). After every pattern has been matched, any remaining input is copied into one last output file.

By default, csplit prints the number of bytes written to each output file after it has been created.

The types of pattern arguments are:

‘n’ Create an output file containing the input up to but not including line n (a positive integer). If followed by a repeat count, also create an output file containing the next line lines of the input file once for each repeat.

‘/regexp/[offset]’

Create an output file containing the current line up to (but not including) the next line of the input file that contains a match for regexp. The optional offset is a ‘+’ or ‘-’ followed by a positive integer. If it is given, the input up to the matching line plus or minus offset is put into the output file, and the line after that begins the next section of input.

‘%regexp%[offset]’

Like the previous type, except that it does not create an output file, so that section of the input file is effectively ignored.

‘{repeat-count}’

Repeat the previous pattern repeat-count additional times. repeat-count can either be a positive integer or an asterisk, meaning repeat as many times as necessary until the input is exhausted.

The output files’ names consist of a prefix (‘xx’ by default) followed by a suffix. By default, the suffix is an ascending sequence of two-digit decimal numbers from ‘00’ to ‘99’. In any case, concatenating the output files in sorted order by filename produces the original input file.

By default, if csplit encounters an error or receives a hangup, interrupt, quit, or terminate signal, it removes any output files that it has created so far before it exits.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-f prefix’
‘--prefix=prefix’

Use prefix as the output file name prefix.

‘-b suffix’
‘--suffix=suffix’

Use suffix as the output file name suffix. When this option is specified, the suffix string must include exactly one printf(3)-style conversion specification, possibly including format specification flags, a field width, a precision specifications, or all of these kinds of modifiers. The format letter must convert a binary integer argument to readable form; thus, only ‘d’, ‘i’, ‘u’, ‘o’, ‘x’, and ‘X’ conversions are allowed. The entire suffix is given (with the current output file number) to sprintf(3) to form the file name suffixes for each of the individual output files in turn. If this option is used, the ‘--digits’ option is ignored.

‘-n digits’
‘--digits=digits’

Use output file names containing numbers that are digits digits long instead of the default 2.

‘-k’
‘--keep-files’

Do not remove output files when errors are encountered.

‘-z’
‘--elide-empty-files’

Suppress the generation of zero-length output files. (In cases where the section delimiters of the input file are supposed to mark the first lines of each of the sections, the first output file will generally be a zero-length file unless you use this option.) The output file sequence numbers always run consecutively starting from 0, even when this option is specified.

‘-s’
‘-q’
‘--silent’

‘--quiet’ Do not print counts of output file sizes.

6 Summarizing files

These commands generate just a few numbers representing entire contents of files.

6.1 wc: Print byte, word, and line counts

wc counts the number of bytes, characters, whitespace-separated words, and newlines in each given file, or standard input if none are given or for a file of ‘-’. Synopsis:

     wc [option]... [file]...

wc prints one line of counts for each file, and if the file was given as an argument, it prints the file name following the counts. If more than one file is given, wc prints a final line containing the cumulative counts, with the file name ‘total’. The counts are printed in this order: newlines, words, characters, bytes. By default, each count is output right-justified in a 7-byte field with one space between fields so that the numbers and file names line up nicely in columns. However, posix requires that there be exactly one space separating columns. You can make wc use the posix-mandated output format by setting the POSIXLY_CORRECT environment variable.

By default, wc prints three counts: the newline, words, and byte counts. Options can specify that only certain counts be printed. Options do not undo others previously given, so

     wc --bytes --words

prints both the byte counts and the word counts.

With the --max-line-length option, wc prints the length of the longest line per file, and if there is more than one file it prints the maximum (not the sum) of those lengths.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c’

‘--bytes’ Print only the byte counts.

‘-m’

‘--chars’ Print only the character counts.

‘-w’

‘--words’ Print only the word counts.

‘-l’

‘--lines’ Print only the newline counts.

‘-L’
‘--max-line-length’

Print only the maximum line lengths.

6.2 sum: Print checksum and block counts

sum computes a 16-bit checksum for each given file, or standard input if none are given or for a file of ‘-’. Synopsis:

     sum [option]... [file]...

sum prints the checksum for each file followed by the number of blocks in the file (rounded up). If more than one file is given, file names are also printed (by default). (With the ‘--sysv’ option, corresponding file names are printed when there is at least one file argument.)

By default, gnu sum computes checksums using an algorithm compatible with BSD sum and prints file sizes in units of 1024-byte blocks.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-r’ Use the default (BSD compatible) algorithm. This option is included for compatibility with the System V sum. Unless ‘-s’ was also given, it has no effect.

‘-s’

‘--sysv’ Compute checksums using an algorithm compatible with System V sum’s default, and print file sizes in units of 512-byte blocks.

sum is provided for compatibility; the cksum program (see next section) is preferable in new applications.

6.3 cksum: Print CRC checksum and byte counts

cksum computes a cyclic redundancy check (CRC) checksum for each given file, or standard input if none are given or for a file of ‘-’. Synopsis:

     cksum [option]... [file]...

cksum prints the CRC checksum for each file along with the number of bytes in the file, and the filename unless no arguments were given.

cksum is typically used to ensure that files transferred by unreliable means (e.g., netnews) have not been corrupted, by comparing the cksum output for the received files with the cksum output for the original files (typically given in the distribution).

The CRC algorithm is specified by the posix standard. It is not compatible with the BSD or System V sum algorithms (see the previous section); it is more robust.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

6.4 md5sum: Print or check message-digests

md5sum computes a 128-bit checksum (or fingerprint or message-digest) for each specified file. If a file is specified as ‘-’ or if no files are given md5sum computes the checksum for the standard input. md5sum can also determine whether a file and checksum are consistent. Synopses:

md5sum [option]... [file]...
md5sum [option]... --check [file]

For each file, ‘md5sum’ outputs the MD5 checksum, a flag indicating a binary or text input file, and the filename. If file is omitted or specified as ‘-’, standard input is read.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--binary’

Treat all input files as binary. This option has no effect on Unix systems, since they don’t distinguish between binary and text files. This option is useful on systems that have different internal and external character representations. On MS-DOS and MS-Windows, this is the default.

‘-c’

‘--check’ Read filenames and checksum information from the single file (or from stdin if no file was specified) and report whether each named file and the corresponding checksum data are consistent. The input to this mode of md5sum is usually the output of a prior, checksum-generating run of ‘md5sum’. Each valid line of input consists of an MD5 checksum, a binary/text flag, and then a filename. Binary files are marked with ‘*’, text with ‘ ’. For each such line, md5sum reads the named file and computes its MD5 checksum. Then, if the computed message digest does not match the one on the line with the filename, the file is noted as having failed the test. Otherwise, the file passes the test. By default, for each valid line, one line is written to standard output indicating whether the named file passed the test. After all checks have been performed, if there were any failures, a warning is issued to standard error. Use the ‘--status’ option to inhibit that output. If any listed file cannot be opened or read, if any valid line has an MD5 checksum inconsistent with the associated file, or if no valid line is found, md5sum exits with nonzero status. Otherwise, it exits successfully.

‘--status’

This option is useful only when verifying checksums. When verifying checksums, don’t generate the default one-line-per-file diagnostic and don’t output the warning summarizing any failures. Failures to open or read a file still evoke individual diagnostics to standard error. If all listed files are readable and are consistent with the associated MD5 checksums, exit successfully. Otherwise exit with a status code indicating there was a failure.

‘-t’

‘--text’ Treat all input files as text files. This is the reverse of ‘--binary’.

‘-w’

‘--warn’ When verifying checksums, warn about improperly formatted MD5 checksum lines. This option is useful only if all but a few lines in the checked input are valid.

7 Operating on sorted files

These commands work with (or produce) sorted files.

7.1 sort: Sort text files

sort sorts, merges, or compares all the lines from the given files, or standard input if none are given or for a file of ‘-’. By default, sort writes the results to standard output. Synopsis:

     sort [option]... [file]...

sort has three modes of operation: sort (the default), merge, and check for sortedness. The following options change the operation mode:

‘-c’

‘--check’ Check whether the given files are already sorted: if they are not all sorted, print an error message and exit with a status of 1. Otherwise, exit successfully.

‘-m’

‘--merge’ Merge the given files by sorting them as a group. Each input file must always be individually sorted. It always works to sort instead of merge; merging is provided because it is faster, in the case where it works.

A pair of lines is compared as follows: if any key fields have been specified, sort compares each pair of fields, in the order specified on the command line, according to the associated ordering options, until a difference is found or no fields are left. Unless otherwise specified, all comparisons use the character collating sequence specified by the LC_COLLATE locale.^{1}

If any of the global options ‘bdfgiMnr’ are given but no key fields are specified, sort compares the entire lines according to the global options.

Finally, as a last resort when all keys compare equal (or if no ordering options were specified at all), sort compares the entire lines. The last resort comparison honors the ‘--reverse’ (‘-r’) global option. The ‘--stable’ (‘-s’) option disables this last-resort comparison so that lines in which all fields compare equal are left in their original relative order. If no fields or global options are specified, ‘--stable’ (‘-s’) has no effect.

gnu sort (as specified for all gnu utilities) has no limits on input line length or restrictions on bytes allowed within lines. In addition, if the final byte of an input file is not a newline, gnu sort silently supplies one. A line’s trailing newline is not part of the line for comparison purposes.

Upon any error, sort exits with a status of ‘2’.

If the environment variable TMPDIR is set, sort uses its value as the directory for temporary files instead of ‘/tmp’. The ‘--temporary-directory’ (‘-T’) option in turn overrides the environment variable.

1 If you use a non-posix locale (e.g., by setting LC_ALL to ‘en_US’), then sort may produce output that is

sorted differently than you’re accustomed to. In that case, set the LC_ALL environment variable to ‘C’.

Note that setting only LC_COLLATE has two problems. First, it is ineffective if LC_ALL is also set. Second,

it has undefined behavior if LC_CTYPE (or LANG, if LC_CTYPE is unset) is set to an incompatible value. For

example, you get undefined behavior if LC_CTYPE is ja_JP.PCK but LC_COLLATE is en_US.UTF-8.

The following options affect the ordering of output lines. They may be specified globally or as part of a specific key field. If no key fields are specified, global options apply to comparison of entire lines; otherwise the global options are inherited by key fields that do not specify any special options of their own. In pre-posix versions of sort, global options affect only later key fields, so portable shell scripts should specify global options first.

‘-b’
‘--ignore-leading-blanks’

Ignore leading blanks when finding sort keys in each line. The LC_CTYPE locale determines character types.

‘-d’
‘--dictionary-order’

Sort in phone directory order: ignore all characters except letters, digits and blanks when sorting. The LC_CTYPE locale determines character types.

‘-f’
‘--ignore-case’

Fold lowercase characters into the equivalent uppercase characters when comparing so that, for example, ‘b’ and ‘B’ sort as equal. The LC_CTYPE locale determines character types.

‘-g’
‘--general-numeric-sort’

Sort numerically, using the standard C function strtod to convert a prefix of each line to a double-precision floating point number. This allows floating point numbers to be specified in scientific notation, like 1.0e-34 and 10e100. The LC_NUMERIC locale determines the decimal-point character. Do not report overflow, underflow, or conversion errors. Use the following collating sequence:

• Lines that do not start with numbers (all considered to be equal).

• NaNs (“Not a Number” values, in IEEE floating point arithmetic) in a consistent but machine-dependent order.

• Minus infinity.

• Finite numbers in ascending numeric order (with −0 and +0 equal).

• Plus infinity.

Use this option only if there is no alternative; it is much slower than ‘--numeric-sort’ (‘-n’) and it can lose information when converting to floating point.

‘-i’
‘--ignore-nonprinting’

Ignore nonprinting characters. The LC_CTYPE locale determines character types.

‘-M’
‘--month-sort’

An initial string, consisting of any amount of whitespace, followed by a month name abbreviation, is folded to UPPER case and compared in the order ‘JAN’ < ‘FEB’ < … < ‘DEC’. Invalid names compare low to valid names. The LC_TIME locale category determines the month spellings.

‘-n’
‘--numeric-sort’

Sort numerically: the number begins each line; specifically, it consists of optional whitespace, an optional ‘-’ sign, and zero or more digits possibly separated by thousands separators, optionally followed by a decimal-point character and zero or more digits. The LC_NUMERIC locale specifies the decimal-point character and thousands separator.

Numeric sort uses what might be considered an unconventional method to compare strings representing floating point numbers. Rather than first converting each string to the C double type and then comparing those values, sort aligns the decimal-point characters in the two strings and compares the strings a character at a time. One benefit of using this approach is its speed. In practice this is much more efficient than performing the two corresponding string-to-double (or even string-to-integer) conversions and then comparing doubles. In addition, there is no corresponding loss of precision. Converting each string to double before comparison would limit precision to about 16 digits on most systems.

Neither a leading ‘+’ nor exponential notation is recognized. To compare such strings numerically, use the ‘--general-numeric-sort’ (‘-g’) option.

‘-r’
‘--reverse’

Reverse the result of comparison, so that lines with greater key values appear earlier in the output instead of later.

Other options are:

‘-o output-file’
‘--output=output-file’

Write output to output-file instead of standard output. If necessary, sort reads input before opening output-file, so you can safely sort a file in place by using

commands like sort -o F F and cat F | sort -o F.

On newer systems, ‘-o’ cannot appear after an input file if POSIXLY_CORRECT is set, e.g., ‘sort F -o F’. Portable scripts should specify ‘-o output-file’ before any input files.

‘-S size’
‘--buffer-size=size’

Use a main-memory sort buffer of the given size. By default, size is in units of 1024 bytes. Appending ‘%’ causes size to be interpreted as a percentage of physical memory. Appending ‘K’ multiplies size by 1024 (the default), ‘M’ by 1,048,576, ‘G’ by 1,073,741,824, and so on for ‘T’, ‘P’, ‘E’, ‘Z’, and ‘Y’. Appending ‘b’ causes size to be interpreted as a byte count, with no multiplication.

This option can improve the performance of sort by causing it to start with a larger or smaller sort buffer than the default. However, this option affects only the initial buffer size. The buffer grows beyond size if sort encounters input lines larger than size.

‘-t separator’
‘--field-separator=separator’

Use character separator as the field separator when finding the sort keys in each line. By default, fields are separated by the empty string between a nonwhitespace character and a whitespace character. That is, given the input line ‘ foo bar’, sort breaks it into fields ‘ foo’ and ‘ bar’. The field separator is not considered to be part of either the field preceding or the field following. But note that sort fields that extend to the end of the line, as ‘-k 2’, or sort fields consisting of a range, as ‘-k 2,3’, retain the field separators present between the endpoints of the range.

‘-T tempdir’
‘--temporary-directory=tempdir’

Use directory tempdir to store temporary files, overriding the TMPDIR environment variable. If this option is given more than once, temporary files are stored in all the directories given. If you have a large sort or merge that is I/O-bound, you can often improve performance by using this option to specify directories on different disks and controllers.

‘-u’
‘--unique’

Normally, output only the first of a sequence of lines that compare equal. For the ‘--check’ (‘-c’) option, check that no pair of consecutive lines compares equal.

‘-k pos1[,pos2]’
‘--key=pos1[,pos2]’

Specify a sort field that consists of the part of the line between pos1 and pos2 (or the end of the line, if pos2 is omitted), inclusive. Fields and character positions are numbered starting with 1. So to sort on the second field, you’d use ‘--key=2,2’ (‘-k 2,2’). See below for more examples.

‘-z’
‘--zero-terminated’

Treat the input as a set of lines, each terminated by a zero byte (ascii nul (Null) character) instead of an ascii lf (Line Feed). This option can be useful in conjunction with ‘perl -0’ or ‘find -print0’ and ‘xargs -0’ which do the same in order to reliably handle arbitrary pathnames (even those which contain Line Feed characters.)

Historical (BSD and System V) implementations of sort have differed in their interpretation of some options, particularly ‘-b’, ‘-f’, and ‘-n’. gnu sort follows the posix behavior, which is usually (but not always!) like the System V behavior. According to posix, ‘-n’ no longer implies ‘-b’. For consistency, ‘-M’ has been changed in the same way. This may affect the meaning of character positions in field specifications in obscure cases. The only fix is to add an explicit ‘-b’.

A position in a sort field specified with the ‘-k’ option has the form ‘f.c’, where f is the number of the field to use and c is the number of the first character from the beginning of the field. In a start position, an omitted ‘.c’ stands for the field’s first character. In an end position, an omitted or zero ‘.c’ stands for the field’s last character. If the ‘-b’ option was specified, the ‘.c’ part of a field specification is counted from the first nonblank character of the field.

A sort key position may also have any of the option letters ‘Mbdfinr’ appended to it, in which case the global ordering options are not used for that particular field. The ‘-b’ option may be independently attached to either or both of the start and end positions of a field specification, and if it is inherited from the global options it will be attached to both. Keys may span multiple fields.

On older systems, sort supports an obsolete origin-zero syntax ‘+pos1 [-pos2]’ for specifying sort keys. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-k’ instead.

Here are some examples to illustrate various combinations of options.

• Sort in descending (reverse) numeric order.

     sort -nr

• Sort alphabetically, omitting the first and second fields. This uses a single key composed of the characters beginning at the start of field three and extending to the end of each line.

     sort -k 3

• Sort numerically on the second field and resolve ties by sorting alphabetically on the

third and fourth characters of field five. Use ‘:’ as the field delimiter.

     sort -t : -k 2,2n -k 5.3,5.4

Note that if you had written ‘-k 2’ instead of ‘-k 2,2’ sort would have used all characters beginning in the second field and extending to the end of the line as the primary numeric key. For the large majority of applications, treating keys spanning more than one field as numeric will not do what you expect. Also note that the ‘n’ modifier was applied to the field-end specifier for the first key. It would have been equivalent to specify ‘-k 2n,2’ or ‘-k 2n,2n’. All modifiers except ‘b’ apply to the associated field, regardless of whether the modifier character is attached to the field-start and/or the field-end part of the key specifier.

• Sort the password file on the fifth field and ignore any leading white space. Sort lines

with equal values in field five on the numeric user ID in field three.

     sort -t : -k 5b,5 -k 3,3n /etc/passwd

An alternative is to use the global numeric modifier ‘-n’.

     sort -t : -n -k 5b,5 -k 3,3 /etc/passwd

• Generate a tags file in case-insensitive sorted order.

     find src -type f -print0 | sort -t / -z -f | xargs -0 etags --append

The use of ‘-print0’, ‘-z’, and ‘-0’ in this case means that pathnames that contain Line Feed characters will not get broken up by the sort operation. Finally, to ignore both leading and trailing white space, you could have applied the ‘b’ modifier to the field-end specifier for the first key,

     sort -t : -n -k 5b,5b -k 3,3 /etc/passwd

or by using the global ‘-b’ modifier instead of ‘-n’ and an explicit ‘n’ with the second key specifier.

         sort -t : -b -k 5,5 -k 3,3n /etc/passwd

7.2 uniq: Uniquify files

uniq writes the unique lines in the given ‘input’, or standard input if nothing is given or for an input name of ‘-’. Synopsis:

     uniq [option]... [input [output]]

By default, uniq prints the unique lines in a sorted file, i.e., discards all but one of identical successive lines. Optionally, it can instead show only lines that appear exactly once, or lines that appear more than once.

The input need not be sorted, but duplicate input lines are detected only if they are adjacent. If you want to discard non-adjacent duplicate lines, perhaps you want to use sort

-u.

Comparisons use the character collating sequence specified by the LC_COLLATE locale category.

If no output file is specified, uniq writes to standard output.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-f n’
‘--skip-fields=n’

Skip n fields on each line before checking for uniqueness. Fields are sequences of non-space non-tab characters that are separated from each other by at least one space or tab. On older systems, uniq supports an obsolete option ‘-n’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-f n’ instead.

‘-s n’
‘--skip-chars=n’

Skip n characters before checking for uniqueness. If you use both the field and character skipping options, fields are skipped over first. On older systems, uniq supports an obsolete option ‘+n’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-s n’ instead.

‘-c’

‘--count’ Print the number of times each line occurred along with the line.

‘-i’
‘--ignore-case’

Ignore differences in case when comparing lines.

‘-d’
‘--repeated’

Print one copy of each duplicate line.

‘-D’
‘--all-repeated[=delimit-method]’

Print all copies of each duplicate line. This option is useful mainly in conjunction with other options e.g., to ignore case or to compare only selected fields. The optional delimit-method tells how to delimit groups of duplicate lines, and must be one of the following:

‘none’ Do not delimit groups of duplicate lines. This is equivalent to

          ‘--all-repeated’ (‘-D’).

‘prepend’ Output a newline before each group of duplicate lines.

‘separate’

Separate groups of duplicate lines with a single newline. This is the same as using ‘prepend’, except that there is no newline before the first group, and hence may be better suited for output direct to users.

Note that when groups are delimited and the input stream contains two or more consecutive blank lines, then the output is ambiguous. To avoid that, filter the input through ‘tr -s '\n'’ to replace each sequence of consecutive newlines with a single newline. This is a gnu extension.

‘-u’
‘--unique’

Print non-duplicate lines.

‘-w n’
‘--check-chars=n’

Compare n characters on each line (after skipping any specified fields and characters). By default the entire rest of the lines are compared.

7.3 comm: Compare two sorted files line by line

comm writes to standard output lines that are common, and lines that are unique, to two input files; a file name of ‘-’ means standard input. Synopsis:

comm [option]... file1 file2

Before comm can be used, the input files must be sorted using the collating sequence specified by the LC_COLLATE locale. If an input file ends in a non-newline character, a newline is silently appended. The sort command with no options always outputs a file that is suitable input to comm.

With no options, comm produces three column output. Column one contains lines unique to file1, column two contains lines unique to file2, and column three contains lines common to both files. Columns are separated by a single TAB character.

The options ‘-1’, ‘-2’, and ‘-3’ suppress printing of the corresponding columns. Also see Chapter 2 [Common options], page 2.

Unlike some other comparison utilities, comm has an exit status that does not depend on the result of the comparison. Upon normal completion comm produces an exit code of zero. If there is an error it exits with nonzero status.

7.4 tsort: Topological sort

tsort performs a topological sort on the given file, or standard input if no input file is given or for a file of ‘-’. For more details and some history, see Section 7.5 [tsort background], page 39. Synopsis:

tsort [option] [file]

tsort reads its input as pairs of strings, separated by blanks, indicating a partial ordering. The output is a total ordering that corresponds to the given partial ordering.

For example

tsort <<EOF
a b c
d
e f
b c d e
EOF

will produce the output

a
b
c
d
e
f

Consider a more realistic example. You have a large set of functions all in one file, and they may all be declared static except one. Currently that one (say main) is the first function defined in the file, and the ones it calls directly follow it, followed by those they call, etc. Let’s say that you are determined to take advantage of prototypes, so you have to choose between declaring all of those functions (which means duplicating a lot of information from the definitions) and rearranging the functions so that as many as possible are defined before they are used. One way to automate the latter process is to get a list for each function of the functions it calls directly. Many programs can generate such lists. They describe a call graph. Consider the following list, in which a given line indicates that the function on the left calls the one on the right directly.

main parse_options
main tail_file
main tail_forever
tail_file pretty_name
tail_file write_header
tail_file tail
tail_forever recheck
tail_forever pretty_name

tail_forever write_header
tail_forever dump_remainder
tail tail_lines
tail tail_bytes
tail_lines start_lines
tail_lines dump_remainder
tail_lines file_lines
tail_lines pipe_lines
tail_bytes xlseek
tail_bytes start_bytes
tail_bytes dump_remainder
tail_bytes pipe_bytes
file_lines dump_remainder
recheck pretty_name

then you can use tsort to produce an ordering of those functions that satisfies your requirement.

example$ tsort call-graph | tac
dump_remainder
start_lines
file_lines
pipe_lines
xlseek
start_bytes
pipe_bytes
tail_lines
tail_bytes
pretty_name
write_header
tail
recheck
parse_options
tail_file
tail_forever
main

tsort detects any cycles in the input and writes the first cycle encountered to standard error.

Note that for a given partial ordering, generally there is no unique total ordering. In the context of the call graph above, the function parse_options may be placed anywhere in the list as long as it precedes main.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

7.5 tsort: Background

tsort exists because very early versions of the Unix linker processed an archive file exactly once, and in order. As ld read each object in the archive, it decided whether it was needed in the program based on whether it defined any symbols which were undefined at that point in the link.

This meant that dependencies within the archive had to be handled specially. For example, scanf probably calls read. That means that in a single pass through an archive, it was important for scanf.o to appear before read.o, because otherwise a program which calls scanf but not read might end up with an unexpected unresolved reference to read.

The way to address this problem was to first generate a set of dependencies of one object file on another. This was done by a shell script called lorder. The GNU tools don’t provide a version of lorder, as far as I know, but you can still find it in BSD distributions.

Then you ran tsort over the lorder output, and you used the resulting sort to define the order in which you added objects to the archive.

This whole procedure has been obsolete since about 1980, because Unix archives now contain a symbol table (traditionally built by ranlib, now generally built by ar itself), and the Unix linker uses the symbol table to effectively make multiple passes over an archive file.

Anyhow, that’s where tsort came from. To solve an old problem with the way the linker handled archive files, which has since been solved in different ways.

7.6 ptx: Produce permuted indexes

ptx reads a text file and essentially produces a permuted index, with each keyword in its context. The calling sketch is either one of:

     ptx [option ...] [file ...]
     ptx -G [option ...] [input [output]]

The ‘-G’ (or its equivalent: ‘--traditional’) option disables all gnu extensions and reverts to traditional mode, thus introducing some limitations and changing several of the program’s default option values. When ‘-G’ is not specified, gnu extensions are always enabled. gnu extensions to ptx are documented wherever appropriate in this document. For the full list, see See Section 7.6.5 [Compatibility in ptx], page 45.

Individual options are explained in the following sections.

When gnu extensions are enabled, there may be zero, one or several files after the options. If there is no file, the program reads the standard input. If there is one or several files, they give the name of input files which are all read in turn, as if all the input files were concatenated. However, there is a full contextual break between each file and, when automatic referencing is requested, file names and line numbers refer to individual text input files. In all cases, the program outputs the permuted index to the standard output.

When gnu extensions are not enabled, that is, when the program operates in traditional mode, there may be zero, one or two parameters besides the options. If there are no parameters, the program reads the standard input and outputs the permuted index to the standard output. If there is only one parameter, it names the text input to be read instead of the standard input. If two parameters are given, they give respectively the name of the input file to read and the name of the output file to produce. Be very careful to note that, in this case, the contents of file given by the second parameter is destroyed. This behavior is dictated by System V ptx compatibility; gnu Standards normally discourage output parameters not introduced by an option.

Note that for any file named as the value of an option or as an input text file, a single dash - may be used, in which case standard input is assumed. However, it would not make sense to use this convention more than once per program invocation.

7.6.1 General options

‘-C’
‘--copyright’

Print a short note about the copyright and copying conditions, then exit without further processing.

‘-G’
‘--traditional’

As already explained, this option disables all gnu extensions to ptx and switches to traditional mode.

‘--help’ Print a short help on standard output, then exit without further processing.

‘--version’

Print the program version on standard output, then exit without further processing.

7.6.2 Charset selection

As it is set up now, the program assumes that the input file is coded using 8-bit ISO 8859-1 code, also known as Latin-1 character set, unless it is compiled for MS-DOS, in which case it uses the character set of the IBM-PC. (gnu ptx is not known to work on smaller MS-DOS machines anymore.) Compared to 7-bit ascii, the set of characters which are letters is different; this alters the behavior of regular expression matching. Thus, the default regular expression for a keyword allows foreign or diacriticized letters. Keyword sorting, however, is still crude; it obeys the underlying character set ordering quite blindly.

‘-f’
‘--ignore-case’

Fold lower case letters to upper case for sorting.

7.6.3 Word selection and input processing

‘-b file’
‘--break-file=file’

This option provides an alternative (to ‘-W’) method of describing which characters make up words. It introduces the name of a file which contains a list of characters which cannot be part of one word; this file is called the Break file.

Any character which is not part of the Break file is a word constituent. If both options ‘-b’ and ‘-W’ are specified, then ‘-W’ has precedence and ‘-b’ is ignored. When gnu extensions are enabled, the only way to avoid newline as a break character is to write all the break characters in the file with no newline at all, not even at the end of the file. When gnu extensions are disabled, spaces, tabs and newlines are always considered as break characters even if not included in the Break file.

‘-i file’
‘--ignore-file=file’

The file associated with this option contains a list of words which will never be taken as keywords in concordance output. It is called the Ignore file. The file contains exactly one word in each line; the end of line separation of words is not subject to the value of the ‘-S’ option. There is a default Ignore file used by ptx when this option is not specified, usually found in ‘/usr/local/lib/eign’ if this has not been changed at installation time. If you want to deactivate the default Ignore file, specify /dev/null instead.

‘-o file’
‘--only-file=file’

The file associated with this option contains a list of words which will be retained in concordance output; any word not mentioned in this file is ignored. The file is called the Only file. The file contains exactly one word in each line; the end of line separation of words is not subject to the value of the ‘-S’ option. There is no default for the Only file. When both an Only file and an Ignore file are specified, a word is considered a keyword only if it is listed in the Only file and not in the Ignore file.

‘-r’
‘--references’

On each input line, the leading sequence of non-white space characters will be taken to be a reference that has the purpose of identifying this input line in the resulting permuted index. For more information about reference production, see See Section 7.6.4 [Output formatting in ptx], page 42. Using this option changes the default value for option ‘-S’. Using this option, the program does not try very hard to remove references from contexts in output, but it succeeds in doing so when the context ends exactly at the newline. If option ‘-r’ is used with ‘-S’ default value, or when gnu extensions are disabled, this condition is always met and references are completely excluded from the output contexts.

‘-S regexp’
‘--sentence-regexp=regexp’

This option selects which regular expression will describe the end of a line or the end of a sentence. In fact, this regular expression is not the only distinction between end of lines or end of sentences, and input line boundaries have no special significance outside this option. By default, when gnu extensions are enabled and if ‘-r’ option is not used, end of sentences are used. In this case, this regex is imported from gnu Emacs:

     [.?!][]\"')}]*\\($\\|\t\\|  \\)[ \t\n]*

Whenever gnu extensions are disabled or if ‘-r’ option is used, end of lines are used; in this case, the default regexp is just:

     \n

Using an empty regexp is equivalent to completely disabling end of line or end of sentence recognition. In this case, the whole file is considered to be a single big line or sentence. The user might want to disallow all truncation flag generation as well, through option ‘-F ""’. See section “Syntax of Regular Expressions” in The GNU Emacs Manual. When the keywords happen to be near the beginning of the input line or sentence, this often creates an unused area at the beginning of the output context line; when the keywords happen to be near the end of the input line or sentence, this often creates an unused area at the end of the output context line. The program tries to fill those unused areas by wrapping around context in them; the tail of the input line or sentence is used to fill the unused area on the left of the output line; the head of the input line or sentence is used to fill the unused area on the right of the output line. As a matter of convenience to the user, many usual backslashed escape sequences from the C language are recognized and converted to the corresponding characters by ptx itself.

‘-W regexp’
‘--word-regexp=regexp’

This option selects which regular expression will describe each keyword. By default, if gnu extensions are enabled, a word is a sequence of letters; the regexp used is ‘\w+’. When gnu extensions are disabled, a word is by default anything which ends with a space, a tab or a newline; the regexp used is ‘[^

\t\n]+’.

An empty regexp is equivalent to not using this option. See section “Syntax of Regular Expressions” in The GNU Emacs Manual. As a matter of convenience to the user, many usual backslashed escape sequences, as found in the C language, are recognized and converted to the corresponding characters by ptx itself.

7.6.4 Output formatting

Output format is mainly controlled by the ‘-O’ and ‘-T’ options described in the table below. When neither ‘-O’ nor ‘-T’ are selected, and if gnu extensions are enabled, the program chooses an output format suitable for a dumb terminal. Each keyword occurrence is output to the center of one line, surrounded by its left and right contexts. Each field is properly justified, so the concordance output can be readily observed. As a special feature, if automatic references are selected by option ‘-A’ and are output before the left context, that is, if option ‘-R’ is not selected, then a colon is added after the reference; this nicely interfaces with gnu Emacs next-error processing. In this default output format, each white space character, like newline and tab, is merely changed to exactly one space, with no special attempt to compress consecutive spaces. This might change in the future. Except for those white space characters, every other character of the underlying set of 256 characters is transmitted verbatim.

Output format is further controlled by the following options.

‘-g number’
‘--gap-size=number’

Select the size of the minimum white space gap between the fields on the output line.

‘-w number’
‘--width=number’

Select the maximum output width of each final line. If references are used, they are included or excluded from the maximum output width depending on the value of option ‘-R’. If this option is not selected, that is, when references are output before the left context, the maximum output width takes into account the maximum length of all references. If this option is selected, that is, when references are output after the right context, the maximum output width does not take into account the space taken by references, nor the gap that precedes them.

‘-A’
‘--auto-reference’

Select automatic references. Each input line will have an automatic reference made up of the file name and the line ordinal, with a single colon between them. However, the file name will be empty when standard input is being read. If both ‘-A’ and ‘-r’ are selected, then the input reference is still read and skipped, but the automatic reference is used at output time, overriding the input reference.

‘-R’
‘--right-side-refs’

In the default output format, when option ‘-R’ is not used, any references produced by the effect of options ‘-r’ or ‘-A’ are placed to the far right of output lines, after the right context. With default output format, when the ‘-R’ option is specified, references are rather placed at the beginning of each output line, before the left context. For any other output format, option ‘-R’ is ignored, with one exception: with ‘-R’ the width of references is not taken into account in total output width given by ‘-w’. This option is automatically selected whenever gnu extensions are disabled.

‘-F string’
‘--flac-truncation=string’

This option will request that any truncation in the output be reported using the string string. Most output fields theoretically extend towards the beginning or the end of the current line, or current sentence, as selected with option ‘-S’. But there is a maximum allowed output line width, changeable through option ‘-w’, which is further divided into space for various output fields. When a field has to be truncated because it cannot extend beyond the beginning or the end of the current line to fit in, then a truncation occurs. By default, the string used is a single slash, as in ‘-F /’. string may have more than one character, as in ‘-F ...’. Also, in the particular case when string is empty (‘-F ""’), truncation flagging is disabled, and no truncation marks are appended in this case. As a matter of convenience to the user, many usual backslashed escape sequences, as found in the C language, are recognized and converted to the corresponding characters by ptx itself.

‘-M string’
‘--macro-name=string’

Select another string to be used instead of ‘xx’, while generating output suitable

for nroff, troff or TEX.
‘-O’
‘--format=roff’

Choose an output format suitable for nroff or troff processing. Each output line will look like:

     .xx "tail" "before" "keyword_and_after" "head" "ref"

so it will be possible to write a ‘.xx’ roff macro to take care of the output typesetting. This is the default output format when gnu extensions are disabled. Option ‘-M’ can be used to change ‘xx’ to another macro name. In this output format, each non-graphical character, like newline and tab, is merely changed to exactly one space, with no special attempt to compress consecutive spaces. Each quote character: " is doubled so it will be correctly processed by nroff or troff.

‘-T’
‘--format=tex’

Choose an output format suitable for TEX processing. Each output line will look like:

     \xx {tail}{before}{keyword}{after}{head}{ref}

so it will be possible to write a \xx definition to take care of the output typesetting. Note that when references are not being produced, that is, neither option ‘-A’ nor option ‘-r’ is selected, the last parameter of each \xx call is inhibited. Option ‘-M’ can be used to change ‘xx’ to another macro name. In this output format, some special characters, like $, %, &, # and _are automatically protected with a backslash. Curly brackets {, } are protected with a backslash and a pair of dollar signs (to force mathematical mode). The backslash itself produces the sequence \backslash{}. Circumflex and tilde diacritical marks produce the sequence ^\{ } and ~\{ } respectively. Other diacriticized characters of the underlying character set produce an appropriate TEX sequence as far as possible. The other non-graphical characters, like newline and tab, and all other characters which are not part of ascii, are merely changed to exactly one space, with no special attempt to compress consecutive spaces. Let me know how to improve this special character processing for TEX.

7.6.5 The gnu extensions to ptx

This version of ptx contains a few features which do not exist in System V ptx. These extra features are suppressed by using the ‘-G’ command line option, unless overridden by other command line options. Some gnu extensions cannot be recovered by overriding, so the simple rule is to avoid ‘-G’ if you care about gnu extensions. Here are the differences between this program and System V ptx.

• This program can read many input files at once, it always writes the resulting concordance on standard output. On the other hand, System V ptx reads only one file and sends the result to standard output or, if a second file parameter is given on the command, to that file. Having output parameters not introduced by options is a dangerous practice which gnu avoids as far as possible. So, for using ptx portably between gnu and System V, you should always use it with a single input file, and always expect the result on standard output. You might also want to automatically configure in a ‘-G’ option to ptx calls in products using ptx, if the configurator finds that the installed ptx accepts ‘-G’.

• The only options available in System V ptx are options ‘-b’, ‘-f’, ‘-g’, ‘-i’, ‘-o’, ‘-r’, ‘-t’ and ‘-w’. All other options are gnu extensions and are not repeated in this enumeration. Moreover, some options have a slightly different meaning when gnu extensions are enabled, as explained below.

• By default, concordance output is not formatted for troff or nroff. It is rather formatted for a dumb terminal. troff or nroff output may still be selected through option ‘-O’.

• Unless ‘-R’ option is used, the maximum reference width is subtracted from the total output line width. With gnu extensions disabled, width of references is not taken into account in the output line width computations.

• All 256 characters, even NULs, are always read and processed from input file with no adverse effect, even if gnu extensions are disabled. However, System V ptx does not accept 8-bit characters, a few control characters are rejected, and the tilde ~ is also rejected.

• Input line length is only limited by available memory, even if gnu extensions are disabled. However, System V ptx processes only the first 200 characters in each line.

• The break (non-word) characters default to be every character except all letters of the underlying character set, diacriticized or not. When gnu extensions are disabled, the break characters default to space, tab and newline only.

• The program makes better use of output line width. If gnu extensions are disabled, the program rather tries to imitate System V ptx, but still, there are some slight disposition glitches this program does not completely reproduce.

• The user can specify both an Ignore file and an Only file. This is not allowed with System V ptx.

8 Operating on fields within a line

8.1 cut: Print selected parts of lines

cut writes to standard output selected parts of each line of each input file, or standard input if no files are given or for a file name of ‘-’. Synopsis:

     cut [option]... [file]...

In the table which follows, the byte-list, character-list, and field-list are one or more numbers or ranges (two numbers separated by a dash) separated by commas. Bytes, characters, and fields are numbered starting at 1. Incomplete ranges may be given: ‘-m’ means ‘1-m’; ‘n-’ means ‘n’ through end of line or last field. The list elements can be repeated, can overlap, and can be specified in any order; but the selected input is written in the same order that it is read, and is written exactly once.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b byte-list’
‘--bytes=byte-list’

Print only the bytes in positions listed in byte-list. Tabs and backspaces are treated like any other character; they take up 1 byte.

‘-c character-list’
‘--characters=character-list’

Print only characters in positions listed in character-list. The same as ‘-b’ for now, but internationalization will change that. Tabs and backspaces are treated like any other character; they take up 1 character.

‘-f field-list’
‘--fields=field-list’

Print only the fields listed in field-list. Fields are separated by a TAB character by default. Also print any line that contains no delimiter character, unless the ‘--only-delimited’ (‘-s’) option is specified

‘-d input_delim_byte’
‘--delimiter=input_delim_byte’

For ‘-f’, fields are separated in the input by the first character in input delim byte (default is TAB).

‘-n’ Do not split multi-byte characters (no-op for now).

‘-s’
‘--only-delimited’

For ‘-f’, do not print lines that do not contain the field separator character. Normally, any line without a field separator is printed verbatim.

‘--output-delimiter=output_delim_string’

For ‘-f’, output fields are separated by output delim string. The default is to use the input delimiter.

8.2 paste: Merge lines of files

paste writes to standard output lines consisting of sequentially corresponding lines of each given file, separated by a TAB character. Standard input is used for a file name of ‘-’ or if no input files are given.

For example:

$ cat num2
1
2
$ cat let3
a
b
c
$ paste num2 let3
1       a
2       b
        c

Synopsis:

paste [option]... [file]...

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-s’
‘--serial’

Paste the lines of one file at a time rather than one line from each file. Using the above example data:

          $ paste -s num2 let3
          1       2
          a       b       c
‘-d delim-list’
‘--delimiters=delim-list’

Consecutively use the characters in delim-list instead of TAB to separate merged lines. When delim-list is exhausted, start again at its beginning. Using the above example data:

          $ paste -d '%_' num2 let3 num2
          1%a_1
          2%b_2
          %c_

8.3 join: Join lines on a common field

join writes to standard output a line for each pair of input lines that have identical join fields. Synopsis:

join [option]... file1 file2

Either file1 or file2 (but not both) can be ‘-’, meaning standard input. file1 and file2 should be sorted on the join fields.

Normally, the sort order is that of the collating sequence specified by the LC_COLLATE locale. Unless the ‘-t’ option is given, the sort comparison ignores blanks at the start of the join field, as in sort -b. If the ‘--ignore-case’ option is given, the sort comparison ignores the case of characters in the join field, as in sort -f.

However, as a GNU extension, if the input has no unpairable lines the sort order can be any order that considers two fields to be equal if and only if the sort comparison described above considers them to be equal. For example:

     $ cat file1
     a a1
     c c1
     b b1
     $ cat file2
     a a2
     c c2
     b b2
     $ join file1 file2
     a a1 a2
     c c1 c2
     b b1 b2

The defaults are: the join field is the first field in each line; fields in the input are separated by one or more blanks, with leading blanks on the line ignored; fields in the output are separated by a space; each output line consists of the join field, the remaining fields from file1, then the remaining fields from file2.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a file-number’

Print a line for each unpairable line in file file-number (either ‘1’ or ‘2’), in addition to the normal output.

‘-e string’

Replace those output fields that are missing in the input with string.

‘-i’
‘--ignore-case’

Ignore differences in case when comparing keys. With this option, the lines of the input files must be ordered in the same way. Use ‘sort -f’ to produce this ordering.

‘-1 field’
‘-j1 field’

Join on field field (a positive integer) of file 1.

‘-2 field’
‘-j2 field’

Join on field field (a positive integer) of file 2.

‘-j field’

Equivalent to ‘-1 field -2 field’.

‘-o field-list...’

Construct each output line according to the format in field-list. Each element in field-list is either the single character ‘0’ or has the form m.n where the file number, m, is ‘1’ or ‘2’ and n is a positive field number. A field specification of ‘0’ denotes the join field. In most cases, the functionality of the ‘0’ field spec may be reproduced using the explicit m.n that corresponds to the join field. However, when printing unpairable lines (using either of the ‘-a’ or ‘-v’ options), there is no way to specify the join field using m.n in fieldlist if there are unpairable lines in both files. To give join that functionality, posix invented the ‘0’ field specification notation. The elements in field-list are separated by commas or blanks. Multiple field-list arguments can be given after a single ‘-o’ option; the values of all lists given with ‘-o’ are concatenated together. All output lines – including those printed because of any -a or -v option – are subject to the specified field-list.

‘-t char’ Use character char as the input and output field separator.

‘-v file-number’

Print a line for each unpairable line in file file-number (either ‘1’ or ‘2’), instead of the normal output.

In addition, when gnu join is invoked with exactly one argument, the ‘--help’ and ‘--version’ options are recognized. See Chapter 2 [Common options], page 2.

9 Operating on characters

This commands operate on individual characters.

9.1 tr: Translate, squeeze, and/or delete characters

Synopsis:

tr [option]... set1 [set2]

tr copies standard input to standard output, performing one of the following operations:

• translate, and optionally squeeze repeated characters in the result,

• squeeze repeated characters,

• delete characters,

• delete characters, then squeeze repeated characters from the result. The set1 and (if given) set2 arguments define ordered sets of characters, referred to

below as set1 and set2. These sets are the characters of the input that tr operates on. The ‘--complement’ (‘-c’) option replaces set1 with its complement (all of the characters that are not in set1).

9.1.1 Specifying sets of characters

The format of the set1 and set2 arguments resembles the format of regular expressions; however, they are not regular expressions, only lists of characters. Most characters simply represent themselves in these strings, but the strings can contain the shorthands listed below, for convenience. Some of them can be used only in set1 or set2, as noted below. Backslash escapes

A backslash followed by a character not listed below causes an error message. ‘\a’ Control-G. ‘\b’ Control-H. ‘\f’ Control-L. ‘\n’ Control-J. ‘\r’ Control-M. ‘\t’ Control-I. ‘\v’ Control-K. ‘\ooo’ The character with the value given by ooo, which is 1 to 3 octal

digits,

‘\\’ A backslash.

Ranges The notation ‘m-n’ expands to all of the characters from m through n, in ascending order. m should collate before n; if it doesn’t, an error results. As an example, ‘0-9’ is the same as ‘0123456789’. gnu tr does not support the System V syntax that uses square brackets to enclose ranges. Translations specified in that format sometimes work as expected, since the brackets are often transliterated to themselves. However, they should be avoided because they sometimes behave unexpectedly. For example, ‘tr -d '[0-9]'’ deletes brackets as well as digits. Many historically common and even accepted uses of ranges are not portable. For example, on ebcdic hosts using the ‘A-Z’ range will not do what most would expect because ‘A’ through ‘Z’ are not contiguous as they are in ascii. If you can rely on a posix compliant version of tr, then the best way to work around this is to use character classes (see below). Otherwise, it is most portable (and most ugly) to enumerate the members of the ranges.

Repeated characters The notation ‘[c*n]’ in set2 expands to n copies of character c. Thus, ‘[y*6]’ is the same as ‘yyyyyy’. The notation ‘[c*]’ in string2 expands to as many copies of c as are needed to make set2 as long as set1. If n begins with ‘0’, it is interpreted in octal, otherwise in decimal.

Character classes The notation ‘[:class:]’ expands to all of the characters in the (predefined) class class. The characters expand in no particular order, except for the upper and lower classes, which expand in ascending order. When the ‘--delete’ (‘-d’) and ‘--squeeze-repeats’ (‘-s’) options are both given, any character class can be used in set2. Otherwise, only the character classes lower and upper are accepted in set2, and then only if the corresponding character class (upper and lower, respectively) is specified in the same relative position in set1. Doing this specifies case conversion. The class names are given below; an error results when an invalid class name is given.

alnum Letters and digits.

alpha Letters.

blank Horizontal whitespace.

cntrl Control characters.

digit Digits.

graph Printable characters, not including space.

lower Lowercase letters.

print Printable characters, including space.

punct Punctuation characters.

space Horizontal or vertical whitespace.

upper Uppercase letters.

xdigit Hexadecimal digits.

Equivalence classes The syntax ‘[=c=]’ expands to all of the characters that are equivalent to c, in no particular order. Equivalence classes are a relatively recent invention intended to support non-English alphabets. But there seems to be no standard way to define them or determine their contents. Therefore, they are not fully implemented in gnu tr; each character’s equivalence class consists only of that character, which is of no particular use.

9.1.2 Translating

tr performs translation when set1 and set2 are both given and the ‘--delete’ (‘-d’) option is not given. tr translates each character of its input that is in set1 to the corresponding character in set2. Characters not in set1 are passed through unchanged. When a character appears more than once in set1 and the corresponding characters in set2 are not all the same, only the final one is used. For example, these two commands are equivalent:

     tr aaa xyz
     tr a z

A common use of tr is to convert lowercase characters to uppercase. This can be done in many ways. Here are three of them:

     tr abcdefghijklmnopqrstuvwxyz ABCDEFGHIJKLMNOPQRSTUVWXYZ
     tr a-z A-Z
     tr '[:lower:]' '[:upper:]'

But note that using ranges like a-z above is not portable.

When tr is performing translation, set1 and set2 typically have the same length. If set1 is shorter than set2, the extra characters at the end of set2 are ignored.

On the other hand, making set1 longer than set2 is not portable; posix says that the result is undefined. In this situation, BSD tr pads set2 to the length of set1 by repeating the last character of set2 as many times as necessary. System V tr truncates set1 to the length of set2.

By default, gnu tr handles this case like BSD tr. When the ‘--truncate-set1’ (‘-t’) option is given, gnu tr handles this case like the System V tr instead. This option is ignored for operations other than translation.

Acting like System V tr in this case breaks the relatively common BSD idiom:

     tr -cs A-Za-z0-9 '\012'

because it converts only zero bytes (the first element in the complement of set1), rather than all non-alphanumerics, to newlines. By the way, the above idiom is not portable because it uses ranges. Assuming a posix compliant tr, here is a better way to write it:

     tr -cs '[:alnum:]' '[\n*]'

9.1.3 Squeezing repeats and deleting

When given just the ‘--delete’ (‘-d’) option, tr removes any input characters that are in set1.

When given just the ‘--squeeze-repeats’ (‘-s’) option, tr replaces each input sequence of a repeated character that is in set1 with a single occurrence of that character.

When given both ‘--delete’ and ‘--squeeze-repeats’, tr first performs any deletions using set1, then squeezes repeats from any remaining characters using set2.

The ‘--squeeze-repeats’ option may also be used when translating, in which case tr first performs translation, then squeezes repeats from any remaining characters using set2.

Here are some examples to illustrate various combinations of options:

• Remove all zero bytes:

     tr -d '\000'

• Put all words on lines by themselves. This converts all non-alphanumeric characters to newlines, then squeezes each string of repeated newlines into a single newline:

     tr -cs '[:alnum:]' '[\n*]'

• Convert each sequence of repeated newlines to a single newline:

     tr -s '\n'

• Find doubled occurrences of words in a document. For example, people often write “the the” with the duplicated words separated by a newline. The bourne shell script below works first by converting each sequence of punctuation and blank characters to a single newline. That puts each “word” on a line by itself. Next it maps all uppercase characters to lower case, and finally it runs uniq with the ‘-d’ option to print out only the words that were adjacent duplicates.

     #!/bin/sh
     cat "$@" \
       | tr -s '[:punct:][:blank:]' '\n' \
       | tr '[:upper:]' '[:lower:]' \
       | uniq -d

• Deleting a small set of characters is usually straightforward. For example, to remove all ‘a’s, ‘x’s, and ‘M’s you would do this:

     tr -d axM

However, when ‘-’ is one of those characters, it can be tricky because ‘-’ has special meanings. Performing the same task as above but also removing all ‘-’ characters, we might try tr -d -axM, but that would fail because tr would try to interpret ‘-a’ as a command-line option. Alternatively, we could try putting the hyphen inside the string, tr -d a-xM, but that wouldn’t work either because it would make tr interpret a-x as the range of characters ‘a’…‘x’ rather than the three. One way to solve the problem is to put the hyphen at the end of the list of characters:

     tr -d axM-

More generally, use the character class notation [=c=] with ‘-’ (or any other character) in place of the ‘c’:

     tr -d '[=-=]axM'

Note how single quotes are used in the above example to protect the square brackets from interpretation by a shell.

9.1.4 Warning messages

Setting the environment variable POSIXLY_CORRECT turns off the following warning and error messages, for strict compliance with posix. Otherwise, the following diagnostics are issued:

1. When the ‘--delete’ option is given but ‘--squeeze-repeats’ is not, and set2 is given, gnu tr by default prints a usage message and exits, because set2 would not be used. The posix specification says that set2 must be ignored in this case. Silently ignoring arguments is a bad idea.

2. When an ambiguous octal escape is given. For example, ‘\400’ is actually ‘\40’ followed by the digit ‘0’, because the value 400 octal does not fit into a single byte.

gnu tr does not provide complete BSD or System V compatibility. For example, it is impossible to disable interpretation of the posix constructs ‘[:alpha:]’, ‘[=c=]’, and ‘[c*10]’. Also, gnu tr does not delete zero bytes automatically, unlike traditional Unix versions, which provide no way to preserve zero bytes.

9.2 expand: Convert tabs to spaces

expand writes the contents of each given file, or standard input if none are given or for a file of ‘-’, to standard output, with tab characters converted to the appropriate number of spaces. Synopsis:

     expand [option]... [file]...

By default, expand converts all tabs to spaces. It preserves backspace characters in the output; they decrement the column count for tab calculations. The default action is equivalent to ‘-t 8’ (set tabs every 8 columns).

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-t tab1[,tab2]...’
‘--tabs=tab1[,tab2]...’

If only one tab stop is given, set the tabs tab1 spaces apart (default is 8). Otherwise, set the tabs at columns tab1, tab2, … (numbered from 0), and replace any tabs beyond the last tabstop given with single spaces. Tabstops can be separated by blanks as well as by commas.

On older systems, expand supports an obsolete option ‘-tab1[,tab2]...’, where tabstops must be separated by commas. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-t

          tab1[,tab2]...’ instead.
‘-i’
‘--initial’

Only convert initial tabs (those that precede all non-space or non-tab characters) on each line to spaces.

9.3 unexpand: Convert spaces to tabs

unexpand writes the contents of each given file, or standard input if none are given or for a file of ‘-’, to standard output, with strings of two or more space or tab characters converted to as many tabs as possible followed by as many spaces as are needed. Synopsis:

     unexpand [option]... [file]...

By default, unexpand converts only initial spaces and tabs (those that precede all non space or tab characters) on each line. It preserves backspace characters in the output; they decrement the column count for tab calculations. By default, tabs are set at every 8th column.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-t tab1[,tab2]...’
‘--tabs=tab1[,tab2]...’

If only one tab stop is given, set the tabs tab1 spaces apart instead of the default 8. Otherwise, set the tabs at columns tab1, tab2, … (numbered from 0), and leave spaces and tabs beyond the tabstops given unchanged. Tabstops can be separated by blanks as well as by commas. This option implies the ‘-a’ option. On older systems, unexpand supports an obsolete option ‘-tab1[,tab2]...’, where tabstops must be separated by commas. (Unlike ‘-t’, this obsolete option does not imply ‘-a’.) posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘--first-only -t tab1[,tab2]...’ instead.

‘-a’

‘--all’ Convert all strings of two or more spaces or tabs, not just initial ones, to tabs.

10 Directory listing

This chapter describes the ls command and its variants dir and vdir, which list information about files.

10.1 ls: List directory contents

The ls program lists information about files (of any type, including directories). Options and file arguments can be intermixed arbitrarily, as usual.

For non-option command-line arguments that are directories, by default ls lists the contents of directories, not recursively, and omitting files with names beginning with ‘.’. For other non-option arguments, by default ls lists just the file name. If no non-option argument is specified, ls operates on the current directory, acting as if it had been invoked with a single argument of ‘.’.

By default, the output is sorted alphabetically, according to the locale settings in effect.^{1} If standard output is a terminal, the output is in columns (sorted vertically) and control characters are output as question marks; otherwise, the output is listed one per line and control characters are output as-is.

Because ls is such a fundamental program, it has accumulated many options over the years. They are described in the subsections below; within each section, options are listed alphabetically (ignoring case). The division of options into the subsections is not absolute, since some options affect more than one aspect of ls’s operation.

Also see Chapter 2 [Common options], page 2.

10.1.1 Which files are listed

These options determine which files ls lists information for. By default, any files and the contents of any directories on the command line are shown.

‘-a’

‘--all’ List all files in directories, including files that start with ‘.’.

‘-A’
‘--almost-all’

List all files in directories except for ‘.’ and ‘..’.

‘-B’
‘--ignore-backups’

Do not list files that end with ‘~’, unless they are given on the command line.

‘-d’
‘--directory’

List just the names of directories, as with other types of files, rather than listing their contents. Do not follow symbolic links listed on the command line unless

          the ‘--dereference-command-line’ (‘-H’) or ‘--dereference’ (‘-L’) options

are specified.

1 If you use a non-posix locale (e.g., by setting LC_ALL to ‘en_US’), then ls may produce output that is

sorted differently than you’re accustomed to. In that case, set the LC_ALL environment variable to ‘C’.

‘-H’
‘--dereference-command-line’

If a command line argument specifies a symbolic link, show information for the file the link references rather than for the link itself.

‘-I PATTERN’
‘--ignore=PATTERN’

Do not list files whose names match the shell pattern (not regular expression) pattern unless they are given on the command line. As in the shell, an initial ‘.’ in a file name does not match a wildcard at the start of pattern. Sometimes it is useful to give this option several times. For example,

     $ ls --ignore='.??*' --ignore='.[^.]' --ignore='#*'

The first option ignores names of length 3 or more that start with ‘.’, the second ignores all two-character names that start with ‘.’ except ‘..’, and the third ignores names that start with ‘#’.

‘-L’
‘--dereference’

When showing file information for a symbolic link, show information for the file the link references rather than the link itself.

‘-R’
‘--recursive’

List the contents of all directories recursively.

10.1.2 What information is listed

These options affect the information that ls displays. By default, only file names are shown.

‘--author’

List each file’s author when producing long format directory listings. In GNU/Hurd, file authors can differ from their owners, but in other operating systems the two are the same.

‘-D’

‘--dired’ With the long listing (‘-l’) format, print an additional line after the main

output:

     //DIRED// beg1 end1 beg2 end2 ...

The begN and endN are unsigned integers that record the byte position of the beginning and end of each file name in the output. This makes it easy for Emacs to find the names, even when they contain unusual characters such as space or newline, without fancy searching. If directories are being listed recursively (-R), output a similar line with offsets for each subdirectory name:

     //SUBDIRED// beg1 end1 ...

Finally, output a line of the form:

//DIRED-OPTIONS// --quoting-style=word

where word is the quoting style (see Section 10.1.7 [Formatting the file names], page 65). Here is an actual example:

$ mkdir -p a/sub/deeper a/sub2
$ touch a/f1 a/f2
$ touch a/sub/deeper/file
$ ls -gloRF --dired a
a:
total 8
-rw-r--r--    1        0 Nov  9 18:30 f1
-rw-r--r--    1        0 Nov  9 18:30 f2
drwxr-xr-x    3     4096 Nov  9 18:30 sub/
drwxr-xr-x    2     4096 Nov  9 18:30 sub2/
a/sub:
total 4
drwxr-xr-x    2     4096 Nov  9 18:30 deeper/
a/sub/deeper:
total 0
-rw-r--r--    1        0 Nov  9 18:30 file
a/sub2:
total 0
//DIRED// 55 57 98 100 141 144 186 190 252 258 327 331
//SUBDIRED// 2 3 195 200 263 275 335 341
//DIRED-OPTIONS// --quoting-style=literal

Note that the pairs of offsets on the ‘//DIRED//’ line above delimit these names: ‘f1’, ‘f2’, ‘sub’, ‘sub2’, ‘deeper’, ‘file’. The offsets on the ‘//SUBDIRED//’ line delimit the following directory names: ‘a’, ‘a/sub’, ‘a/sub/deeper’, ‘a/sub2’. Here is an example of how to extract the fifth entry name, ‘deeper’, corresponding to the pair of offsets, 252 and 258:

$ ls -gloRF --dired a > out
$ dd bs=1 skip=252 count=6 < out 2>/dev/null; echo
deeper

Note that although the listing above includes a trailing slash for the ‘deeper’ entry, the offsets select the name without the trailing slash. However, if you invoke ls with ‘--dired’ along with an option like ‘--escape’ (aka ‘-b’) and operate on a file whose name contains special characters, notice that the backslash is included:

$ touch 'a b'
$ ls -blog --dired 'a b'
-rw-r--r--    1        0 Nov  9 18:41 a\ b
//DIRED// 40 44

     //DIRED-OPTIONS// --quoting-style=escape

If you use a quoting style that adds quote marks (e.g., ‘--quoting-style=c’), then the offsets include the quote marks. So beware that the user may select the quoting style via the environment variable QUOTING_STYLE. Hence, applications using ‘--dired’ should either specify an explicit ‘--quoting-style=literal’ option (aka ‘-N’ or ‘--literal’) on the command line, or else be prepared to parse the escaped names.

‘--full-time’

Produce long format directory listings, and list times in full. It is equivalent

to using ‘--format=long’ with ‘--time-style=full-iso’ (see Section 10.1.6

[Formatting file timestamps], page 64).

‘-g’ Produce long format directory listings, but don’t display owner information.

‘-G’
‘--no-group’

Inhibit display of group information in a long format directory listing. (This is the default in some non-gnu versions of ls, so we provide this option for compatibility.)

‘-h’
‘--human-readable’

Append a size letter to each size, such as ‘M’ for mebibytes. Powers of 1024 are used, not 1000; ‘M’ stands for 1,048,576 bytes. This option is equivalent to ‘--block-size=human’ (see Section 2.2 [Block size], page 3). Use the ‘--si’ option if you prefer powers of 1000.

‘-i’

‘--inode’ Print the inode number (also called the file serial number and index number) of each file to the left of the file name. (This number uniquely identifies each file within a particular filesystem.)

‘-l’
‘--format=long’
‘--format=verbose’

In addition to the name of each file, print the file type, permissions, number of hard links, owner name, group name, size, and timestamp (see Section 10.1.6 [Formatting file timestamps], page 64), normally the modification time.

Normally the size is printed as a byte count without punctuation, but this can be overridden (see Section 2.2 [Block size], page 3). For example, ‘-h’ prints an abbreviated, human-readable count, and ‘--block-size="'1"’ prints a byte count with the thousands separator of the current locale.

For each directory that is listed, preface the files with a line ‘total blocks’, where blocks is the total disk allocation for all files in that directory. The block size currently defaults to 1024 bytes, but this can be overridden (see Section 2.2 [Block size], page 3). The blocks computed counts each hard link separately; this is arguably a deficiency.

The permissions listed are similar to symbolic mode specifications (see Section 26.2 [Symbolic Modes], page 135). But ls combines multiple bits into the third character of each set of permissions as follows:

‘s’ If the setuid or setgid bit and the corresponding executable bit are both set.

‘S’ If the setuid or setgid bit is set but the corresponding executable bit is not set.

‘t’ If the sticky bit and the other-executable bit are both set.

‘T’ If the sticky bit is set but the other-executable bit is not set.

‘x’ If the executable bit is set and none of the above apply.

‘-’ Otherwise.

Following the permission bits is a single character that specifies whether an alternate access method applies to the file. When that character is a space, there is no alternate access method. When it is a printing character (e.g., ‘+’), then there is such a method.

‘-n’
‘--numeric-uid-gid’

Produce long format directory listings, but display numeric UIDs and GIDs instead of the owner and group names.

‘-o’ Produce long format directory listings, but don’t display group information. It is equivalent to using ‘--format=long’ with ‘--no-group’ .

‘-s’

‘--size’ Print the disk allocation of each file to the left of the file name. This is the amount of disk space used by the file, which is usually a bit more than the file’s size, but it can be less if the file has holes. Normally the disk allocation is printed in units of 1024 bytes, but this can be overridden (see Section 2.2 [Block size], page 3). For files that are NFS-mounted from an HP-UX system to a BSD system, this option reports sizes that are half the correct values. On HP-UX systems, it reports sizes that are twice the correct values for files that are NFS-mounted from BSD systems. This is due to a flaw in HP-UX; it also affects the HP-UX ls program.

‘--si’ Append an SI-style abbreviation to each size, such as ‘MB’ for megabytes. Powers of 1000 are used, not 1024; ‘MB’ stands for 1,000,000 bytes. This option is

equivalent to ‘--block-size=si’. Use the ‘-h’ or ‘--human-readable’ option

if you prefer powers of 1024.

10.1.3 Sorting the output

These options change the order in which ls sorts the information it outputs. By default, sorting is done by character code (e.g., ascii order).

‘-c’
‘--time=ctime’
‘--time=status’
‘--time=use’

If the long listing format (e.g., ‘-l’, ‘-o’) is being used, print the status change time (the ‘ctime’ in the inode) instead of the modification time. When explicitly sorting by time (‘--sort=time’ or ‘-t’) or when not using a long listing format, sort according to the status change time.

‘-f’ Primarily, like ‘-U’—do not sort; list the files in whatever order they are stored in the directory. But also enable ‘-a’ (list all files) and disable ‘-l’, ‘--color’, and ‘-s’ (if they were specified before the ‘-f’).

‘-r’
‘--reverse’

Reverse whatever the sorting method is—e.g., list files in reverse alphabetical order, youngest first, smallest first, or whatever.

‘-S’
‘--sort=size’

Sort by file size, largest first.

‘-t’
‘--sort=time’

Sort by modification time (the ‘mtime’ in the inode), newest first.

‘-u’
‘--time=atime’
‘--time=access’

If the long listing format (e.g., ‘--format=long’) is being used, print the last access time (the ‘atime’ in the inode). When explicitly sorting by time (‘--sort=time’ or ‘-t’) or when not using a long listing format, sort according to the access time.

‘-U’
‘--sort=none’

Do not sort; list the files in whatever order they are stored in the directory. (Do not do any of the other unrelated things that ‘-f’ does.) This is especially useful when listing very large directories, since not doing any sorting can be noticeably faster.

‘-v’
‘--sort=version’

Sort by version name and number, lowest first. It behaves like a default sort, except that each sequence of decimal digits is treated numerically as an index/version number. (See Section 10.1.4 [More details about version sort], page 62.)

‘-X’
‘--sort=extension’

Sort directory contents alphabetically by file extension (characters after the last ‘.’); files with no extension are sorted first.

10.1.4 More details about version sort

The version sort takes into account the fact that file names frequently include indices or version numbers. Standard sorting functions usually do not produce the ordering that people expect because comparisons are made on a character-by-character basis. The version sort addresses this problem, and is especially useful when browsing directories that contain many files with indices/version numbers in their names:

> ls -1            > ls -1v
foo.zml-1.gz       foo.zml-1.gz
foo.zml-100.gz     foo.zml-2.gz
foo.zml-12.gz      foo.zml-6.gz
foo.zml-13.gz      foo.zml-12.gz
foo.zml-2.gz       foo.zml-13.gz
foo.zml-25.gz      foo.zml-25.gz
foo.zml-6.gz       foo.zml-100.gz

Note also that numeric parts with leading zeroes are considered as fractional one:

> ls -1            > ls -1v
abc-1.007.tgz      abc-1.007.tgz
abc-1.012b.tgz     abc-1.01a.tgz
abc-1.01a.tgz      abc-1.012b.tgz

10.1.5 General output formatting

These options affect the appearance of the overall output.

‘-1’
‘--format=single-column’

List one file per line. This is the default for ls when standard output is not a terminal.

‘-C’
‘--format=vertical’

List files in columns, sorted vertically. This is the default for ls if standard output is a terminal. It is always the default for the dir and d programs. gnu ls uses variable width columns to display as many files as possible in the fewest lines.

‘--color [=when]’

Specify whether to use color for distinguishing file types. when may be omitted, or one of:

• none - Do not use color at all. This is the default.

• auto - Only use color if standard output is a terminal.

• always - Always use color.

Specifying ‘--color’ and no when is equivalent to ‘--color=always’. Piping a colorized listing through a pager like more or less usually produces unreadable results. However, using more -f does seem to work.

‘-F’
‘--classify’
‘--indicator-style=classify’

Append a character to each file name indicating the file type. Also, for regular files that are executable, append ‘*’. The file type indicators are ‘/’ for directories, ‘@’ for symbolic links, ‘|’ for FIFOs, ‘=’ for sockets, and nothing for regular files. Do not follow symbolic links listed on the command line unless

the ‘--dereference-command-line’ (‘-H’) or ‘--dereference’ (‘-L’) options

are specified.

‘--indicator-style=word’

Append a character indicator with style word to entry names, as follows:

‘none’ Do not append any character indicator; this is the default.

‘file-type’

Append ‘/’ for directories, ‘@’ for symbolic links, ‘|’ for FIFOs, ‘=’ for sockets, and nothing for regular files. This is the same as the

          ‘-p’ or ‘--file-type’ option.
‘classify’

Append ‘*’ for executable regular files, otherwise behave as for ‘file-type’. This is the same as the ‘-F’ or ‘--classify’ option.

‘-k’ Print file sizes in 1024-byte blocks, overriding the default block size (see Section 2.2 [Block size], page 3). This option is equivalent to ‘--block-size=1K’.

‘-m’
‘--format=commas’

List files horizontally, with as many as will fit on each line, separated by ‘, ’ (a comma and a space).

‘-p’
‘--file-type’
‘--indicator-style=file-type’

Append a character to each file name indicating the file type. This is like ‘-F’, except that executables are not marked.

‘-x format’
‘--format=across’
‘--format=horizontal’

List the files in columns, sorted horizontally.

‘-T cols’
‘--tabsize=cols’

Assume that each tabstop is cols columns wide. The default is 8. ls uses tabs where possible in the output, for efficiency. If cols is zero, do not use tabs at all.

‘-w’
‘--width=cols’

Assume the screen is cols columns wide. The default is taken from the terminal settings if possible; otherwise the environment variable COLUMNS is used if it is set; otherwise the default is 80.

10.1.6 Formatting file timestamps

By default, file timestamps are listed in abbreviated form. Most locales use a timestamp like ‘2002-03-30 23:45’. However, the default posix locale uses a date like ‘Mar 30 2002’ for non-recent timestamps, and a date-without-year and time like ‘Mar 30 23:45’ for recent timestamps.

A timestamp is considered to be recent if it is less than six months old, and is not dated in the future. If a timestamp dated today is not listed in recent form, the timestamp is in the future, which means you probably have clock skew problems which may break programs like make that rely on file timestamps.

The following option changes how file timestamps are printed.

‘--time-style=style’

List timestamps in style style. The style should be one of the following:

‘+format’ List timestamps using format, where format is interpreted like the format argument of date (see Section 21.1 [date invocation],

page 117). For example, ‘--time-style="+%Y-%m-%d %H:%M:S"’

causes ls to list timestamps like ‘2002-03-30 23:45:56’. As with date, format’s interpretation is affected by the LC_TIME locale category. If format contains two format strings separated by a newline, the former is used for non-recent files and the latter for recent files; if you want output columns to line up, you may need to insert spaces in one of the two formats.

‘full-iso’

List timestamps in full using iso 8601 date, time, and time zone format with nanosecond precision, e.g., ‘2002-03-30 23:45:56.477817180 -0700’. This style is equivalent to

‘+%Y-%m-%d %H:%M:%S.%N %z’.

This is useful because the time output includes all the information that is available from the operating system. For example, this can help explain make’s behavior, since gnu make uses the full timestamp to determine whether a file is out of date.

‘long-iso’

List iso 8601 date and time in minutes, e.g., ‘2002-03-30 23:45’. These timestamps are shorter than ‘full-iso’ timestamps, and are usually good enough for everyday work. This style is equivalent to

‘%Y-%m-%d %H:%M’.

‘iso’ List iso 8601 dates for non-recent timestamps (e.g., ‘2002-03-30 ’), and iso 8601 month, day, hour, and minute for recent timestamps (e.g., ‘03-30 23:45’). These timestamps are uglier than ‘long-iso’ timestamps, but they carry nearly the same information in a smaller space and their brevity helps ls output fit within traditional 80-column output lines. The following two ls invocations are equivalent:

     newline='
     '
     ls -l --time-style="+%Y-%m-%d $newline%m-%d %H:%M"
     ls -l --time-style="iso"

‘locale’ List timestamps in a locale-dependent form. For example, a Finnish locale might list non-recent timestamps like ‘maalis 30 2002’ and recent timestamps like ‘maalis 30 23:45’. Locale-dependent timestamps typically consume more space than ‘iso’ timestamps and are harder for programs to parse because locale conventions vary so widely, but they are easier for many people to read. The LC_TIME locale category specifies the timestamp format. The default posix locale uses timestamps like ‘Mar 30 2002’ and ‘Mar 30 23:45’; in this locale, the following two ls invocations are equivalent:

     newline='
     '
     ls -l --time-style="+%b %e  %Y$newline%b %e %H:%M"
     ls -l --time-style="locale"

Other locales behave differently. For example, in a German locale, ‘--time-style="locale"’ might be equivalent to

‘--time-style="+%e. %b %Y $newline%e. %b %H:%M"’ and might

generate timestamps like ‘30. Mär 2002 ’ and ‘30. Mär 23:45’.

‘posix-style’

List posix-locale timestamps if the LC_TIME locale category is posix, style timestamps otherwise. For example, the default style, which is ‘posix-long-iso’, lists timestamps like ‘Mar 30 2002’ and ‘Mar 30 23:45’ when in the posix locale, and like

‘2002-03-30 23:45’ otherwise.

You can specify the default value of the ‘--time-style’ option with the environment variable TIME_STYLE; if TIME_STYLE is not set the default style is ‘posix-long-iso’. gnu Emacs 21 and later can parse iso dates, but older Emacs versions do not, so if you are using an older version of Emacs and specify a non-posix locale, you may need to set

‘TIME_STYLE="locale"’.

10.1.7 Formatting the file names

These options change how file names themselves are printed.

‘-b’
‘--escape’
‘--quoting-style=escape’

Quote nongraphic characters in file names using alphabetic and octal backslash sequences like those used in C.

‘-N’
‘--literal’
‘--quoting-style=literal’

Do not quote file names.

‘-q’
‘--hide-control-chars’

Print question marks instead of nongraphic characters in file names. This is the default if the output is a terminal and the program is ls.

‘-Q’
‘--quote-name’
‘--quoting-style=c’

Enclose file names in double quotes and quote nongraphic characters as in C.

‘--quoting-style=word’

Use style word to quote output names. The word should be one of the following:

‘literal’ Output names as-is; this is the same as the ‘-N’ or ‘--literal’ option.

‘shell’ Quote names for the shell if they contain shell metacharacters or would cause ambiguous output.

‘shell-always’

Quote names for the shell, even if they would normally not require quoting.

‘c’ Quote names as for a C language string; this is the same as the ‘-Q’

          or ‘--quote-name’ option.

‘escape’ Quote as with ‘c’ except omit the surrounding double-quote characters; this is the same as the ‘-b’ or ‘--escape’ option.

‘clocale’ Quote as with ‘c’ except use quotation marks appropriate for the locale.

‘locale’ Like ‘clocale’, but quote `like this' instead of "like this" in

the default C locale. This looks nicer on many displays.

You can specify the default value of the ‘--quoting-style’ option with the environment variable QUOTING_STYLE. If that environment variable is not set, the default value is ‘literal’, but this default may change to ‘shell’ in a future version of this package.

‘--show-control-chars’

Print nongraphic characters as-is in file names. This is the default unless the output is a terminal and the program is ls.

10.2 dir: Briefly list directory contents

dir (also installed as d) is equivalent to ls -C -b; that is, by default files are listed in columns, sorted vertically, and special characters are represented by backslash escape sequences.

See Section 10.1 [ls invocation], page 56.

10.3 vdir: Verbosely list directory contents

vdir (also installed as v) is equivalent to ls -l -b; that is, by default files are listed in long format and special characters are represented by backslash escape sequences.

10.4 dircolors: Color setup for ls

dircolors outputs a sequence of shell commands to set up the terminal for color output from ls (and dir, etc.). Typical usage:

     eval `dircolors [option]... [file]`

If file is specified, dircolors reads it to determine which colors to use for which file types and extensions. Otherwise, a precompiled database is used. For details on the format

of these files, run ‘dircolors --print-database’.

The output is a shell command to set the LS_COLORS environment variable. You can specify the shell syntax to use on the command line, or dircolors will guess it from the value of the SHELL environment variable.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--sh’
‘--bourne-shell’

Output Bourne shell commands. This is the default if the SHELL environment variable is set and does not end with ‘csh’ or ‘tcsh’.

‘-c’
‘--csh’
‘--c-shell’

Output C shell commands. This is the default if SHELL ends with csh or tcsh.

‘-p’
‘--print-database’

Print the (compiled-in) default color configuration database. This output is itself a valid configuration file, and is fairly descriptive of the possibilities.

11 Basic operations

This chapter describes the commands for basic file manipulation: copying, moving (renaming), and deleting (removing).

11.1 cp: Copy files and directories

cp copies files (or, optionally, directories). The copy is completely independent of the original. You can either copy one file to another, or copy arbitrarily many files to a destination directory. Synopsis:

     cp [option]... source dest
     cp [option]... source... directory

If the last argument names an existing directory, cp copies each source file into that directory (retaining the same name). Otherwise, if only two files are given, it copies the first onto the second. It is an error if the last argument is not a directory and more than two non-option arguments are given.

Generally, files are written just as they are read. For exceptions, see the ‘--sparse’ option below.

By default, cp does not copy directories. However, the ‘-R’, ‘-a’, and ‘-r’ options cause cp to copy recursively by descending into source directories and copying files to corresponding destination directories.

By default, cp follows symbolic links only when not copying recursively. This default can be overridden with the ‘--archive’ (‘-a’), ‘-d’, ‘--dereference’ (‘-L’), ‘--no-dereference’ (‘-P’), and ‘-H’ options. If more than one of these options is specified, the last one silently overrides the others.

By default, cp copies the contents of special files only when not copying recursively. This default can be overridden with the ‘--copy-contents’ option.

cp generally refuses to copy a file onto itself, with the following exception: if ‘--force --backup’ is specified with source and dest identical, and referring to a regular file, cp will make a backup file, either regular or numbered, as specified in the usual ways (see Section 2.1 [Backup options], page 2). This is useful when you simply want to make a backup of an existing file before changing it.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’
‘--archive’

Preserve as much as possible of the structure and attributes of the original files in the copy (but do not attempt to preserve internal directory structure; i.e., ‘ls -U’ may list the entries in a copied directory in a different order). Equivalent

          to ‘-dpPR’.
‘-b’
‘--backup[=method]’

See Section 2.1 [Backup options], page 2. Make a backup of each file that would otherwise be overwritten or removed. As a special case, cp makes a backup of source when the force and backup options are given and source and dest are the same name for an existing, regular file. One useful application of this combination of options is this tiny Bourne shell script:

     #!/bin/sh
     # Usage: backup FILE...
     # Create a gnu-style backup of each listed FILE.
     for i; do
       cp --backup --force "$i" "$i"
     done
‘--copy-contents’

If copying recursively, copy the contents of any special files (e.g., FIFOs and device files) as if they were regular files. This means trying to read the data in each source file and writing it to the destination. It is usually a mistake to use this option, as it normally has undesirable effects on special files like FIFOs and the ones typically found in the ‘/dev’ directory. In most cases, cp -R --copy-contents will hang indefinitely trying to read from FIFOs and special files like ‘/dev/console’, and it will fill up your destination disk if you use it to copy ‘/dev/zero’. This option has no effect unless copying recursively, and it does not affect the copying of symbolic links.

‘-d’ Copy symbolic links as symbolic links rather than copying the files that they point to, and preserve hard links between source files in the copies. Equivalent

to ‘--no-dereference --preserve=links’.
‘-f’

‘--force’ When copying without this option and an existing destination file cannot be opened for writing, the copy fails. However, with ‘--force’), when a destination file cannot be opened, cp then unlinks it and tries to open it again. Contrast this behavior with that enabled by ‘--link’ and ‘--symbolic-link’, whereby the destination file is never opened but rather is unlinked unconditionally. Also see the description of ‘--remove-destination’.

‘-H’ If a command line argument specifies a symbolic link, then copy the file it points to rather than the symbolic link itself. However, copy (preserving its nature) any symbolic link that is encountered via recursive traversal.

‘-i’
‘--interactive’

Prompt whether to overwrite existing regular destination files.

‘-l’

‘--link’ Make hard links instead of copies of non-directories.

‘-L’
‘--dereference’

Always follow symbolic links.

‘-P’
‘--no-dereference’

Copy symbolic links as symbolic links rather than copying the files that they point to.

‘-p’
‘--preserve[=attribute_list]’

Preserve the specified attributes of the original files. If specified, the attribute list must be a comma-separated list of one or more of the following strings:

‘mode’ Preserve the permission attributes.

‘ownership’

Preserve the owner and group. On most modern systems, only the super-user may change the owner of a file, and regular users may preserve the group ownership of a file only if they happen to be a member of the desired group.

‘timestamps’

Preserve the times of last access and last modification.

‘links’ Preserve in the destination files any links between corresponding

source files.

‘all’ Preserve all file attributes. Equivalent to specifying all of the above.

Using ‘--preserve’ with no attribute list is equivalent to ‘--preserve=mode,ownership,timestamps’.

In the absence of this option, each destination file is created with the permissions of the corresponding source file, minus the bits set in the umask and minus the set-user-id and set-group-id bits. See Chapter 26 [File permissions], page 135.

‘--no-preserve=attribute_list’

Do not preserve the specified attributes. The attribute list has the same form

as for ‘--preserve’.
‘--parents’

Form the name of each destination file by appending to the target directory a slash and the specified name of the source file. The last argument given to cp must be the name of an existing directory. For example, the command:

     cp --parents a/b/c existing_dir

copies the file ‘a/b/c’ to ‘existing_dir/a/b/c’, creating any missing intermediate directories.

‘--reply[=how]’

Using ‘--reply=yes’ makes cp act as if ‘yes’ were given as a response to every prompt about a destination file. That effectively cancels any preceding ‘--interactive’ or ‘-i’ option. Specify ‘--reply=no’ to make cp act as if ‘no’ were given as a response to every prompt about a destination file. Specify ‘--reply=query’ to make cp prompt the user about each existing destination file.

‘-R’
‘-r’
‘--recursive’

Copy directories recursively. Symbolic links are not followed by default; see the

‘--archive’ (‘-a’), ‘-d’, ‘--dereference’ (‘-L’), ‘--no-dereference’ (‘-P’),

and ‘-H’ options. Special files are copied by creating a destination file of the same type as the source; see the ‘--copy-contents’ option. It is not portable to use ‘-r’ to copy symbolic links or special files. On some non-gnu systems, ‘-r’ implies the equivalent of ‘-L’ and ‘--copy-contents’ for historical reasons. Also, it is not portable to use ‘-R’ to copy symbolic links unless you also specify ‘-P’, as posix allows implementations that dereference symbolic links by default.

‘--remove-destination’

Remove each existing destination file before attempting to open it (contrast with ‘-f’ above).

‘--sparse=when’

A sparse file contains holes—a sequence of zero bytes that does not occupy any physical disk blocks; the ‘read’ system call reads these as zeroes. This can both save considerable disk space and increase speed, since many binary files contain lots of consecutive zero bytes. By default, cp detects holes in input source files via a crude heuristic and makes the corresponding output file sparse as well. The when value can be one of the following:

‘auto’ The default behavior: the output file is sparse if the input file is sparse.

‘always’ Always make the output file sparse. This is useful when the input file resides on a filesystem that does not support sparse files (the most notable example is ‘efs’ filesystems in SGI IRIX 5.3 and earlier), but the output file is on another type of filesystem.

‘never’ Never make the output file sparse. This is useful in creating a file for use with the mkswap command, since such a file must not have any holes.

‘--strip-trailing-slashes’

Remove any trailing slashes from each source argument. See Section 2.4 [Trailing slashes], page 6.

‘-s’
‘--symbolic-link’

Make symbolic links instead of copies of non-directories. All source file names must be absolute (starting with ‘/’) unless the destination files are in the current directory. This option merely results in an error message on systems that do not support symbolic links.

‘-S suffix’
‘--suffix=suffix’

Append suffix to each backup file made with ‘-b’. See Section 2.1 [Backup options], page 2.

‘--target-directory=directory’

Specify the destination directory. See Section 2.3 [Target directory], page 5.

‘-v’
‘--verbose’

Print the name of each file before copying it.

‘-V method’
‘--version-control=method’

Change the type of backups made with ‘-b’. The method argument can be ‘none’ (or ‘off’), ‘numbered’ (or ‘t’), ‘existing’ (or ‘nil’), or ‘never’ (or ‘simple’). See Section 2.1 [Backup options], page 2.

‘-x’
‘--one-file-system’

Skip subdirectories that are on different filesystems from the one that the copy started on. However, mount point directories are copied.

11.2 dd: Convert and copy a file

dd copies a file (from standard input to standard output, by default) with a changeable I/O block size, while optionally performing conversions on it. Synopsis:

     dd [option]...

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

The numeric-valued options below (bytes and blocks) can be followed by a multiplier: ‘b’=512, ‘c’=1, ‘w’=2, ‘xm’=m, or any of the standard block size suffixes like ‘k’=1024 (see Section 2.2 [Block size], page 3).

Use different dd invocations to use different block sizes for skipping and I/O. For example, the following shell commands copy data in 512 KiB blocks between a disk and a tape, but do not save or restore a 4 KiB label at the start of the disk:

     disk=/dev/rdsk/c0t1d0s2
     tape=/dev/rmt/0
     # Copy all but the label from disk to tape.
     (dd bs=4k skip=1 count=0 && dd bs=512k) <$disk >$tape
     # Copy from tape back to disk, but leave the disk label alone.
     (dd bs=4k seek=1 count=0 && dd bs=512k) <$tape >$disk

‘if=file’ Read from file instead of standard input.

‘of=file’ Write to file instead of standard output. Unless ‘conv=notrunc’ is given, dd truncates file to zero bytes (or the size specified with ‘seek=’).

‘ibs=bytes’

Read bytes bytes at a time.

‘obs=bytes’

Write bytes bytes at a time.

‘bs=bytes’

Both read and write bytes bytes at a time. This overrides ‘ibs’ and ‘obs’.

‘cbs=bytes’

Convert bytes bytes at a time.

‘skip=blocks’

Skip blocks ‘ibs’-byte blocks in the input file before copying.

‘seek=blocks’

Skip blocks ‘obs’-byte blocks in the output file before copying.

‘count=blocks’

Copy blocks ‘ibs’-byte blocks from the input file, instead of everything until the end of the file.

‘conv=conversion[,conversion]...’

Convert the file as specified by the conversion argument(s). (No spaces around any comma(s).) Conversions:

‘ascii’ Convert ebcdic to ascii.

‘ebcdic’ Convert ascii to ebcdic.

‘ibm’ Convert ascii to alternate ebcdic.

‘block’ For each line in the input, output ‘cbs’ bytes, replacing the input newline with a space and padding with spaces as necessary.

‘unblock’ Replace trailing spaces in each ‘cbs’-sized input block with a newline.

‘lcase’ Change uppercase letters to lowercase.

‘ucase’ Change lowercase letters to uppercase.

‘swab’ Swap every pair of input bytes. gnu dd, unlike others, works when an odd number of bytes are read—the last byte is simply copied (since there is nothing to swap it with).

‘noerror’ Continue after read errors.

‘notrunc’ Do not truncate the output file.

‘sync’ Pad every input block to size of ‘ibs’ with trailing zero bytes. When used with ‘block’ or ‘unblock’, pad with spaces instead of zero bytes.

11.3 install: Copy files and set attributes

install copies files while setting their permission modes and, if possible, their owner and group. Synopses:

install [option]... source dest
install [option]... source... directory
install -d [option]... directory...

In the first of these, the source file is copied to the dest target file. In the second, each of the source files are copied to the destination directory. In the last, each directory (and any missing parent directories) is created.

install is similar to cp, but allows you to control the attributes of destination files. It is typically used in Makefiles to copy programs into their destination directories. It refuses to copy files onto themselves.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--backup[=method]’

See Section 2.1 [Backup options], page 2. Make a backup of each file that would otherwise be overwritten or removed.

‘-c’ Ignored; for compatibility with old Unix versions of install.

‘-d’
‘--directory’

Create each given directory and any missing parent directories, setting the owner, group and mode as given on the command line or to the defaults. It also gives any parent directories it creates those attributes. (This is different from the SunOS 4.x install, which gives directories that it creates the default attributes.)

‘-g group’
‘--group=group’

Set the group ownership of installed files or directories to group. The default is the process’ current group. group may be either a group name or a numeric group id.

‘-m mode’
‘--mode=mode’

Set the permissions for the installed file or directory to mode, which can be either an octal number, or a symbolic mode as in chmod, with 0 as the point of departure (see Chapter 26 [File permissions], page 135). The default mode is ‘u=rwx,go=rx’—read, write, and execute for the owner, and read and execute for group and other.

‘-o owner’
‘--owner=owner’

If install has appropriate privileges (is run as root), set the ownership of installed files or directories to owner. The default is root. owner may be either a user name or a numeric user ID.

‘-p’
‘--preserve-timestamps’

Set the time of last access and the time of last modification of each installed file to match those of each corresponding original file. When a file is installed without this option, its last access and last modification times are both set to the time of installation. This option is useful if you want to use the last modification times of installed files to keep track of when they were last built as opposed to when they were last installed.

‘-s’

‘--strip’ Strip the symbol tables from installed binary executables.

‘-S suffix’
‘--suffix=suffix’

Append suffix to each backup file made with ‘-b’. See Section 2.1 [Backup options], page 2.

‘--target-directory=directory’

Specify the destination directory. See Section 2.3 [Target directory], page 5.

‘-v’
‘--verbose’

Print the name of each file before copying it.

‘-V method’
‘--version-control=method’

Change the type of backups made with ‘-b’. The method argument can be ‘none’ (or ‘off’), ‘numbered’ (or ‘t’), ‘existing’ (or ‘nil’), or ‘never’ (or ‘simple’). See Section 2.1 [Backup options], page 2.

11.4 mv: Move (rename) files

mv moves or renames files (or directories). Synopsis:

     mv [option]... source dest
     mv [option]... source... directory

If the last argument names an existing directory, mv moves each other given file into a file with the same name in that directory. Otherwise, if only two files are given, it renames the first as the second. It is an error if the last argument is not a directory and more than two files are given.

mv can move any type of file from one filesystem to another. Prior to version 4.0 of the fileutils, mv could move only regular files between filesystems. For example, now mv can move an entire directory hierarchy including special device files from one partition to another. It first uses some of the same code that’s used by cp -a to copy the requested directories and files, then (assuming the copy succeeded) it removes the originals. If the copy fails, then the part that was copied to the destination partition is removed. If you were to copy three directories from one partition to another and the copy of the first directory succeeded, but the second didn’t, the first would be left on the destination partition and the second and third would be left on the original partition.

If a destination file exists but is normally unwritable, standard input is a terminal, and the ‘-f’ or ‘--force’ option is not given, mv prompts the user for whether to replace the file. (You might own the file, or have write permission on its directory.) If the response does not begin with ‘y’ or ‘Y’, the file is skipped.

Warning: If you try to move a symlink that points to a directory, and you specify the symlink with a trailing slash, then mv doesn’t move the symlink but instead moves the directory referenced by the symlink. See Section 2.4 [Trailing slashes], page 6.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--backup[=method]’

See Section 2.1 [Backup options], page 2. Make a backup of each file that would otherwise be overwritten or removed.

‘-f’

‘--force’ Do not prompt the user before removing a destination file.

‘-i’
‘--interactive’

Prompt whether to overwrite each existing destination file, regardless of its permissions. If the response does not begin with ‘y’ or ‘Y’, the file is skipped.

‘--reply[=how]’

Specifying ‘--reply=yes’ is equivalent to using ‘--force’. Specify ‘--reply=no’ to make mv act as if ‘no’ were given as a response to every prompt about a destination file. Specify ‘--reply=query’ to make mv prompt the user about each existing destination file.

‘-u’
‘--update’

Do not move a non-directory that has an existing destination with the same or newer modification time.

‘-v’
‘--verbose’

Print the name of each file before moving it.

‘--strip-trailing-slashes’

Remove any trailing slashes from each source argument. See Section 2.4 [Trailing slashes], page 6.

‘-S suffix’
‘--suffix=suffix’

Append suffix to each backup file made with ‘-b’. See Section 2.1 [Backup options], page 2.

‘--target-directory=directory’

Specify the destination directory. See Section 2.3 [Target directory], page 5.

‘-V method’
‘--version-control=method’

Change the type of backups made with ‘-b’. The method argument can be ‘none’ (or ‘off’), ‘numbered’ (or ‘t’), ‘existing’ (or ‘nil’), or ‘never’ (or ‘simple’). See Section 2.1 [Backup options], page 2.

11.5 rm: Remove files or directories

rm removes each given file. By default, it does not remove directories. Synopsis:

     rm [option]... [file]...

If a file is unwritable, standard input is a terminal, and the ‘-f’ or ‘--force’ option is not given, or the ‘-i’ or ‘--interactive’ option is given, rm prompts the user for whether to remove the file. If the response does not begin with ‘y’ or ‘Y’, the file is skipped.

Warning: If you use rm to remove a file, it is usually possible to recover the contents of that file. If you want more assurance that the contents are truly unrecoverable, consider using shred.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-d’
‘--directory’

Attempt to remove directories using the unlink function rather than the rmdir function, and don’t require a directory to be empty before trying to unlink it. This works only if you have appropriate privileges and if your operating system supports unlink for directories. Because unlinking a directory causes any files in the deleted directory to become unreferenced, it is wise to fsck the filesystem after doing this.

‘-f’

‘--force’ Ignore nonexistent files and never prompt the user. Ignore any previous

          ‘--interactive’ (‘-i’) option.
‘-i’
‘--interactive’

Prompt whether to remove each file. If the response does not begin with ‘y’ or ‘Y’, the file is skipped. Ignore any previous ‘--force’ (‘-f’) option.

‘-r’
‘-R’
‘--recursive’

Remove the contents of directories recursively.

‘-v’
‘--verbose’

Print the name of each file before removing it.

One common question is how to remove files whose names begin with a ‘-’. gnu rm, like every program that uses the getopt function to parse its arguments, lets you use the ‘--’ option to indicate that all following arguments are non-options. To remove a file called ‘-f’ in the current directory, you could type either:

     rm -- -f

or:

     rm ./-f

The Unix rm program’s use of a single ‘-’ for this purpose predates the development of the getopt standard syntax.

11.6 shred: Remove files more securely

shred overwrites devices or files, to help prevent even very expensive hardware from recovering the data.

Ordinarily when you remove a file (see Section 11.5 [rm invocation], page 76), the data is not actually destroyed. Only the index listing where the file is stored is destroyed, and the storage is made available for reuse. There are undelete utilities that will attempt to reconstruct the index and can bring the file back if the parts were not reused.

On a busy system with a nearly-full drive, space can get reused in a few seconds. But there is no way to know for sure. If you have sensitive data, you may want to be sure that recovery is not possible by actually overwriting the file with non-sensitive data.

However, even after doing that, it is possible to take the disk back to a laboratory and use a lot of sensitive (and expensive) equipment to look for the faint “echoes” of the original data underneath the overwritten data. If the data has only been overwritten once, it’s not even that hard.

The best way to remove something irretrievably is to destroy the media it’s on with acid, melt it down, or the like. For cheap removable media like floppy disks, this is the preferred method. However, hard drives are expensive and hard to melt, so the shred utility tries to achieve a similar effect non-destructively.

This uses many overwrite passes, with the data patterns chosen to maximize the damage they do to the old data. While this will work on floppies, the patterns are designed for best effect on hard drives. For more details, see the source code and Peter Gutmann’s paper Secure Deletion of Data from Magnetic and Solid-State Memory, from the proceedings of the Sixth USENIX Security Symposium (San Jose, California, 22–25 July, 1996). The paper is

also available online http://www.cs.auckland.ac.nz/~pgut001/pubs/secure_del.html.

Please note that shred relies on a very important assumption: that the filesystem overwrites data in place. This is the traditional way to do things, but many modern filesystem designs do not satisfy this assumption. Exceptions include:

• Log-structured or journaled filesystems, such as those supplied with AIX and Solaris, and JFS, ReiserFS, XFS, Ext3, etc.

• Filesystems that write redundant data and carry on even if some writes fail, such as RAID-based filesystems.

• Filesystems that make snapshots, such as Network Appliance’s NFS server.

• Filesystems that cache in temporary locations, such as NFS version 3 clients.

• Compressed filesystems.

If you are not sure how your filesystem operates, then you should assume that it does not overwrite data in place, which means that shred cannot reliably operate on regular files in your filesystem.

Generally speaking, it is more reliable to shred a device than a file, since this bypasses the problem of filesystem design mentioned above. However, even shredding devices is not always completely reliable. For example, most disks map out bad sectors invisibly to the application; if the bad sectors contain sensitive data, shred won’t be able to destroy it.

shred makes no attempt to detect or report this problem, just as it makes no attempt to do anything about backups. However, since it is more reliable to shred devices than files, shred by default does not truncate or remove the output file. This default is more suitable for devices, which typically cannot be truncated and should not be removed.

Finally, consider the risk of backups and mirrors. File system backups and remote mirrors may contain copies of the file that cannot be removed, and that will allow a shredded file to be recovered later. So if you keep any data you may later want to destroy using shred, be sure that it is not backed up or mirrored.

shred [option]... file[...]

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-f’

‘--force’ Override file permissions if necessary to allow overwriting.

‘-NUMBER’
‘-n NUMBER’
‘--iterations=NUMBER’

By default, shred uses 25 passes of overwrite. This is enough for all of the useful overwrite patterns to be used at least once. You can reduce this to save time, or increase it if you have a lot of time to waste.

‘-s BYTES’
‘--size=BYTES’

Shred the first BYTES bytes of the file. The default is to shred the whole file. BYTES can be followed by a size specification like ‘K’, ‘M’, or ‘G’ to specify a multiple. See Section 2.2 [Block size], page 3.

‘-u’
‘--remove’

After shredding a file, truncate it (if possible) and then remove it. If a file has multiple links, only the named links will be removed.

‘-v’
‘--verbose’

Display status updates as sterilization proceeds.

‘-x’

‘--exact’ Normally, shred rounds the file size up to the next multiple of the filesystem block size to fully erase the last block of the file. This option suppresses that behavior. Thus, by default if you shred a 10-byte file on a system with 512-byte blocks, the resulting file will be 512 bytes long. With this option, shred does not increase the size of the file.

‘-z’

‘--zero’ Normally, the last pass that shred writes is made up of random data. If this would be conspicuous on your hard drive (for example, because it looks like encrypted data), or you just think it’s tidier, the ‘--zero’ option adds an additional overwrite pass with all zero bits. This is in addition to the number of passes specified by the ‘--iterations’ option.

‘-’ Shred standard output.

This argument is considered an option. If the common ‘--’ option has been used to indicate the end of options on the command line, then ‘-’ will be interpreted as an ordinary file name. The intended use of this is to shred a removed temporary file. For example

     i=`tempfile -m 0600`
     exec 3<>"$i"

               rm -- "$i"
               echo "Hello, world" >&3
               shred - >&3
               exec 3>-

Note that the shell command ‘shred - >file’ does not shred the contents of file, since it truncates file before invoking shred. Use the command ‘shred file’ or (if using a Bourne-compatible shell) the command ‘shred - 1<>file’ instead.

You might use the following command to erase all trace of the filesystem you’d created on the floppy disk in your first drive. That command takes about 20 minutes to erase a “1.44MB” (actually 1440 KiB) floppy.

     shred --verbose /dev/fd0

Similarly, to erase all data on a selected partition of your hard disk, you could give a command like this:

     shred --verbose /dev/sda5

12 Special file types

This chapter describes commands which create special types of files (and rmdir, which removes directories, one special file type).

Although Unix-like operating systems have markedly fewer special file types than others, not everything can be treated only as the undifferentiated byte stream of normal files. For example, when a file is created or removed, the system must record this information, which it does in a directory—a special type of file. Although you can read directories as normal files, if you’re curious, in order for the system to do its job it must impose a structure, a certain order, on the bytes of the file. Thus it is a “special” type of file.

Besides directories, other special file types include named pipes (FIFOs), symbolic links, sockets, and so-called special files.

12.1 link: Make a hard link via the link syscall

link creates a single hard link at a time. It is a minimalist interface to the systemprovided link function. See section “Hard Links” in The GNU C Library Reference Manual. Synopsis:

     link filename linkname

filename must specify an existing file, and linkname must specify a nonexistent entry in an existing directory. link simply calls link (filename, linkname) to create the link.

12.2 ln: Make links between files

ln makes links between files. By default, it makes hard links; with the ‘-s’ option, it makes symbolic (or soft) links. Synopses:

     ln [option]... target [linkname]
     ln [option]... target... directory

• If the last argument names an existing directory, ln creates a link to each target file in that directory, using the targets’ names. (But see the description of the

    ‘--no-dereference’ option below.)

• If two filenames are given, ln creates a link from the second to the first.

• If one target is given, ln creates a link to that file in the current directory.

• It is an error if the last argument is not a directory and more than two files are given. Without ‘-f’ or ‘-i’ (see below), ln will not remove an existing file. Use the ‘--backup’ option to make ln rename existing files.

A hard link is another name for an existing file; the link and the original are indistinguishable. Technically speaking, they share the same inode, and the inode contains all the information about a file—indeed, it is not incorrect to say that the inode is the file. On all existing implementations, you cannot make a hard link to a directory, and hard links cannot cross filesystem boundaries. (These restrictions are not mandated by posix, however.)

Symbolic links (symlinks for short), on the other hand, are a special file type (which not all kernels support: System V release 3 (and older) systems lack symlinks) in which the link file actually refers to a different file, by name. When most operations (opening, reading, writing, and so on) are passed the symbolic link file, the kernel automatically dereferences the link and operates on the target of the link. But some operations (e.g., removing) work on the link file itself, rather than on its target. See section “Symbolic Links” in The GNU C Library Reference Manual.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-b’
‘--backup[=method]’

See Section 2.1 [Backup options], page 2. Make a backup of each file that would otherwise be overwritten or removed.

‘-d’
‘-F’
‘--directory’

Allow the super-user to make hard links to directories.

‘-f’

‘--force’ Remove existing destination files.

‘-i’
‘--interactive’

Prompt whether to remove existing destination files.

‘-n’
‘--no-dereference’

When given an explicit destination that is a symlink to a directory, treat that destination as if it were a normal file. When the destination is an actual directory (not a symlink to one), there is no ambiguity. The link is created in that directory. But when the specified destination is a symlink to a directory, there are two ways to treat the user’s request. ln can treat the destination just as it would a normal directory and create the link in it. On the other hand, the destination can be viewed as a non-directory—as the symlink itself. In that case, ln must delete or backup that symlink before creating the new link. The default is to treat a destination that is a symlink to a directory just like a directory.

‘-s’
‘--symbolic’

Make symbolic links instead of hard links. This option merely produces an error message on systems that do not support symbolic links.

‘-S suffix’
‘--suffix=suffix’

Append suffix to each backup file made with ‘-b’. See Section 2.1 [Backup options], page 2.

‘--target-directory=directory’

Specify the destination directory. See Section 2.3 [Target directory], page 5.

‘-v’
‘--verbose’

Print the name of each file before linking it.

‘-V method’
‘--version-control=method’

Change the type of backups made with ‘-b’. The method argument can be ‘none’ (or ‘off’), ‘numbered’ (or ‘t’), ‘existing’ (or ‘nil’), or ‘never’ (or ‘simple’). See Section 2.1 [Backup options], page 2.

Examples:

     ln -s /some/name  # creates link ./name pointing to /some/name
     ln -s /some/name myname  # creates link ./myname pointing to /some/name
     ln -s a b ..      # creates links ../a and ../b pointing to ./a and ./b

12.3 mkdir: Make directories

mkdir creates directories with the specified names. Synopsis:

     mkdir [option]... name...

If a name is an existing file but not a directory, mkdir prints a warning message on stderr and will exit with a status of 1 after processing any remaining names. The same is done when a name is an existing directory and the -p option is not given. If a name is an existing directory and the -p option is given, mkdir will ignore it. That is, mkdir will not print a warning, raise an error, or change the mode of the directory (even if the -m option is given), and will move on to processing any remaining names.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-m mode’
‘--mode=mode’

Set the mode of created directories to mode, which is symbolic as in chmod and uses ‘a=rwx’ (read, write and execute allowed for everyone) minus the bits set in the umask for the point of the departure. See Chapter 26 [File permissions], page 135.

‘-p’
‘--parents’

Make any missing parent directories for each argument. The mode for parent directories is set to the umask modified by ‘u+wx’. Ignore arguments corresponding to existing directories.

‘-v’
‘--verbose’

Print a message for each created directory. This is most useful with ‘--parents’.

12.4 mkfifo: Make FIFOs (named pipes)

mkfifo creates FIFOs (also called named pipes) with the specified names. Synopsis:

     mkfifo [option] name...

A FIFO is a special file type that permits independent processes to communicate. One process opens the FIFO file for writing, and another for reading, after which data can flow as with the usual anonymous pipe in shells or elsewhere.

The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘-m mode’
‘--mode=mode’

Set the mode of created FIFOs to mode, which is symbolic as in chmod and uses ‘a=rw’ (read and write allowed for everyone) minus the bits set in the umask for the point of departure. See Chapter 26 [File permissions], page 135.

12.5 mknod: Make block or character special files

mknod creates a FIFO, character special file, or block special file with the specified name. Synopsis:

     mknod [option]... name type [major minor]

Unlike the phrase “special file type” above, the term special file has a technical meaning on Unix: something that can generate or receive data. Usually this corresponds to a physical piece of hardware, e.g., a printer or a disk. (These files are typically created at systemconfiguration time.) The mknod command is what creates files of this type. Such devices can be read either a character at a time or a “block” (many characters) at a time, hence we say there are block special files and character special files.

The arguments after name specify the type of file to make:

‘p’ for a FIFO

‘b’ for a block special file

‘c’ for a character special file

When making a block or character special file, the major and minor device numbers must be given after the file type.

The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘-m mode’
‘--mode=mode’

Set the mode of created files to mode, which is symbolic as in chmod and uses ‘a=rw’ minus the bits set in the umask as the point of departure. See Chapter 26 [File permissions], page 135.

12.6 rmdir: Remove empty directories

rmdir removes empty directories. Synopsis:

     rmdir [option]... directory...

If any directory argument does not refer to an existing empty directory, it is an error. The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘--ignore-fail-on-non-empty’

Ignore each failure to remove a directory that is solely because the directory is non-empty.

‘-p’
‘--parents’

Remove directory, then try to remove each component of directory. So, for example, ‘rmdir -p a/b/c’ is similar to ‘rmdir a/b/c a/b a’. As such, it fails if any of those directories turns out not to be empty. Use the ‘--ignore-fail-on-non-empty’ option to make it so such a failure does not evoke a diagnostic and does not cause rmdir to exit unsuccessfully.

‘-v’
‘--verbose’

Give a diagnostic for each successful removal. directory is removed.

See Section 11.5 [rm invocation], page 76, for how to remove non-empty directories (recursively).

12.7 unlink: Remove files via the unlink syscall

unlink deletes a single specified file name. It is a minimalist interface to the systemprovided unlink function. See section “Deleting Files” in The GNU C Library Reference Manual. Synopsis:

     unlink filename

On some systems unlink can be used to delete the name of a directory. On others, it can be used that way only by a privileged user. In the GNU system unlink can never delete the name of a directory.

By default, unlink honors the ‘--help’ and ‘--version’ options. That makes it a little harder to remove files named --help and --version, so when the environment variable POSIXLY_CORRECT is set, unlink treats such a command line arguments not as an option, but as an operand.

13 Changing file attributes

A file is not merely its contents, a name, and a file type (see Chapter 12 [Special file types], page 81). A file also has an owner (a userid), a group (a group id), permissions (what the owner can do with the file, what people in the group can do, and what everyone else can do), various timestamps, and other information. Collectively, we call these a file’s attributes.

These commands change file attributes.

13.1 chown: Change file owner and group

chown changes the user and/or group ownership of each given file to new-owner or to the user and group of an existing reference file. Synopsis:

     chown [option]... {new-owner | --reference=ref_file} file...

If used, new-owner specifies the new owner and/or group as follows (with no embedded white space):

     [owner] [ [:] [group] ]

Specifically:

owner If only an owner (a user name or numeric user id) is given, that user is made

the owner of each given file, and the files’ group is not changed.

owner‘:’group If the owner is followed by a colon and a group (a group name or numeric group id), with no spaces between them, the group ownership of the files is changed as well (to group).

owner‘:’ If a colon but no group name follows owner, that user is made the owner of the

files and the group of the files is changed to owner’s login group.

‘:’group If the colon and following group are given, but the owner is omitted, only the group of the files is changed; in this case, chown performs the same function as

          chgrp.

You may use ‘.’ in place of the ‘:’ separator. This is a gnu extension for compatibility with older scripts. New scripts should avoid the use of ‘.’ because gnu chown may fail if owner contains ‘.’ characters.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c’
‘--changes’

Verbosely describe the action for each file whose ownership actually changes.

‘-f’
‘--silent’

‘--quiet’ Do not print error messages about files whose ownership cannot be changed.

‘--from=old-owner’

Change a file’s ownership only if it has current attributes specified by old-owner. old-owner has the same form as new-owner described above. This option is useful primarily from a security standpoint in that it narrows considerably the window of potential abuse. For example, to reflect a UID numbering change for one user’s files without an option like this, root might run

     find / -owner OLDUSER -print0 | xargs -0 chown NEWUSER

But that is dangerous because the interval between when the find tests the existing file’s owner and when the chown is actually run may be quite large. One way to narrow the gap would be to invoke chown for each file as it is found:

     find / -owner OLDUSER -exec chown NEWUSER {} \;

But that is very slow if there are many affected files. With this option, it is safer (the gap is narrower still) though still not perfect:

     chown -R --from=OLDUSER NEWUSER /
‘--dereference’

Do not act on symbolic links themselves but rather on what they point to.

‘-h’
‘--no-dereference’

Act on symbolic links themselves instead of what they point to. This is the default. This mode relies on the lchown system call. On systems that do not provide the lchown system call, chown fails when a file specified on the command line is a symbolic link. By default, no diagnostic is issued for symbolic links encountered during a recursive traversal, but see ‘--verbose’.

‘--reference=ref_file’

Change the user and group of each file to be the same as those of ref file. If ref file is a symbolic link, do not use the user and group of the symbolic link, but rather those of the file it refers to.

‘-v’
‘--verbose’

Output a diagnostic for every file processed. If a symbolic link is encountered during a recursive traversal on a system without the lchown system call, and ‘--no-dereference’ is in effect, then issue a diagnostic saying neither the symbolic link nor its referent is being changed.

‘-R’
‘--recursive’

Recursively change ownership of directories and their contents.

13.2 chgrp: Change group ownership

chgrp changes the group ownership of each given file to group (which can be either a group name or a numeric group id) or to the group of an existing reference file. Synopsis:

chgrp [option]... {group | --reference=ref_file} file...

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c’
‘--changes’

Verbosely describe the action for each file whose group actually changes.

‘-f’
‘--silent’

‘--quiet’ Do not print error messages about files whose group cannot be changed.

‘--dereference’

Do not act on symbolic links themselves but rather on what they point to.

‘-h’
‘--no-dereference’

Act on symbolic links themselves instead of what they point to. This is the default. This mode relies on the lchown system call. On systems that do not provide the lchown system call, chgrp fails when a file specified on the command line is a symbolic link. By default, no diagnostic is issued for symbolic links encountered during a recursive traversal, but see ‘--verbose’.

‘--reference=ref_file’

Change the group of each file to be the same as that of ref file. If ref file is a symbolic link, do not use the group of the symbolic link, but rather that of the file it refers to.

‘-v’
‘--verbose’

Output a diagnostic for every file processed. If a symbolic link is encountered during a recursive traversal on a system without the lchown system call, and ‘--no-dereference’ is in effect, then issue a diagnostic saying neither the symbolic link nor its referent is being changed.

‘-R’
‘--recursive’

Recursively change the group ownership of directories and their contents.

13.3 chmod: Change access permissions

chmod changes the access permissions of the named files. Synopsis:

     chmod [option]... {mode | --reference=ref_file} file...

chmod never changes the permissions of symbolic links, since the chmod system call cannot change their permissions. This is not a problem since the permissions of symbolic links are never used. However, for each symbolic link listed on the command line, chmod changes the permissions of the pointed-to file. In contrast, chmod ignores symbolic links encountered during recursive directory traversals.

If used, mode specifies the new permissions. For details, see the section on Chapter 26 [File permissions], page 135.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c’
‘--changes’

Verbosely describe the action for each file whose permissions actually changes.

‘-f’
‘--silent’

‘--quiet’ Do not print error messages about files whose permissions cannot be changed.

‘-v’
‘--verbose’

Verbosely describe the action or non-action taken for every file.

‘--reference=ref_file’

Change the mode of each file to be the same as that of ref file. See Chapter 26 [File permissions], page 135. If ref file is a symbolic link, do not use the mode of the symbolic link, but rather that of the file it refers to.

‘-R’
‘--recursive’

Recursively change permissions of directories and their contents.

13.4 touch: Change file timestamps

touch changes the access and/or modification times of the specified files. Synopsis:

     touch [option]... file...

On older systems, touch supports an obsolete syntax, as follows. If the first file would be a valid argument to the ‘-t’ option and no timestamp is given with any of the ‘-d’, ‘-r’, or ‘-t’ options and the ‘--’ argument is not given, that argument is interpreted as the time for the other files instead of as a file name. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-t’ instead.

Any file that does not exist is created empty.

If changing both the access and modification times to the current time, touch can change the timestamps for files that the user running it does not own but has write permission for. Otherwise, the user must own the files.

Although touch provides options for changing two of the times – the times of last access and modification – of a file, there is actually a third one as well: the inode change time. This is often referred to as a file’s ctime. The inode change time represents the time when the file’s meta-information last changed. One common example of this is when the permissions of a file change. Changing the permissions doesn’t access the file, so the atime doesn’t change, nor does it modify the file, so the mtime doesn’t change. Yet, something about the file itself has changed, and this must be noted somewhere. This is the job of the ctime field. This is necessary, so that, for example, a backup program can make a fresh copy of the file, including the new permissions value. Another operation that modifies a file’s ctime without affecting the others is renaming. In any case, it is not possible, in normal operations, for a user to change the ctime field to a user-specified value.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’
‘--time=atime’
‘--time=access’
‘--time=use’

Change the access time only.

‘-c’
‘--no-create’

Do not create files that do not exist.

‘-d’
‘--date=time’

Use time instead of the current time. It can contain month names, time zones, ‘am’ and ‘pm’, etc. See Chapter 27 [Date input formats], page 141.

‘-f’ Ignored; for compatibility with BSD versions of touch.

‘-m’
‘--time=mtime’
‘--time=modify’

Change the modification time only.

‘-r file’
‘--reference=file’

Use the times of the reference file instead of the current time.

‘-t [[CC]YY]MMDDhhmm[.ss]’

Use the argument (optional four-digit or two-digit years, months, days, hours, minutes, optional seconds) instead of the current time. If the year is specified with only two digits, then CC is 20 for years in the range 0 … 68, and 19 for years in 69 … 99. If no digits of the year are specified, the argument is interpreted as a date in the current year.

14 Disk usage

No disk can hold an infinite amount of data. These commands report on how much disk storage is in use or available. (This has nothing much to do with how much main memory, i.e., RAM, a program is using when it runs; for that, you want ps or pstat or swap or some such command.)

14.1 df: Report filesystem disk space usage

df reports the amount of disk space used and available on filesystems. Synopsis:

     df [option]... [file]...

With no arguments, df reports the space used and available on all currently mounted filesystems (of all types). Otherwise, df reports on the filesystem containing each argument file.

Normally the disk space is printed in units of 1024 bytes, but this can be overridden (see Section 2.2 [Block size], page 3). Non-integer quantities are rounded up to the next higher unit.

If an argument file is a disk device file containing a mounted filesystem, df shows the space available on that filesystem rather than on the filesystem containing the device node (i.e., the root filesystem). gnu df does not attempt to determine the disk usage on unmounted filesystems, because on most kinds of systems doing so requires extremely nonportable intimate knowledge of filesystem structures.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’

‘--all’ Include in the listing filesystems that have a size of 0 blocks, which are omitted by default. Such filesystems are typically special-purpose pseudo-filesystems, such as automounter entries. Also, filesystems of type “ignore” or “auto”, supported by some operating systems, are only included if this option is specified.

‘-B size’
‘--block-size=size’

Scale sizes by size before printing them (see Section 2.2 [Block size], page 3). For example, ‘-BG’ prints sizes in units of 1,073,741,824 bytes.

‘-h’
‘--human-readable’

Append a size letter to each size, such as ‘M’ for mebibytes. Powers of 1024 are used, not 1000; ‘M’ stands for 1,048,576 bytes. Use the ‘-H’ or ‘--si’ option if you prefer powers of 1000.

‘-H’

‘--si’ Append an SI-style abbreviation to each size, such as ‘MB’ for megabytes. Powers of 1000 are used, not 1024; ‘MB’ stands for 1,000,000 bytes. Use the ‘-h’ or ‘--human-readable’ option if you prefer powers of 1024.

‘-i’
‘--inodes’

List inode usage information instead of block usage. An inode (short for index node) contains information about a file such as its owner, permissions, timestamps, and location on the disk.

‘-k’ Print sizes in 1024-byte blocks, overriding the default block size (see Section 2.2 [Block size], page 3). This option is equivalent to ‘--block-size=1K’.

‘-l’

‘--local’ Limit the listing to local filesystems. By default, remote filesystems are also listed.

‘--no-sync’

Do not invoke the sync system call before getting any usage data. This may make df run significantly faster on systems with many disks, but on some systems (notably SunOS) the results may be slightly out of date. This is the default.

‘-P’
‘--portability’

Use the posix output format. This is like the default format except for the following:

1. The information about each filesystem is always printed on exactly one line; a mount device is never put on a line by itself. This means that if the mount device name is more than 20 characters long (e.g., for some network mounts), the columns are misaligned.

2. The labels in the header output line are changed to conform to posix.

‘--sync’ Invoke the sync system call before getting any usage data. On some systems (notably SunOS), doing this yields more up to date results, but in general this option makes df much slower, especially when there are many or very busy filesystems.

‘-t fstype’
‘--type=fstype’

Limit the listing to filesystems of type fstype. Multiple filesystem types can be specified by giving multiple ‘-t’ options. By default, nothing is omitted.

‘-T’
‘--print-type’

Print each filesystem’s type. The types printed here are the same ones you can include or exclude with ‘-t’ and ‘-x’. The particular types printed are whatever is supported by the system. Here are some of the common names (this list is certainly not exhaustive):

‘nfs’ An NFS filesystem, i.e., one mounted over a network from another machine. This is the one type name which seems to be used uniformly by all systems.

          ‘4.2, ufs, efs...’

A filesystem on a locally-mounted hard disk. (The system might even support more than one type here; Linux does.)

          ‘hsfs, cdfs’

A filesystem on a CD-ROM drive. HP-UX uses ‘cdfs’, most other systems use ‘hsfs’ (‘hs’ for “High Sierra”).

‘pcfs’ An MS-DOS filesystem, usually on a diskette.

‘-x fstype’
‘--exclude-type=fstype’

Limit the listing to filesystems not of type fstype. Multiple filesystem types can be eliminated by giving multiple ‘-x’ options. By default, no filesystem types are omitted.

‘-v’ Ignored; for compatibility with System V versions of df.

14.2 du: Estimate file space usage

du reports the amount of disk space used by the specified files and for each subdirectory (of directory arguments). Synopsis:

     du [option]... [file]...

With no arguments, du reports the disk space for the current directory. Normally the disk space is printed in units of 1024 bytes, but this can be overridden (see Section 2.2 [Block size], page 3). Non-integer quantities are rounded up to the next higher unit.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’

‘--all’ Show counts for all files, not just directories.

‘-b’

‘--bytes’ Print sizes in bytes, overriding the default block size (see Section 2.2 [Block size], page 3).

‘-B size’
‘--block-size=size’

Scale sizes by size before printing them (see Section 2.2 [Block size], page 3). For example, ‘-BG’ prints sizes in units of 1,073,741,824 bytes.

‘-c’

‘--total’ Print a grand total of all arguments after all arguments have been processed. This can be used to find out the total disk usage of a given set of files or directories.

‘-D’
‘--dereference-args’

Dereference symbolic links that are command line arguments. Does not affect other symbolic links. This is helpful for finding out the disk usage of directories, such as ‘/usr/tmp’, which are often symbolic links.

‘-h’
‘--human-readable’

Append a size letter to each size, such as ‘M’ for mebibytes. Powers of 1024 are used, not 1000; ‘M’ stands for 1,048,576 bytes. Use the ‘-H’ or ‘--si’ option if you prefer powers of 1000.

‘-H’

‘--si’ Append an SI-style abbreviation to each size, such as ‘MB’ for megabytes. Powers of 1000 are used, not 1024; ‘MB’ stands for 1,000,000 bytes. Use the ‘-h’ or ‘--human-readable’ option if you prefer powers of 1024.

‘-k’ Print sizes in 1024-byte blocks, overriding the default block size (see Section 2.2 [Block size], page 3). This option is equivalent to ‘--block-size=1K’.

‘-l’
‘--count-links’

Count the size of all files, even if they have appeared already (as a hard link).

‘-L’
‘--dereference’

Dereference symbolic links (show the disk space used by the file or directory that the link points to instead of the space used by the link).

‘--max-depth=DEPTH’

Show the total for each directory (and file if –all) that is at most MAX DEPTH levels down from the root of the hierarchy. The root is at level 0, so du --max-depth=0 is equivalent to du -s.

‘-s’
‘--summarize’

Display only a total for each argument.

‘-S’
‘--separate-dirs’

Report the size of each directory separately, not including the sizes of subdirectories.

‘-x’
‘--one-file-system’

Skip directories that are on different filesystems from the one that the argument being processed is on.

‘--exclude=PATTERN’

When recursing, skip subdirectories or files matching PATTERN. For example, du --exclude='*.o' excludes files whose names end in ‘.o’.

‘-X FILE’
‘--exclude-from=FILE’

Like ‘--exclude’, except take the patterns to exclude from FILE, one per line. If FILE is ‘-’, take the patterns from standard input.

On BSD systems, du reports sizes that are half the correct values for files that are NFSmounted from HP-UX systems. On HP-UX systems, it reports sizes that are twice the correct values for files that are NFS-mounted from BSD systems. This is due to a flaw in HP-UX; it also affects the HP-UX du program.

14.3 stat: Report file or filesystem status

stat displays information about the specified file(s). Synopsis:

stat [option]... [file]...

With no option, stat reports all information about the given files. But it also can be used to report the information of the filesystems the given files are located on. If the files are links, stat can also give information about the files the links point to.

‘-f’
‘--filesystem’

Report information about the filesystems where the given files are located instead of information about the files themselves.

‘-L’
‘--dereference’

Change how stat treats symbolic links. With this option, stat acts on the file referenced by each symbolic link argument. Without it, stat acts on any symbolic link argument directly.

‘-t’

‘--terse’ Print the information in terse form, suitable for parsing by other programs.

‘-c’
‘--format’

Allow user to specify the output format. Interpreted sequences for file stat are:

• %n - File name

• %N - Quoted File name with dereference if symbolic link

• %d - Device number in decimal

• %D - Device number in hex

• %i - Inode number

• %a - Access rights in octal

• %A - Access rights in human readable form

• %f - raw mode in hex

• %F - File type

• %h - Number of hard links

• %u - User Id of owner

• %U - User name of owner

• %g - Group Id of owner

• %G - Group name of owner

• %t - Major device type in hex • %T - Minor device type in hex

• %s - Total size, in bytes

• %b - Number of blocks allocated

• %o - IO block size

• %x - Time of last access

• %X - Time of last access as seconds since Epoch

• %y - Time of last modification

• %Y - Time of last modification as seconds since Epoch • %z - Time of last change

• %Z - Time of last change as seconds since Epoch Interpreted sequences for filesystem stat are:

• %n - File name

• %i - File System id in hex

• %l - Maximum length of filenames

• %t - Type in hex

• %T - Type in human readable form

• %b - Total data blocks in file system

• %f - Free blocks in file system

• %a - Free blocks available to non-superuser

• %s - Optimal transfer block size

• %c - Total file nodes in file system

14.4 sync: Synchronize data on disk with memory

sync writes any data buffered in memory out to disk. This can include (but is not limited to) modified superblocks, modified inodes, and delayed reads and writes. This must be implemented by the kernel; The sync program does nothing but exercise the sync system call.

The kernel keeps data in memory to avoid doing (relatively slow) disk reads and writes. This improves performance, but if the computer crashes, data may be lost or the filesystem corrupted as a result. sync ensures everything in memory is written to disk.

Any arguments are ignored, except for a lone ‘--help’ or ‘--version’ (see Chapter 2 [Common options], page 2).

15 Printing text

This section describes commands that display text strings.

15.1 echo: Print a line of text

echo writes each given string to standard output, with a space between each and a newline after the last one. Synopsis:

echo [option]... [string]...

The program accepts the following options. Also see Chapter 2 [Common options], page 2. ‘-n’ Do not output the trailing newline. ‘-e’ Enable interpretation of the following backslash-escaped characters in each

string: ‘\a’ alert (bell) ‘\b’ backspace ‘\c’ suppress trailing newline ‘\f’ form feed ‘\n’ new line ‘\r’ carriage return ‘\t’ horizontal tab ‘\v’ vertical tab ‘\\’ backslash ‘\nnn’ the character whose ascii code is nnn (octal); if nnn is not a valid

octal number, it is printed literally.

15.2 printf: Format and print data

printf does formatted printing of text. Synopsis:

printf format [argument]...

printf prints the format string, interpreting ‘%’ directives and ‘\’ escapes in the same way as the C printf function. The format argument is re-used as necessary to convert all of the given arguments.

printf has one additional directive, ‘%b’, which prints its argument string with ‘\’ escapes interpreted in the same way as in the format string.

printf interprets ‘\0ooo’ in format as an octal number (if ooo is 0 to 3 octal digits) specifying a character to print, and ‘\xhh’ as a hexadecimal number (if hh is 1 to 2 hex digits) specifying a character to print.

printf interprets two character syntaxes introduced in ISO C 99: ‘\u’ for 16-bit Unicode characters, specified as 4 hex digits hhhh, and ‘\U’ for 32-bit Unicode characters, specified as 8 hex digits hhhhhhhh. printf outputs the Unicode characters according to the LC CTYPE part of the current locale, i.e. depending on the values of the environment variables LC_ALL,

LC_CTYPE, LANG.

The processing of ‘\u’ and ‘\U’ requires a full-featured iconv facility. It is activated on systems with glibc 2.2 (or newer), or when libiconv is installed prior to this package. Otherwise the use of ‘\u’ and ‘\U’ will give an error message.

An additional escape, ‘\c’, causes printf to produce no further output.

The only options are a lone ‘--help’ or ‘--version’. See Chapter 2 [Common options], page 2.

The Unicode character syntaxes are useful for writing strings in a locale independent way. For example, a string containing the Euro currency symbol

     $ /usr/local/bin/printf '\u20AC 14.95'

will be output correctly in all locales supporting the Euro symbol (ISO-8859-15, UTF-8, and others). Similarly, a Chinese string

     $ /usr/local/bin/printf '\u4e2d\u6587'

will be output correctly in all Chinese locales (GB2312, BIG5, UTF-8, etc).

Note that in these examples, the full pathname of printf has been given, to distinguish it from the GNU bash builtin function printf.

For larger strings, you don’t need to look up the hexadecimal code values of each character one by one. ascii characters mixed with \u escape sequences is also known as the JAVA source file encoding. You can use GNU recode 3.5c (or newer) to convert strings to this encoding. Here is how to convert a piece of text into a shell script which will output this text in a locale-independent way:

     $ LC_CTYPE=zh_CN.big5 /usr/local/bin/printf \
         '\u4e2d\u6587\n' > sample.txt
     $ recode BIG5..JAVA < sample.txt \
         | sed -e "s|^|/usr/local/bin/printf '|" -e "s|$|\\\\n'|" \
         > sample.sh

15.3 yes: Print a string until interrupted

yes prints the command line arguments, separated by spaces and followed by a newline, forever until it is killed. If no arguments are given, it prints ‘y’ followed by a newline forever until killed.

The only options are a lone ‘--help’ or ‘--version’. See Chapter 2 [Common options], page 2.

16 Conditions

This section describes commands that are primarily useful for their exit status, rather than their output. Thus, they are often used as the condition of shell if statements, or as the last command in a pipeline.

16.1 false: Do nothing, unsuccessfully

false does nothing except return an exit status of 1, meaning failure. It can be used as a place holder in shell scripts where an unsuccessful command is needed.

By default, false honors the ‘--help’ and ‘--version’ options. However, that is contrary to posix, so when the environment variable POSIXLY_CORRECT is set, false ignores all command line arguments, including ‘--help’ and ‘--version’.

This version of false is implemented as a C program, and is thus more secure and faster than a shell script implementation, and may safely be used as a dummy shell for the purpose of disabling accounts.

16.2 true: Do nothing, successfully

true does nothing except return an exit status of 0, meaning success. It can be used as a place holder in shell scripts where a successful command is needed, although the shell built-in command : (colon) may do the same thing faster. In most modern shells, true is a built-in command, so when you use ‘true’ in a script, you’re probably using the built-in command, not the one documented here.

By default, true honors the ‘--help’ and ‘--version’ options. However, that is contrary to posix, so when the environment variable POSIXLY_CORRECT is set, true ignores all command line arguments, including ‘--help’ and ‘--version’.

This version of true is implemented as a C program, and is thus more secure and faster than a shell script implementation, and may safely be used as a dummy shell for the purpose of disabling accounts.

16.3 test: Check file types and compare values

test returns a status of 0 (true) or 1 (false) depending on the evaluation of the conditional expression expr. Each part of the expression must be a separate argument.

test has file status checks, string operators, and numeric comparison operators.

Because most shells have a built-in command by the same name, using the unadorned command name in a script or interactively may get you different functionality than that described here.

Besides the options below, test accepts a lone ‘--help’ or ‘--version’. See Chapter 2 [Common options], page 2. A single non-option argument is also allowed: test returns true if the argument is not null.

16.3.1 File type tests

These options test for particular types of files. (Everything’s a file, but not all files are the same!) ‘-b file’ True if file exists and is a block special device. ‘-c file’ True if file exists and is a character special device. ‘-d file’ True if file exists and is a directory. ‘-f file’ True if file exists and is a regular file.

‘-h file’

‘-L file’ True if file exists and is a symbolic link. ‘-p file’ True if file exists and is a named pipe. ‘-S file’ True if file exists and is a socket. ‘-t [fd]’ True if fd is opened on a terminal. If fd is omitted, it defaults to 1 (standard

output).

16.3.2 Access permission tests

These options test for particular access permissions. ‘-g file’ True if file exists and has its set-group-id bit set. ‘-k file’ True if file has its sticky bit set. ‘-r file’ True if file exists and is readable. ‘-u file’ True if file exists and has its set-user-id bit set. ‘-w file’ True if file exists and is writable. ‘-x file’ True if file exists and is executable. ‘-O file’ True if file exists and is owned by the current effective user id. ‘-G file’ True if file exists and is owned by the current effective group id.

16.3.3 File characteristic tests

These options test other file characteristics. ‘-e file’ True if file exists. ‘-s file’ True if file exists and has a size greater than zero.

‘file1 -nt file2’

True if file1 is newer (according to modification date) than file2, or if file1 exists and file2 does not.

‘file1 -ot file2’

True if file1 is older (according to modification date) than file2, or if file2 exists and file1 does not.

‘file1 -ef file2’

True if file1 and file2 have the same device and inode numbers, i.e., if they are hard links to each other.

16.3.4 String tests

These options test string characteristics. Strings are not quoted for test, though you may need to quote them to protect characters with special meaning to the shell, e.g., spaces.

‘-z string’

True if the length of string is zero.

‘-n string’

‘string’ True if the length of string is nonzero.

‘string1 = string2’

True if the strings are equal.

‘string1 != string2’

True if the strings are not equal.

16.3.5 Numeric tests

Numeric relationals. The arguments must be entirely numeric (possibly negative), or the special expression -l string, which evaluates to the length of string.

‘arg1 -eq arg2’
‘arg1 -ne arg2’
‘arg1 -lt arg2’
‘arg1 -le arg2’
‘arg1 -gt arg2’
‘arg1 -ge arg2’

These arithmetic binary operators return true if arg1 is equal, not-equal, lessthan, less-than-or-equal, greater-than, or greater-than-or-equal than arg2, respectively.

For example:

     test -1 -gt -2 && echo yes
     ⇒ yes
     test -l abc -gt 1 && echo yes
     ⇒ yes
     test 0x100 -eq 1
      error test: integer expression expected before -eq

16.3.6 Connectives for test

The usual logical connectives.

‘! expr’ True if expr is false.

‘expr1 -a expr2’

True if both expr1 and expr2 are true.

‘expr1 -o expr2’

True if either expr1 or expr2 is true.

16.4 expr: Evaluate expressions

expr evaluates an expression and writes the result on standard output. Each token of the expression must be a separate argument.

Operands are either numbers or strings. expr converts anything appearing in an operand position to an integer or a string depending on the operation being applied to it.

Strings are not quoted for expr itself, though you may need to quote them to protect characters with special meaning to the shell, e.g., spaces.

Operators may be given as infix symbols or prefix keywords. Parentheses may be used for grouping in the usual manner (you must quote parentheses to avoid the shell evaluating them, however).

Exit status:

0 if the expression is neither null nor 0, 1 if the expression is null or 0, 2 for invalid expressions.

16.4.1 String expressions

expr supports pattern matching and other string operators. These have lower precedence than both the numeric and relational operators (in the next sections).

‘string : regex’

Perform pattern matching. The arguments are converted to strings and the second is considered to be a (basic, a la GNU grep) regular expression, with a ^implicitly prepended. The first argument is then matched against this regular expression. If the match succeeds and regex uses ‘\(’ and ‘\)’, the : expression returns the part of string that matched the subexpression; otherwise, it returns the number of characters matched. If the match fails, the : operator returns the null string if ‘\(’ and ‘\)’ are used in regex, otherwise 0. Only the first ‘\( ... \)’ pair is relevant to the return value; additional pairs are meaningful only for grouping the regular expression operators. In the regular expression, \+, \?, and \| are operators which respectively match one or more, zero or one, or separate alternatives. SunOS and other expr’s treat these as regular characters. (posix allows either behavior.) See section “Regular Expression Library” in Regex, for details of regular expression syntax. Some examples are in Section 16.4.4 [Examples of expr], page 104.

‘match string regex’

An alternative way to do pattern matching. This is the same as

‘string : regex’.
‘substr string position length’

Returns the substring of string beginning at position with length at most length. If either position or length is negative, zero, or non-numeric, returns the null string.

‘index string charset’

Returns the first position in string where the first character in charset was found. If no character in charset is found in string, return 0.

‘length string’

Returns the length of string.

‘+ token’ Interpret token as a string, even if it is a keyword like match or an operator like /. This makes it possible to test expr length + "$x" or expr + "$x" : '.*/\(.\)' and have it do the right thing even if the value of $x happens to be (for example) / or index. This operator is a GNU extension. Portable shell

scripts should use " $token" : ' \(.*\)' instead of + "$token".

To make expr interpret keywords as strings, you must use the quote operator.

16.4.2 Numeric expressions

expr supports the usual numeric operators, in order of increasing precedence. The string operators (previous section) have lower precedence, the connectives (next section) have higher.

‘+ -’ Addition and subtraction. Both arguments are converted to numbers; an error occurs if this cannot be done.

‘* / %’ Multiplication, division, remainder. Both arguments are converted to numbers; an error occurs if this cannot be done.

16.4.3 Relations for expr

expr supports the usual logical connectives and relations. These are higher precedence than either the string or numeric operators (previous sections). Here is the list, lowestprecedence operator first.

‘|’ Returns its first argument if that is neither null nor 0, otherwise its second argument.

‘&’ Return its first argument if neither argument is null or 0, otherwise 0.

‘< <= = == != >= >’

Compare the arguments and return 1 if the relation is true, 0 otherwise. == is a synonym for =. expr first tries to convert both arguments to numbers and do a numeric comparison; if either conversion fails, it does a lexicographic comparison using the character collating sequence specified by the LC_COLLATE locale.

16.4.4 Examples of using expr

Here are a few examples, including quoting for shell metacharacters. To add 1 to the shell variable foo, in Bourne-compatible shells:

foo=`expr $foo + 1`

To print the non-directory part of the file name stored in $fname, which need not contain a /.

expr $fname : '.*/\(.*\)' '|' $fname

An example showing that \+ is an operator:

expr aaa : 'a\+'
⇒ 3
expr abc : 'a\(.\)c'
⇒ b
expr index abcdef cz
⇒ 3
expr index index a
 error expr: syntax error
expr index quote index a
⇒ 0

17 Redirection

Unix shells commonly provide several forms of redirection—ways to change the input source or output destination of a command. But one useful redirection is performed by a separate command, not by the shell; it’s described here.

17.1 tee: Redirect output to multiple files

The tee command copies standard input to standard output and also to any files given as arguments. This is useful when you want not only to send some data down a pipe, but also to save a copy. Synopsis:

     tee [option]... [file]...

If a file being written to does not already exist, it is created. If a file being written to already exists, the data it previously contained is overwritten unless the -a option is used.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’
‘--append’

Append standard input to the given files rather than overwriting them.

‘-i’
‘--ignore-interrupts’

Ignore interrupt signals.

18 File name manipulation

This section describes commands that manipulate file names.

18.1 basename: Strip directory and suffix from a file name

basename removes any leading directory components from name. Synopsis:

  basename name [suffix]

If suffix is specified and is identical to the end of name, it is removed from name as well. basename prints the result on standard output.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

18.2 dirname: Strip non-directory suffix from a file name

dirname prints all but the final slash-delimited component of a string (presumably a filename). Synopsis:

  dirname name

If name is a single component, dirname prints ‘.’ (meaning the current directory). The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options],

page 2.

18.3 pathchk: Check file name portability

pathchk checks portability of filenames. Synopsis:

  pathchk [option]... name...

For each name, pathchk prints a message if any of these conditions is true:

1. one of the existing directories in name does not have search (execute) permission,

2. the length of name is larger than its filesystem’s maximum file name length,

3. the length of one component of name, corresponding to an existing directory name, is larger than its filesystem’s maximum length for a file name component. The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘-p’
‘--portability’

Instead of performing length checks on the underlying filesystem, test the length of each file name and its components against the posix minimum limits for portability. Also check that the file name contains no characters not in the portable file name character set.

Exit status:

0 if all specified file names passed all of the tests, 1 otherwise.

19 Working context

This section describes commands that display or alter the context in which you are working: the current directory, the terminal settings, and so forth. See also the user-related commands in the next section.

19.1 pwd: Print working directory

pwd prints the fully resolved name of the current directory. That is, all components of the printed name will be actual directory names—none will be symbolic links.

Because most shells have a built-in command by the same name, using the unadorned command name in a script or interactively may get you different functionality than that described here.

The only options are a lone ‘--help’ or ‘--version’. See Chapter 2 [Common options], page 2.

19.2 stty: Print or change terminal characteristics

stty prints or changes terminal characteristics, such as baud rate. Synopses:

     stty [option] [setting]...
     stty [option]

If given no line settings, stty prints the baud rate, line discipline number (on systems that support it), and line settings that have been changed from the values set by ‘stty sane’. By default, mode reading and setting are performed on the tty line connected to standard input, although this can be modified by the ‘--file’ option.

stty accepts many non-option arguments that change aspects of the terminal line operation, as described below.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’

‘--all’ Print all current settings in human-readable form. This option may not be used in combination with any line settings.

‘-F device’
‘--file=device’

Set the line opened by the filename specified in device instead of the tty line connected to standard input. This option is necessary because opening a posix tty requires use of the O_NONDELAY flag to prevent a posix tty from blocking until the carrier detect line is high if the clocal flag is not set. Hence, it is not always possible to allow the shell to open the device in the traditional manner.

‘-g’

‘--save’ Print all current settings in a form that can be used as an argument to another stty command to restore the current settings. This option may not be used in combination with any line settings.

Many settings can be turned off by preceding them with a ‘-’. Such arguments are marked below with “May be negated” in their description. The descriptions themselves refer to the positive case, that is, when not negated (unless stated otherwise, of course).

Some settings are not available on all posix systems, since they use extensions. Such arguments are marked below with “Non-posix” in their description. On non-posix systems, those or other settings also may not be available, but it’s not feasible to document all the variations: just try it and see.

19.2.1 Control settings

Control settings: ‘parenb’ Generate parity bit in output and expect parity bit in input. May be negated. ‘parodd’ Set odd parity (even if negated). May be negated.

‘cs5’
‘cs6’
‘cs7’

‘cs8’ Set character size to 5, 6, 7, or 8 bits.

‘hup’

‘hupcl’ Send a hangup signal when the last process closes the tty. May be negated. ‘cstopb’ Use two stop bits per character (one if negated). May be negated. ‘cread’ Allow input to be received. May be negated. ‘clocal’ Disable modem control signals. May be negated. ‘crtscts’ Enable RTS/CTS flow control. Non-posix. May be negated.

19.2.2 Input settings

‘ignbrk’ Ignore break characters. May be negated. ‘brkint’ Make breaks cause an interrupt signal. May be negated. ‘ignpar’ Ignore characters with parity errors. May be negated. ‘parmrk’ Mark parity errors (with a 255-0-character sequence). May be negated. ‘inpck’ Enable input parity checking. May be negated. ‘istrip’ Clear high (8th) bit of input characters. May be negated. ‘inlcr’ Translate newline to carriage return. May be negated. ‘igncr’ Ignore carriage return. May be negated. ‘icrnl’ Translate carriage return to newline. May be negated. ‘ixon’ Enable XON/XOFF flow control (that is, CTRL-S/CTRL-Q). May be negated.

‘ixoff’

‘tandem’ Enable sending of stop character when the system input buffer is almost full, and start character when it becomes almost empty again. May be negated.

‘iuclc’ Translate uppercase characters to lowercase. Non-posix. May be negated. ‘ixany’ Allow any character to restart output (only the start character if negated).

Non-posix. May be negated.

‘imaxbel’ Enable beeping and not flushing input buffer if a character arrives when the input buffer is full. Non-posix. May be negated.

19.2.3 Output settings

These arguments specify output-related operations. ‘opost’ Postprocess output. May be negated. ‘olcuc’ Translate lowercase characters to uppercase. Non-posix. May be negated. ‘ocrnl’ Translate carriage return to newline. Non-posix. May be negated. ‘onlcr’ Translate newline to carriage return-newline. Non-posix. May be negated. ‘onocr’ Do not print carriage returns in the first column. Non-posix. May be negated. ‘onlret’ Newline performs a carriage return. Non-posix. May be negated. ‘ofill’ Use fill (padding) characters instead of timing for delays. Non-posix. May be

negated.

‘ofdel’ Use delete characters for fill instead of null characters. Non-posix. May be

negated.

‘nl1’

‘nl0’ Newline delay style. Non-posix.

‘cr3’
‘cr2’
‘cr1’

‘cr0’ Carriage return delay style. Non-posix.

‘tab3’
‘tab2’
‘tab1’

‘tab0’ Horizontal tab delay style. Non-posix.

‘bs1’

‘bs0’ Backspace delay style. Non-posix.

‘vt1’

‘vt0’ Vertical tab delay style. Non-posix.

‘ff1’

‘ff0’ Form feed delay style. Non-posix.

19.2.4 Local settings

‘isig’ Enable interrupt, quit, and suspend special characters. May be negated.

‘icanon’ Enable erase, kill, werase, and rprnt special characters. May be negated.

‘iexten’ Enable non-posix special characters. May be negated.

‘echo’ Echo input characters. May be negated.

‘echoe’
‘crterase’

Echo erase characters as backspace-space-backspace. May be negated.

‘echok’ Echo a newline after a kill character. May be negated.

‘echonl’ Echo newline even if not echoing other characters. May be negated.

‘noflsh’ Disable flushing after interrupt and quit special characters. May be negated.

‘xcase’ Enable input and output of uppercase characters by preceding their lowercase equivalents with ‘\’, when icanon is set. Non-posix. May be negated.

‘tostop’ Stop background jobs that try to write to the terminal. Non-posix. May be negated.

‘echoprt’
‘prterase’

Echo erased characters backward, between ‘\’ and ‘/’. Non-posix. May be negated.

‘echoctl’

‘ctlecho’ Echo control characters in hat notation (‘^c’) instead of literally. Non-posix. May be negated.

‘echoke’

‘crtkill’ Echo the kill special character by erasing each character on the line as indicated by the echoprt and echoe settings, instead of by the echoctl and echok settings. Non-posix. May be negated.

19.2.5 Combination settings

Combination settings:

‘evenp’

‘parity’ Same as parenb -parodd cs7. May be negated. If negated, same as -parenb

          cs8.

‘oddp’ Same as parenb parodd cs7. May be negated. If negated, same as -parenb

          cs8.

‘nl’ Same as -icrnl -onlcr. May be negated. If negated, same as icrnl -inlcr

          -igncr onlcr -ocrnl -onlret.

‘ek’ Reset the erase and kill special characters to their default values.

‘sane’ Same as:

               cread -ignbrk brkint -inlcr -igncr icrnl -ixoff
               -iuclc -ixany imaxbel opost -olcuc -ocrnl onlcr
               -onocr -onlret -ofill -ofdel nl0 cr0 tab0 bs0 vt0
               ff0 isig icanon iexten echo echoe echok -echonl
               -noflsh -xcase -tostop -echoprt echoctl echoke

and also sets all special characters to their default values.

‘cooked’ Same as brkint ignpar istrip icrnl ixon opost isig icanon, plus sets the

eof and eol characters to their default values if they are the same as the min and time characters. May be negated. If negated, same as raw.

‘raw’ Same as:

               -ignbrk -brkint -ignpar -parmrk -inpck -istrip
               -inlcr -igncr -icrnl -ixon -ixoff -iuclc -ixany
               -imaxbel -opost -isig -icanon -xcase min 1 time 0

May be negated. If negated, same as cooked.

‘cbreak’ Same as -icanon. May be negated. If negated, same as icanon.

‘pass8’ Same as -parenb -istrip cs8. May be negated. If negated, same as parenb

          istrip cs7.

‘litout’ Same as -parenb -istrip -opost cs8. May be negated. If negated, same as

          parenb istrip opost cs7.

‘decctlq’ Same as -ixany. Non-posix. May be negated.

‘tabs’ Same as tab0. Non-posix. May be negated. If negated, same as tab3.

‘lcase’

‘LCASE’ Same as xcase iuclc olcuc. Non-posix. May be negated.

‘crt’  Same as echoe echoctl echoke.
‘dec’  Same as echoe echoctl echoke -ixany intr ^C erase ^? kill C-u.

19.2.6 Special characters

The special characters’ default values vary from system to system. They are set with the syntax ‘name value’, where the names are listed below and the value can be given either literally, in hat notation (‘^c’), or as an integer which may start with ‘0x’ to indicate hexadecimal, ‘0’ to indicate octal, or any other digit to indicate decimal.

For GNU stty, giving a value of ^- or undef disables that special character. (This is incompatible with Ultrix stty, which uses a value of ‘u’ to disable a special character. GNU stty treats a value ‘u’ like any other, namely to set that special character to ^{⟨}\underline{\overline{U}}^{⟩}.)

‘intr’ Send an interrupt signal.

‘quit’ Send a quit signal.

‘erase’ Erase the last character typed.

‘kill’ Erase the current line. ‘eof’ Send an end of file (terminate the input). ‘eol’ End the line. ‘eol2’ Alternate character to end the line. Non-posix. ‘swtch’ Switch to a different shell layer. Non-posix. ‘start’ Restart the output after stopping it. ‘stop’ Stop the output. ‘susp’ Send a terminal stop signal. ‘dsusp’ Send a terminal stop signal after flushing the input. Non-posix. ‘rprnt’ Redraw the current line. Non-posix. ‘werase’ Erase the last word typed. Non-posix. ‘lnext’ Enter the next character typed literally, even if it is a special character. Non-

posix.

19.2.7 Special settings

‘min n’ Set the minimum number of characters that will satisfy a read until the time

value has expired, when -icanon is set.

‘time n’ Set the number of tenths of a second before reads time out if the minimum

number of characters have not been read, when -icanon is set.

‘ispeed n’

Set the input speed to n.

‘ospeed n’

Set the output speed to n.

‘rows n’ Tell the tty kernel driver that the terminal has n rows. Non-posix.

‘cols n’
‘columns n’

Tell the kernel that the terminal has n columns. Non-posix.

‘size’ Print the number of rows and columns that the kernel thinks the terminal has. (Systems that don’t support rows and columns in the kernel typically use the environment variables LINES and COLUMNS instead; however, GNU stty does not know anything about them.) Non-posix.

‘line n’ Use line discipline n. Non-posix. ‘speed’ Print the terminal speed. ‘n’ Set the input and output speeds to n. n can be one of: 0 50 75 110 134 134.5

150 200 300 600 1200 1800 2400 4800 9600 19200 38400 exta extb. exta is the same as 19200; extb is the same as 38400. 0 hangs up the line if -clocal is set.

19.3 printenv: Print all or some environment variables

printenv prints environment variable values. Synopsis:

printenv [option] [variable]...

If no variables are specified, printenv prints the value of every environment variable. Otherwise, it prints the value of each variable that is set, and nothing for those that are not set.

The only options are a lone ‘--help’ or ‘--version’. See Chapter 2 [Common options], page 2.

Exit status:

0 if all variables specified were found 1 if at least one specified variable was not found 2 if a write error occurred

19.4 tty: Print file name of terminal on standard input

tty prints the file name of the terminal connected to its standard input. It prints ‘not a tty’ if standard input is not a terminal. Synopsis:

tty [option]...

The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘-s’
‘--silent’

‘--quiet’ Print nothing; only return an exit status.

Exit status:

0 if standard input is a terminal 1 if standard input is not a terminal 2 if given incorrect arguments 3 if a write error occurs

20 User information

This section describes commands that print user-related information: logins, groups, and so forth.

20.1 id: Print real and effective uid and gid

id prints information about the given user, or the process running it if no user is specified. Synopsis:

     id [option]... [username]

By default, it prints the real user id, real group id, effective user id if different from the real user id, effective group id if different from the real group id, and supplemental group ids.

Each of these numeric values is preceded by an identifying string and followed by the corresponding user or group name in parentheses.

The options cause id to print only part of the above information. Also see Chapter 2 [Common options], page 2.

‘-g’

‘--group’ Print only the group id.

‘-G’
‘--groups’

Print only the supplementary groups.

‘-n’

‘--name’ Print the user or group name instead of the ID number. Requires -u, -g, or -G.

‘-r’

‘--real’ Print the real, instead of effective, user or group id. Requires -u, -g, or -G.

‘-u’

‘--user’ Print only the user id.

20.2 logname: Print current login name

logname prints the calling user’s name, as found in the file ‘/etc/utmp’, and exits with a status of 0. If there is no ‘/etc/utmp’ entry for the calling process, logname prints an error message and exits with a status of 1.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

20.3 whoami: Print effective user id

whoami prints the user name associated with the current effective user id. It is equivalent to the command ‘id -un’.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

20.4 groups: Print group names a user is in

groups prints the names of the primary and any supplementary groups for each given username, or the current process if no names are given. If names are given, the name of each user is printed before the list of that user’s groups. Synopsis:

     groups [username]...

The group lists are equivalent to the output of the command ‘id -Gn’.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

20.5 users: Print login names of users currently logged in

users prints on a single line a blank-separated list of user names of users currently logged in to the current host. Each user name corresponds to a login session, so if a user has more than one login session, that user’s name will appear the same number of times in the output. Synopsis:

     users [file]

With no file argument, users extracts its information from the file ‘/etc/utmp’. If a file argument is given, users uses that file instead. A common choice is ‘/etc/wtmp’.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

20.6 who: Print who is currently logged in

who prints information about users who are currently logged on. Synopsis:

     who [option] [file] [am i]

If given no non-option arguments, who prints the following information for each user currently logged on: login name, terminal line, login time, and remote hostname or X display.

If given one non-option argument, who uses that instead of ‘/etc/utmp’ as the name of the file containing the record of users logged on. ‘/etc/wtmp’ is commonly given as an argument to who to look at who has previously logged on.

If given two non-option arguments, who prints only the entry for the user running it (determined from its standard input), preceded by the hostname. Traditionally, the two arguments given are ‘am i’, as in ‘who am i’.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-m’ Same as ‘who am i’.

‘-q’

‘--count’ Print only the login names and the number of users logged on. Overrides all other options.

‘-s’ Ignored; for compatibility with other versions of who.

‘-i’
‘-u’

‘--idle’ After the login time, print the number of hours and minutes that the user has been idle. ‘.’ means the user was active in last minute. ‘old’ means the user was idle for more than 24 hours.

‘-l’
‘--lookup’

Attempt to canonicalize hostnames found in utmp through a DNS lookup. This is not the default because it can cause significant delays on systems with automatic dial-up internet access.

‘-H’
‘--heading’

Print a line of column headings.

‘-w’
‘-T’
‘--mesg’
‘--message’
‘--writable’

After each login name print a character indicating the user’s message status: ‘+’ allowing write messages ‘-’ disallowing write messages ‘?’ cannot find terminal device

21 System context

This section describes commands that print or change system-wide information.

21.1 date: Print or set system date and time

Synopses:

     date [option]... [+format]
     date [-u|--utc|--universal] [ MMDDhhmm[[CC]YY][.ss] ]

Invoking date with no format argument is equivalent to invoking ‘date '+%a %b %e

%H:%M:%S %Z %Y'’.

If given an argument that starts with a ‘+’, date prints the current time and date (or the time and date specified by the --date option, see below) in the format defined by that argument, which is the same as in the strftime function. Except for directives, which start with ‘%’, characters in the format string are printed unchanged. The directives are described below.

21.1.1 Time directives

date directives related to times. ‘%H’ hour (00…23) ‘%I’ hour (01…12) ‘%k’ hour ( 0…23) ‘%l’ hour ( 1…12) ‘%M’ minute (00…59) ‘%N’ nanoseconds (000000000…999999999) ‘%p’ locale’s upper case ‘AM’ or ‘PM’ (blank in many locales) ‘%P’ locale’s lower case ‘am’ or ‘pm’ (blank in many locales) ‘%r’ time, 12-hour (hh:mm:ss [AP]M) ‘%R’ time, 24-hour (hh:mm). Same as %H:%M. ‘%s’ seconds since the epoch, i.e., 1 January 1970 00:00:00 UTC (a GNU extension).

Note that this value is the number of seconds between the epoch and the current date as defined by the localtime system call. It isn’t changed by the ‘--date’ option.

‘%S’ second (00…60). The range is [00…60], and not [00…59], in order to ac-

commodate the occasional positive leap second.

‘%T’ time, 24-hour (hh:mm:ss) ‘%X’ locale’s time representation (%H:%M:%S)

‘%z’ RFC-822 style numeric time zone (e.g., -0600 or +0100), or nothing if no time zone is determinable. This value reflects the current time zone. It isn’t changed by the ‘--date’ option.

‘%Z’ time zone (e.g., EDT), or nothing if no time zone is determinable. Note that this value reflects the current time zone. It isn’t changed by the ‘--date’ option.

21.1.2 Date directives

date directives related to dates. ‘%a’ locale’s abbreviated weekday name (Sun…Sat) ‘%A’ locale’s full weekday name, variable length (Sunday…Saturday) ‘%b’ locale’s abbreviated month name (Jan…Dec) ‘%B’ locale’s full month name, variable length (January…December) ‘%c’ locale’s date and time (Sat Nov 04 12:02:33 EST 1989) ‘%C’ century (year divided by 100 and truncated to an integer) (00…99) ‘%d’ day of month (01…31) ‘%D’ date (mm/dd/yy) ‘%e’ blank-padded day of month (1…31) ‘%F’ the ISO 8601 standard date format: %Y-%m-%d. This is the preferred form for

all uses.

‘%g’ The year corresponding to the ISO week number, but without the century (range 00 through 99). This has the same format and value as %y, except that if the ISO week number (see %V) belongs to the previous or next year, that year is used instead.

‘%G’ The year corresponding to the ISO week number. This has the same format and value as %Y, except that if the ISO week number (see %V) belongs to the previous or next year, that year is used instead.

‘%h’ same as %b ‘%j’ day of year (001…366) ‘%m’ month (01…12) ‘%u’ day of week (1…7) with 1 corresponding to Monday ‘%U’ week number of year with Sunday as first day of week (00…53). Days in a

new year preceding the first Sunday are in week zero.

‘%V’ week number of year with Monday as first day of the week as a decimal (01…53). If the week containing January 1 has four or more days in the new year, then it is considered week 1; otherwise, it is week 53 of the previous year, and the next week is week 1. (See the iso 8601 standard.)

‘%w’ day of week (0…6) with 0 corresponding to Sunday ‘%W’ week number of year with Monday as first day of week (00…53). Days in a

new year preceding the first Monday are in week zero.

‘%x’ locale’s date representation (mm/dd/yy)

‘%y’ last two digits of year (00…99)

‘%Y’ year (1970. . ..)

21.1.3 Literal directives

date directives that produce literal strings.

‘%%’ a literal %

‘%n’ a newline

‘%t’ a horizontal tab

21.1.4 Padding

By default, date pads numeric fields with zeroes, so that, for example, numeric months are always output as two digits. GNU date recognizes the following numeric modifiers between the ‘%’ and the directive.

‘-’ (hyphen) do not pad the field; useful if the output is intended for human consumption.

‘_’ (underscore) pad the field with spaces; useful if you need a fixed number of characters in the output, but zeroes are too distracting.

These are GNU extensions.

Here is an example illustrating the differences:

     date +%d/%m -d "Feb 1"
     ⇒ 01/02
     date +%-d/%-m -d "Feb 1"
     ⇒ 1/2
     date +%_d/%_m -d "Feb 1"
     ⇒ 1/ 2

21.1.5 Setting the time

If given an argument that does not start with ‘+’, date sets the system clock to the time and date specified by that argument (as described below). You must have appropriate privileges to set the system clock. The ‘--date’ and ‘--set’ options may not be used with such an argument. The ‘--universal’ option may be used with such an argument to indicate that the specified time and date are relative to Coordinated Universal Time rather than to the local time zone.

The argument must consist entirely of digits, which have the following meaning:

‘MM’ month ‘DD’ day within month

‘hh’ hour

‘mm’ minute

‘CC’ first two digits of year (optional)

‘YY’ last two digits of year (optional)

‘ss’ second (optional)

The ‘--set’ option also sets the system clock; see the next section.

21.1.6 Options for date

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-d datestr’
‘--date=datestr’

Display the time and date specified in datestr instead of the current time and date. datestr can be in almost any common format. It can contain month names, time zones, ‘am’ and ‘pm’, ‘yesterday’, ‘ago’, ‘next’, etc. See Chapter 27 [Date input formats], page 141.

‘-f datefile’
‘--file=datefile’

Parse each line in datefile as with ‘-d’ and display the resulting time and date. If datefile is ‘-’, use standard input. This is useful when you have many dates to process, because the system overhead of starting up the date executable many times can be considerable.

‘-I timespec’
‘--iso-8601[=timespec]’

Display the date using the iso 8601 format, ‘%Y-%m-%d’. The argument timespec specifies the number of additional terms of the time to include. It can be one of the following:

‘auto’ The default behavior: print just the date.

‘hours’ Append the hour of the day to the date.

‘minutes’ Append the hours and minutes.

‘seconds’ Append the hours, minutes, and seconds.

If showing any time terms, then include the time zone using the format ‘%z’. If timespec is omitted with ‘--iso-8601’, the default is ‘auto’. On older systems, gnu date instead supports an obsolete option ‘-I[timespec]’, where timespec defaults to ‘auto’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow ‘-I’ without an argument; use ‘--iso-8601’ instead.

‘-R’
‘--rfc-822’

Display the time and date using the RFC-822-conforming format, ‘%a, %_d %b

          %Y %H:%M:%S %z’.
‘-r file’
‘--reference=file’

Display the time and date reference according to the last modification time of file, instead of the current time and date.

‘-s datestr’
‘--set=datestr’

Set the time and date to datestr. See ‘-d’ above.

‘-u’
‘--utc’
‘--universal’

Use Coordinated Universal Time (utc) by operating as if the TZ environment variable were set to the string ‘UTC0’. Normally, date operates in the time zone indicated by TZ, or the system default if TZ is not set. Coordinated Universal Time is often called “Greenwich Mean Time” (gmt) for historical reasons.

21.1.7 Examples of date

Here are a few examples. Also see the documentation for the ‘-d’ option in the previous section.

• To print the date of the day before yesterday:

         date --date='2 days ago'

• To print the date of the day three months and one day hence:

         date --date='3 months 1 day'

• To print the day of year of Christmas in the current year:

         date --date='25 Dec' +%j

• To print the current full month name and the day of the month:

         date '+%B %d'

But this may not be what you want because for the first nine days of the month, the ‘%d’ expands to a zero-padded two-digit field, for example ‘date -d 1may '+%B %d'’ will print ‘May 01’.

• To print a date without the leading zero for one-digit days of the month, you can use

the (GNU extension) - modifier to suppress the padding altogether.

         date -d 1may '+%B %-d

• To print the current date and time in the format required by many non-GNU versions

of date when setting the system clock:

         date +%m%d%H%M%Y.%S

• To set the system clock forward by two minutes:

         date --set='+2 minutes'

• To print the date in the format specified by RFC-822, use ‘date --rfc’. I just did and

saw this:

     Mon, 25 Mar 1996 23:34:17 -0600

• To convert a date string to the number of seconds since the epoch (which is 1970-01-01 00:00:00 UTC), use the ‘--date’ option with the ‘%s’ format. That can be useful in sorting and/or graphing and/or comparing data by date. The following command outputs the number of the seconds since the epoch for the time two minutes after the epoch:

     date --date='1970-01-01 00:02:00 +0000' +%s
     120

If you do not specify time zone information in the date string, date uses your computer’s idea of the time zone when interpreting the string. For example, if your computer’s time zone is that of Cambridge, Massachusetts, which was then 5 hours (i.e., 18,000 seconds) behind UTC:

     # local time zone used
     date --date='1970-01-01 00:02:00' +%s
     18120

• If you’re sorting or graphing dated data, your raw date values may be represented as seconds since the epoch. But few people can look at the date ‘946684800’ and casually note “Oh, that’s the first second of the year 2000 in Greenwich, England.”

     date --date='2000-01-01 UTC' +%s
     946684800

To convert such an unwieldy number of seconds back to a more readable form, use a command like this:

     # local time zone used
     date -d '1970-01-01 UTC 946684800 seconds' +"%Y-%m-%d %T %z"
     1999-12-31 19:00:00 -0500

21.2 uname: Print system information

uname prints information about the machine and operating system it is run on. If no options are given, uname acts as if the -s option were given. Synopsis:

 uname [option]...

If multiple options or -a are given, the selected information is printed in this order:

 kernel-name nodename kernel-release kernel-version machine processor hardware-
 platform operating-system

The information may contain internal spaces, so such output cannot be parsed reliably.

In the following example, release is ‘2.2.18ss.e820-bda652a #4 SMP Tue Jun 5 11:24:08
PDT 2001’:
 uname -a
 ⇒ Linux dum 2.2.18ss.e820-bda652a #4 SMP Tue Jun 5 11:24:08 PDT 2001 i686 unknown unknown GNU/Linux

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-a’

‘--all’ Print all of the below information.

‘-i’
‘--hardware-platform’

Print the hardware platform name (sometimes called the hardware implementation).

‘-m’
‘--machine’

Print the machine hardware name (sometimes called the hardware class).

‘-n’
‘--nodename’

Print the network node hostname.

‘-p’
‘--processor’

Print the processor type (sometimes called the instruction set architecture or ISA).

‘-o’
‘--operating-system’

Print the name of the operating system.

‘-r’
‘--kernel-release’

Print the kernel release.

‘-s’
‘--kernel-name’

Print the kernel name.

‘-v’
‘--kernel-version’

Print the kernel version.

21.3 hostname: Print or set system name

With no arguments, hostname prints the name of the current host system. With one argument, it sets the current host name to the specified string. You must have appropriate privileges to set the host name. Synopsis:

     hostname [name]

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

21.4 hostid: Print numeric host identifier.

hostid prints the numeric identifier of the current host in hexadecimal. This command accepts no arguments. The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

For example, here’s what it prints on one system I use:

     $ hostid
     1bac013d

On that system, the 32-bit quantity happens to be closely related to the system’s Internet address, but that isn’t always the case.

22 Modified command invocation

This section describes commands that run other commands in some context different than the current one: a modified environment, as a different user, etc.

22.1 chroot: Run a command with a different root directory

chroot runs a command with a specified root directory. On many systems, only the super-user can do this. Synopses:

     chroot newroot [command [args]...]
     chroot option

Ordinarily, filenames are looked up starting at the root of the directory structure, i.e., ‘/’. chroot changes the root to the directory newroot (which must exist) and then runs command with optional args. If command is not specified, the default is the value of the SHELL environment variable or /bin/sh if not set, invoked with the ‘-i’ option.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

Here are a few tips to help avoid common problems in using chroot. To start with a simple example, make command refer to a statically linked binary. If you were to use a dynamically linked executable, then you’d have to arrange to have the shared libraries in the right place under your new root directory.

For example, if you create a statically linked ‘ls’ executable, and put it in /tmp/empty, you can run this command as root:

     $ chroot /tmp/empty /ls -Rl /

Then you’ll see output like this:

     /:
     total 1023
     -rwxr-xr-x    1 0        0         1041745 Aug 16 11:17 ls

If you want to use a dynamically linked executable, say bash, then first run ‘ldd bash’ to see what shared objects it needs. Then, in addition to copying the actual binary, also copy the listed files to the required positions under your intended new root directory. Finally, if the executable requires any other files (e.g., data, state, device files), copy them into place, too.

22.2 env: Run a command in a modified environment

env runs a command with a modified environment. Synopses:

     env [option]... [name=value]... [command [args]...]
     env

Arguments of the form ‘variable=value’ set the environment variable variable to value value. value may be empty (‘variable=’). Setting a variable to an empty value is different from unsetting it.

The first remaining argument specifies the program name to invoke; it is searched for according to the PATH environment variable. Any remaining arguments are passed as arguments to that program.

If no command name is specified following the environment specifications, the resulting environment is printed. This is like specifying a command name of printenv.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-u name’
‘--unset=name’

Remove variable name from the environment, if it was in the environment.

‘-’

‘-i’
‘--ignore-environment’

Start with an empty environment, ignoring the inherited environment.

22.3 nice: Run a command with modified scheduling priority

nice prints or modifies the scheduling priority of a job. Synopsis:

     nice [option]... [command [arg]...]

If no arguments are given, nice prints the current scheduling priority, which it inherited. Otherwise, nice runs the given command with its scheduling priority adjusted. If no adjustment is given, the priority of the command is incremented by 10. You must have appropriate privileges to specify a negative adjustment. The priority can be adjusted by nice over the range of -20 (the highest priority) to 19 (the lowest).

Because most shells have a built-in command by the same name, using the unadorned command name in a script or interactively may get you different functionality than that described here.

The program accepts the following option. Also see Chapter 2 [Common options], page 2.

‘-n adjustment’
‘--adjustment=adjustment’

Add adjustment instead of 10 to the command’s priority.

On older systems, nice supports an obsolete option ‘-adjustment’. posix 1003.1-2001 (see Section 2.5 [Standards conformance], page 6) does not allow this; use ‘-n adjustment’ instead.

22.4 nohup: Run a command immune to hangups

nohup runs the given command with hangup signals ignored, so that the command can continue running in the background after you log out. Synopsis:

     nohup command [arg]...

If standard output is a terminal, it is redirected so that it is appended to the file ‘nohup.out’; if that cannot be written to, it is appended to the file ‘$HOME/nohup.out’. If that cannot be written to, the command is not run.

If nohup creates either ‘nohup.out’ or ‘$HOME/nohup.out’, it creates it with no “group” or “other” access permissions. It does not change the permissions if the output file already existed.

If standard error is a terminal, it is redirected to the same file descriptor as the standard output.

nohup does not automatically put the command it runs in the background; you must do that explicitly, by ending the command line with an ‘&’. Also, nohup does not change the scheduling priority of command; use nice for that, e.g., ‘nohup nice command’.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

Exit status:

126 if command was found but could not be invoked 127 if nohup itself failed or if command could not be found the exit status of command otherwise

22.5 su: Run a command with substitute user and group id

su allows one user to temporarily become another user. It runs a command (often an interactive shell) with the real and effective user id, group id, and supplemental groups of a given user. Synopsis:

     su [option]... [user [arg]...]

If no user is given, the default is root, the super-user. The shell to use is taken from user’s passwd entry, or ‘/bin/sh’ if none is specified there. If user has a password, su prompts for the password unless run by a user with effective user id of zero (the superuser).

By default, su does not change the current directory. It sets the environment variables HOME and SHELL from the password entry for user, and if user is not the super-user, sets USER and LOGNAME to user. By default, the shell is not a login shell.

Any additional args are passed as additional arguments to the shell.

GNU su does not treat ‘/bin/sh’ or any other shells specially (e.g., by setting argv[0] to ‘-su’, passing -c only to certain shells, etc.).

su can optionally be compiled to use syslog to report failed, and optionally successful, su attempts. (If the system supports syslog.) However, GNU su does not check if the user is a member of the wheel group; see below.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-c command’
‘--command=command’

Pass command, a single command line to run, to the shell with a -c option instead of starting an interactive shell.

‘-f’

‘--fast’ Pass the -f option to the shell. This probably only makes sense if the shell run is csh or tcsh, for which the -f option prevents reading the startup file (‘.cshrc’). With Bourne-like shells, the -f option disables file name pattern expansion (globbing), which is not likely to be useful.

‘-’

‘-l’

‘--login’ Make the shell a login shell. This means the following. Unset all environment variables except TERM, HOME, and SHELL (which are set as described above), and USER and LOGNAME (which are set, even for the super-user, as described above), and set PATH to a compiled-in default value. Change to user’s home directory. Prepend ‘-’ to the shell’s name, intended to make it read its login startup file(s).

‘-m’
‘-p’
‘--preserve-environment’

Do not change the environment variables HOME, USER, LOGNAME, or SHELL. Run the shell given in the environment variable SHELL instead of the shell from user’s passwd entry, unless the user running su is not the superuser and user’s shell is restricted. A restricted shell is one that is not listed in the file ‘/etc/shells’, or in a compiled-in list if that file does not exist. Parts of what this option does can be overridden by --login and --shell.

‘-s shell’
‘--shell=shell’

Run shell instead of the shell from user’s passwd entry, unless the user running su is not the superuser and user’s shell is restricted (see ‘-m’ just above).

Why GNU su does not support the ‘wheel’ group

(This section is by Richard Stallman.)

Sometimes a few of the users try to hold total power over all the rest. For example, in 1984, a few users at the MIT AI lab decided to seize power by changing the operator password on the Twenex system and keeping it secret from everyone else. (I was able to thwart this coup and give power back to the users by patching the kernel, but I wouldn’t know how to do that in Unix.)

However, occasionally the rulers do tell someone. Under the usual su mechanism, once someone learns the root password who sympathizes with the ordinary users, he or she can tell the rest. The “wheel group” feature would make this impossible, and thus cement the power of the rulers.

I’m on the side of the masses, not that of the rulers. If you are used to supporting the bosses and sysadmins in whatever they do, you might find this idea strange at first.

23 Process control

23.1 kill: Send a signal to processes

The kill command sends a signal to processes, causing them to terminate or otherwise act upon receiving the signal in some way. Alternatively, it lists information about signals. Synopses:

     kill [-s signal | --signal signal | -signal] pid...
     kill [-l | --list | -t | --table] [signal]...

The first form of the kill command sends a signal to all pid arguments. The default signal to send if none is specified is ‘TERM’. The special signal number ‘0’ does not denote a valid signal, but can be used to test whether the pid arguments specify processes to which a signal could be sent.

If pid is positive, the signal is sent to the process with the process id pid. If pid is zero, the signal is sent to all processes in the process group of the current process. If pid is -1, the signal is sent to all processes for which the user has permission to send a signal. If pid is less than -1, the signal is sent to all processes in the process group that equals the absolute value of pid.

If pid is not positive, a system-dependent set of system processes is excluded from the list of processes to which the signal is sent.

If a negative PID argument is desired as the first one, either a signal must be specified as well, or the option parsing must be interrupted with ‘–’ before the first pid argument. The following three commands are equivalent:

     kill -15 -1
     kill -TERM -1
     kill -- -1

The first form of the kill command succeeds if every pid argument specifies at least one process that the signal was sent to.

The second form of the kill command lists signal information. Either the ‘-l’ or ‘--list’ option, or the ‘-t’ or ‘--table’ option must be specified. Without any signal argument, all supported signals are listed. The output of ‘-l’ or ‘--list’ is a list of the signal names, one per line; if signal is already a name, the signal number is printed instead. The output of ‘-t’ or ‘--table’ is a table of signal numbers, names, and descriptions. This form of the kill command succeeds if all signal arguments are valid and if there is no output error.

The kill command also supports the ‘--help’ and ‘--version’ options. See Chapter 2 [Common options], page 2.

A signal may be a signal name like ‘HUP’, or a signal number like ‘1’, or an exit status of a process terminated by the signal. A signal name can be given in canonical form or prefixed by ‘SIG’. The case of the letters is ignored, except for the ‘-signal’ option which must use upper case to avoid ambiguity with lower case option letters. The following signal names and numbers are supported on all posix compliant systems:

‘HUP’ 1. Hangup.

‘INT’ 2. Terminal interrupt. ‘QUIT’ 3. Terminal quit. ‘ABRT’ 6. Process abort. ‘KILL’ 9. Kill (cannot be caught or ignored). ‘ALRM’ 14. Alarm Clock. ‘TERM’ 15. Termination. Other supported signal names have system-dependent corresponding numbers. All systems conforming to posix 1003.1-2001 also support the following signals: ‘BUS’ Access to an undefined portion of a memory object. ‘CHLD’ Child process terminated, stopped, or continued. ‘CONT’ Continue executing, if stopped. ‘FPE’ Erroneous arithmetic operation. ‘ILL’ Illegal Instruction. ‘PIPE’ Write on a pipe with no one to read it. ‘SEGV’ Invalid memory reference. ‘STOP’ Stop executing (cannot be caught or ignored). ‘TSTP’ Terminal stop. ‘TTIN’ Background process attempting read. ‘TTOU’ Background process attempting write. ‘URG’ High bandwidth data is available at a socket. ‘USR1’ User-defined signal 1. ‘USR2’ User-defined signal 2. posix 1003.1-2001 systems that support the xsi extension also support the following signals: ‘POLL’ Pollable event. ‘PROF’ Profiling timer expired. ‘SYS’ Bad system call. ‘TRAP’ Trace/breakpoint trap. ‘VTALRM’ Virtual timer expired. ‘XCPU’ CPU time limit exceeded. ‘XFSZ’ File size limit exceeded. posix 1003.1-2001 systems that support the xrt extension also support at least eight realtime signals called ‘RTMIN’, ‘RTMIN+1’, …, ‘RTMAX-1’, ‘RTMAX’.

24 Delaying

24.1 sleep: Delay for a specified time

sleep pauses for an amount of time specified by the sum of the values of the command line arguments. Synopsis:

     sleep number[smhd]...

Each argument is a number followed by an optional unit; the default is seconds. The units are: ‘s’ seconds ‘m’ minutes ‘h’ hours ‘d’ days

Historical implementations of sleep have required that number be an integer. However, GNU sleep accepts arbitrary floating point numbers.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

25 Numeric operations

These programs do numerically-related operations.

25.1 factor: Print prime factors

factor prints prime factors. Synopses:

     factor [number]...
     factor option

If no number is specified on the command line, factor reads numbers from standard input, delimited by newlines, tabs, or spaces.

The only options are ‘--help’ and ‘--version’. See Chapter 2 [Common options], page 2.

The algorithm it uses is not very sophisticated, so for some inputs factor runs for a long time. The hardest numbers to factor are the products of large primes. Factoring the product of the two largest 32-bit prime numbers takes over 10 minutes of CPU time on a 400MHz Pentium II.

     $ p=`echo '4294967279 * 4294967291'|bc`
     $ factor $p
     18446743979220271189: 4294967279 4294967291

In contrast, factor factors the largest 64-bit number in just over a tenth of a second:

     $ factor `echo '2^64-1'|bc`
     18446744073709551615: 3 5 17 257 641 65537 6700417

25.2 seq: Print numeric sequences

seq prints a sequence of numbers to standard output. Synopses:

     seq [option]... [first [increment]] last...

seq prints the numbers from first to last by increment. By default, first and increment are both 1, and each number is printed on its own line. All numbers can be reals, not just integers.

The program accepts the following options. Also see Chapter 2 [Common options], page 2.

‘-f format’
‘--format=format’

Print all numbers using format; default ‘%g’. format must contain exactly one of the floating point output formats ‘%e’, ‘%f’, or ‘%g’.

‘-s string’
‘--separator=string’

Separate numbers with string; default is a newline. The output always terminates with a newline.

‘-w’
‘--equal-width’

Print all numbers with the same width, by padding with leading zeroes. (To have other kinds of padding, use ‘--format’).

If you want to use seq to print sequences of large integer values, don’t use the default ‘%g’ format since it can result in loss of precision:

$ seq 1000000 1000001
1e+06
1e+06

Instead, you can use the format, ‘%1.f’, to print large decimal numbers with no exponent and no decimal point.

$ seq --format=%1.f 1000000 1000001
1000000
1000001

If you want hexadecimal output, you can use printf to perform the conversion:

$ printf %x'\n' `seq -f %1.f 1048575 1024 1050623`
fffff
1003ff
1007ff

For very long lists of numbers, use xargs to avoid system limitations on the length of an argument list:

$ seq -f %1.f 1000000 | xargs printf %x'\n' | tail -n 3
f423e
f423f
f4240

To generate octal output, use the printf %o format instead of %x. Note however that using printf works only for numbers smaller than 2^32:

$ printf "%x\n" `seq -f %1.f 4294967295 4294967296`
ffffffff
bash: printf: 4294967296: Numerical result out of range

On most systems, seq can produce whole-number output for values up to 2^53, so here’s a more general approach to base conversion that also happens to be more robust for such large numbers. It works by using bc and setting its output radix variable, obase, to ‘16’ in this case to produce hexadecimal output.

$ (echo obase=16; seq -f %1.f 4294967295 4294967296)|bc
FFFFFFFF
100000000

Be careful when using seq with a fractional increment, otherwise you may see surprising results. Most people would expect to see 0.3 printed as the last number in this example:

$ seq -s' ' 0 .1 .3
0 0.1 0.2

But that doesn’t happen on most systems because seq is implemented using binary floating point arithmetic (via the C double type) – which means some decimal numbers like .1 cannot be represented exactly. That in turn means some nonintuitive conditions like .1 * 3 > .3 will end up being true.

To work around that in the above example, use a slightly larger number as the last value:

     $ seq -s' ' 0 .1 .31
     0 0.1 0.2 0.3

In general, when using an increment with a fractional part, where (last - first) / increment is (mathematically) a whole number, specify a slightly larger (or smaller, if increment is negative) value for last to ensure that last is the final value printed by seq.

26 File permissions

Each file has a set of permissions that control the kinds of access that users have to that file. The permissions for a file are also called its access mode. They can be represented either in symbolic form or as an octal number.

26.1 Structure of File Permissions

There are three kinds of permissions that a user can have for a file:

1. permission to read the file. For directories, this means permission to list the contents of the directory.

2. permission to write to (change) the file. For directories, this means permission to create and remove files in the directory.

3. permission to execute the file (run it as a program). For directories, this means permission to access files in the directory.

There are three categories of users who may have different permissions to perform any of the above operations on a file:

1. the file’s owner;

2. other users who are in the file’s group;

3. everyone else.

Files are given an owner and group when they are created. Usually the owner is the current user and the group is the group of the directory the file is in, but this varies with the operating system, the filesystem the file is created on, and the way the file is created. You can change the owner and group of a file by using the chown and chgrp commands.

In addition to the three sets of three permissions listed above, a file’s permissions have three special components, which affect only executable files (programs) and, on some systems, directories:

1. set the process’s effective user ID to that of the file upon execution (called the setuid bit). No effect on directories.

2. set the process’s effective group ID to that of the file upon execution (called the setgid bit). For directories on some systems, put files created in the directory into the same group as the directory, no matter what group the user who creates them is in.

3. save the program’s text image on the swap device so it will load more quickly when run (called the sticky bit). For directories on some systems, prevent users from removing or renaming a file in a directory unless they own the file or the directory; this is called the restricted deletion flag for the directory.

26.2 Symbolic Modes

Symbolic modes represent changes to files’ permissions as operations on single-character symbols. They allow you to modify either all or selected parts of files’ permissions, optionally based on their previous values, and perhaps on the current umask as well (see Section 26.2.6 [Umask and Protection], page 139).

The format of symbolic modes is:

     [ugoa...][[+-=][rwxXstugo...]...][,...]

The following sections describe the operators and other details of symbolic modes.

26.2.1 Setting Permissions

The basic symbolic operations on a file’s permissions are adding, removing, and setting the permission that certain users have to read, write, and execute the file. These operations have the following format:

     users operation permissions

The spaces between the three parts above are shown for readability only; symbolic modes cannot contain spaces.

The users part tells which users’ access to the file is changed. It consists of one or more of the following letters (or it can be empty; see Section 26.2.6 [Umask and Protection], page 139, for a description of what happens then). When more than one of these letters is given, the order that they are in does not matter.

u the user who owns the file;

g other users who are in the file’s group;

o all other users;

a all users; the same as ‘ugo’.

The operation part tells how to change the affected users’ access to the file, and is one of the following symbols:

+ to add the permissions to whatever permissions the users already have for the file;

- to remove the permissions from whatever permissions the users already have for the file;

= to make the permissions the only permissions that the users have for the file.

The permissions part tells what kind of access to the file should be changed; it is zero or more of the following letters. As with the users part, the order does not matter when more than one letter is given. Omitting the permissions part is useful only with the ‘=’ operation, where it gives the specified users no access at all to the file.

r the permission the users have to read the file;

w the permission the users have to write to the file;

x the permission the users have to execute the file.

For example, to give everyone permission to read and write a file, but not to execute it, use:

     a=rw

To remove write permission for from all users other than the file’s owner, use:

     go-w

The above command does not affect the access that the owner of the file has to it, nor does it affect whether other users can read or execute the file.

To give everyone except a file’s owner no permission to do anything with that file, use the mode below. Other users could still remove the file, if they have write permission on the directory it is in.

     go=

Another way to specify the same thing is:

     og-rxw

26.2.2 Copying Existing Permissions

You can base a file’s permissions on its existing permissions. To do this, instead of using ‘r’, ‘w’, or ‘x’ after the operator, you use the letter ‘u’, ‘g’, or ‘o’. For example, the mode

     o+g

adds the permissions for users who are in a file’s group to the permissions that other users have for the file. Thus, if the file started out as mode 664 (‘rw-rw-r--’), the above mode would change it to mode 666 (‘rw-rw-rw-’). If the file had started out as mode 741 (‘rwxr----x’), the above mode would change it to mode 745 (‘rwxr--r-x’). The ‘-’ and ‘=’ operations work analogously.

26.2.3 Changing Special Permissions

In addition to changing a file’s read, write, and execute permissions, you can change its special permissions. See Section 26.1 [Mode Structure], page 135, for a summary of these permissions.

To change a file’s permission to set the user ID on execution, use ‘u’ in the users part of the symbolic mode and ‘s’ in the permissions part.

To change a file’s permission to set the group ID on execution, use ‘g’ in the users part of the symbolic mode and ‘s’ in the permissions part.

To change a file’s permission to stay permanently on the swap device, use ‘o’ in the users part of the symbolic mode and ‘t’ in the permissions part.

For example, to add set user ID permission to a program, you can use the mode:

     u+s

To remove both set user ID and set group ID permission from it, you can use the mode:

     ug-s

To cause a program to be saved on the swap device, you can use the mode:

     o+t

Remember that the special permissions only affect files that are executable, plus, on some systems, directories (on which they have different meanings; see Section 26.1 [Mode Structure], page 135). Also, the combinations ‘u+t’, ‘g+t’, and ‘o+s’ have no effect.

The ‘=’ operator is not very useful with special permissions; for example, the mode:

     o=t

does cause the file to be saved on the swap device, but it also removes all read, write, and execute permissions that users not in the file’s group might have had for it.

26.2.4 Conditional Executability

There is one more special type of symbolic permission: if you use ‘X’ instead of ‘x’, execute permission is affected only if the file already had execute permission or is a directory. It affects directories’ execute permission even if they did not initially have any execute permissions set.

For example, this mode:

     a+X

gives all users permission to execute files (or search directories) if anyone could before.

26.2.5 Making Multiple Changes

The format of symbolic modes is actually more complex than described above (see Section 26.2.1 [Setting Permissions], page 136). It provides two ways to make multiple changes to files’ permissions.

The first way is to specify multiple operation and permissions parts after a users part in the symbolic mode.

For example, the mode:

     og+rX-w

gives users other than the owner of the file read permission and, if it is a directory or if someone already had execute permission to it, gives them execute permission; and it also denies them write permission to the file. It does not affect the permission that the owner of the file has for it. The above mode is equivalent to the two modes:

     og+rX
     og-w

The second way to make multiple changes is to specify more than one simple symbolic mode, separated by commas. For example, the mode:

     a+r,go-w

gives everyone permission to read the file and removes write permission on it for all users except its owner. Another example:

     u=rwx,g=rx,o=

sets all of the non-special permissions for the file explicitly. (It gives users who are not in the file’s group no permission at all for it.)

The two methods can be combined. The mode:

     a+r,g+x-w

gives all users permission to read the file, and gives users who are in the file’s group permission to execute it, as well, but not permission to write to it. The above mode could be written in several different ways; another is:

     u+r,g+rx,o+r,g-w

26.2.6 The Umask and Protection

If the users part of a symbolic mode is omitted, it defaults to ‘a’ (affect all users), except that any permissions that are set in the system variable umask are not affected. The value of umask can be set using the umask command. Its default value varies from system to system.

Omitting the users part of a symbolic mode is generally not useful with operations other than ‘+’. It is useful with ‘+’ because it allows you to use umask as an easily customizable protection against giving away more permission to files than you intended to.

As an example, if umask has the value 2, which removes write permission for users who are not in the file’s group, then the mode:

     +w

adds permission to write to the file to its owner and to other users who are in the file’s group, but not to other users. In contrast, the mode:

     a+w

ignores umask, and does give write permission for the file to all users.

26.3 Numeric Modes

File permissions are stored internally as integers. As an alternative to giving a symbolic mode, you can give an octal (base 8) number that corresponds to the internal representation of the new mode. This number is always interpreted in octal; you do not have to add a leading 0, as you do in C. Mode 0055 is the same as mode 55.

A numeric mode is usually shorter than the corresponding symbolic mode, but it is limited in that it cannot take into account a file’s previous permissions; it can only set them absolutely.

On most systems, the permissions granted to the user, to other users in the file’s group, and to other users not in the file’s group are each stored as three bits, which are represented as one octal digit. The three special permissions are also each stored as one bit, and they are as a group represented as another octal digit. Here is how the bits are arranged, starting with the lowest valued bit:

     Value in  Corresponding
     Mode      Permission
               Other users not in the file's group:
        1      Execute
        2      Write
        4      Read
               Other users in the file's group:
       10      Execute
       20      Write
       40      Read

               The file's owner:
      100      Execute
      200      Write
      400      Read
               Special permissions:
     1000      Save text image on swap device
     2000      Set group ID on execution
     4000      Set user ID on execution

For example, numeric mode 4755 corresponds to symbolic mode ‘u=rwxs,go=rx’, and numeric mode 664 corresponds to symbolic mode ‘ug=rw,o=r’. Numeric mode 0 corresponds to symbolic mode ‘ugo=’.

27 Date input formats

First, a quote:

Our units of temporal measurement, from seconds on up to months, are so complicated, asymmetrical and disjunctive so as to make coherent mental reckoning in time all but impossible. Indeed, had some tyrannical god contrived to enslave our minds to time, to make it all but impossible for us to escape subjection to sodden routines and unpleasant surprises, he could hardly have done better than handing down our present system. It is like a set of trapezoidal building blocks, with no vertical or horizontal surfaces, like a language in which the simplest thought demands ornate constructions, useless particles and lengthy circumlocutions. Unlike the more successful patterns of language and science, which enable us to face experience boldly or at least level-headedly, our system of temporal calculation silently and persistently encourages our terror of time. … It is as though architects had to measure length in feet, width in meters and height in ells; as though basic instruction manuals demanded a knowledge of five different languages. It is no wonder then that we often look into our own immediate past or future, last Tuesday or a week from Sunday, with feelings of helpless confusion. …

— Robert Grudin, Time and the Art of Living. This section describes the textual date representations that gnu programs accept. These

are the strings you, as a user, can supply as arguments to the various programs. The C interface (via the getdate function) is not described here.

Although the date syntax here can represent any possible time since the year zero, computer integers often cannot represent such a wide range of time. On posix systems, the clock starts at 1970-01-01 00:00:00 utc: posix does not require support for times before the posix Epoch and times far in the future. Traditional Unix systems have 32-bit signed time_t and can represent times from 1901-12-13 20:45:52 through 2038-01-19 03:14:07 utc. Systems with 64-bit signed time_t can represent all the times in the known lifetime of the universe.

27.1 General date syntax

A date is a string, possibly empty, containing many items separated by whitespace. The whitespace may be omitted when no ambiguity arises. The empty string means the beginning of today (i.e., midnight). Order of the items is immaterial. A date string may contain many flavors of items:

• calendar date items

• time of the day items

• time zone items

• day of the week items

• relative items

• pure numbers.

We describe each of these item types in turn, below.

A few numbers may be written out in words in most contexts. This is most useful for specifying day of the week items or relative items (see below). Here is the list: ‘first’ for 1, ‘next’ for 2, ‘third’ for 3, ‘fourth’ for 4, ‘fifth’ for 5, ‘sixth’ for 6, ‘seventh’ for 7, ‘eighth’ for 8, ‘ninth’ for 9, ‘tenth’ for 10, ‘eleventh’ for 11 and ‘twelfth’ for 12. Also, ‘last’ means exactly −1.

When a month is written this way, it is still considered to be written numerically, instead of being “spelled in full”; this changes the allowed strings.

In the current implementation, only English is supported for words and abbreviations

like ‘AM’, ‘DST’, ‘EST’, ‘first’, ‘January’, ‘Sunday’, ‘tomorrow’, and ‘year’.

The output of date is not always acceptable as a date string, not only because of the language problem, but also because there is no standard meaning for time zone items like ‘IST’. When using date to generate a date string intended to be parsed later, specify a date format that is independent of language and that does not use time zone items other than ‘UTC’ and ‘Z’. Here are some ways to do this:

$ LC_ALL=C TZ=UTC0 date
Fri Dec 15 19:48:05 UTC 2000
$ TZ=UTC0 date +"%Y-%m-%d %H:%M:%SZ"
2000-12-15 19:48:05Z
$ date --iso-8601=seconds  # a GNU extension
2000-12-15T11:48:05-0800
$ date --rfc-822  # a GNU extension
Fri, 15 Dec 2000 11:48:05 -0800
$ date +"%Y-%m-%d %H:%M:%S %z"  # %z is a GNU extension.
2000-12-15 11:48:05 -0800

Alphabetic case is completely ignored in dates. Comments may be introduced between round parentheses, as long as included parentheses are properly nested. Hyphens not followed by a digit are currently ignored. Leading zeros on numbers are ignored.

27.2 Calendar date items

A calendar date item specifies a day of the year. It is specified differently, depending on whether the month is specified numerically or literally. All these strings specify the same calendar date:

1972-09-24     # iso 8601.
72-9-24        # Assume 19xx for 69 through 99,
               # 20xx for 00 through 68.
72-09-24       # Leading zeros are ignored.
9/24/72        # Common U.S. writing.
24 September 1972
24 Sept 72     # September has a special abbreviation.
24 Sep 72      # Three-letter abbreviations always allowed.
Sep 24, 1972
24-sep-72
24sep72

The year can also be omitted. In this case, the last specified year is used, or the current year if none. For example:

     9/24
     sep 24

Here are the rules.

For numeric months, the iso 8601 format ‘year-month-day’ is allowed, where year is any positive number, month is a number between 01 and 12, and day is a number between 01 and 31. A leading zero must be present if a number is less than ten. If year is 68 or smaller, then 2000 is added to it; otherwise, if year is less than 100, then 1900 is added to it. The construct ‘month/day/year’, popular in the United States, is accepted. Also ‘month/day’, omitting the year.

Literal months may be spelled out in full: ‘January’, ‘February’, ‘March’, ‘April’, ‘May’,

‘June’, ‘July’, ‘August’, ‘September’, ‘October’, ‘November’ or ‘December’. Literal months

may be abbreviated to their first three letters, possibly followed by an abbreviating dot. It is also permitted to write ‘Sept’ instead of ‘September’.

When months are written literally, the calendar date may be given as any of the following:

     day month year
     day month
     month day year
     day-month-year

Or, omitting the year:

     month day

27.3 Time of day items

A time of day item in date strings specifies the time on a given day. Here are some examples, all of which represent the same time:

     20:02:0
     20:02
     8:02pm
     20:02-0500      # In est (U.S. Eastern Standard Time).

More generally, the time of the day may be given as ‘hour:minute:second’, where hour is a number between 0 and 23, minute is a number between 0 and 59, and second is a number between 0 and 59. Alternatively, ‘:second’ can be omitted, in which case it is taken to be zero.

If the time is followed by ‘am’ or ‘pm’ (or ‘a.m.’ or ‘p.m.’), hour is restricted to run from 1 to 12, and ‘:minute’ may be omitted (taken to be zero). ‘am’ indicates the first half of the day, ‘pm’ indicates the second half of the day. In this notation, 12 is the predecessor of 1: midnight is ‘12am’ while noon is ‘12pm’. (This is the zero-oriented interpretation of ‘12am’ and ‘12pm’, as opposed to the old tradition derived from Latin which uses ‘12m’ for noon and ‘12pm’ for midnight.)

The time may alternatively be followed by a time zone correction, expressed as ‘shhmm’, where s is ‘+’ or ‘-’, hh is a number of zone hours and mm is a number of zone minutes. When a time zone correction is given this way, it forces interpretation of the time relative to Coordinated Universal Time (utc), overriding any previous specification for the time zone or the local time zone. The minute part of the time of the day may not be elided when a time zone correction is used. This is the best way to specify a time zone correction by fractional parts of an hour.

Either ‘am’/‘pm’ or a time zone correction may be specified, but not both.

27.4 Time zone items

A time zone item specifies an international time zone, indicated by a small set of letters, e.g., ‘UTC’ or ‘Z’ for Coordinated Universal Time. Any included periods are ignored. By following a non-daylight-saving time zone by the string ‘DST’ in a separate word (that is, separated by some white space), the corresponding daylight saving time zone may be specified.

Time zone items other than ‘UTC’ and ‘Z’ are obsolescent and are not recommended, because they are ambiguous; for example, ‘EST’ has a different meaning in Australia than in the United States. Instead, it’s better to use unambiguous numeric time zone corrections like ‘-0500’, as described in the previous section.

27.5 Day of week items

The explicit mention of a day of the week will forward the date (only if necessary) to reach that day of the week in the future.

Days of the week may be spelled out in full: ‘Sunday’, ‘Monday’, ‘Tuesday’, ‘Wednesday’, ‘Thursday’, ‘Friday’ or ‘Saturday’. Days may be abbreviated to their first three letters, optionally followed by a period. The special abbreviations ‘Tues’ for ‘Tuesday’, ‘Wednes’ for ‘Wednesday’ and ‘Thur’ or ‘Thurs’ for ‘Thursday’ are also allowed.

A number may precede a day of the week item to move forward supplementary weeks. It is best used in expression like ‘third monday’. In this context, ‘last day’ or ‘next day’ is also acceptable; they move one week before or after the day that day by itself would represent.

A comma following a day of the week item is ignored.

27.6 Relative items in date strings

Relative items adjust a date (or the current date if none) forward or backward. The effects of relative items accumulate. Here are some examples:

     1 year
     1 year ago
     3 years
     2 days

The unit of time displacement may be selected by the string ‘year’ or ‘month’ for moving by whole years or months. These are fuzzy units, as years and months are not all of equal duration. More precise units are ‘fortnight’ which is worth 14 days, ‘week’ worth 7 days, ‘day’ worth 24 hours, ‘hour’ worth 60 minutes, ‘minute’ or ‘min’ worth 60 seconds, and ‘second’ or ‘sec’ worth one second. An ‘s’ suffix on these units is accepted and ignored.

The unit of time may be preceded by a multiplier, given as an optionally signed number. Unsigned numbers are taken as positively signed. No number at all implies 1 for a multiplier. Following a relative item by the string ‘ago’ is equivalent to preceding the unit by a multiplier with value −1.

The string ‘tomorrow’ is worth one day in the future (equivalent to ‘day’), the string ‘yesterday’ is worth one day in the past (equivalent to ‘day ago’).

The strings ‘now’ or ‘today’ are relative items corresponding to zero-valued time displacement, these strings come from the fact a zero-valued time displacement represents the current time when not otherwise changed by previous items. They may be used to stress other items, like in ‘12:00 today’. The string ‘this’ also has the meaning of a zero-valued time displacement, but is preferred in date strings like ‘this thursday’.

When a relative item causes the resulting date to cross a boundary where the clocks were adjusted, typically for daylight-saving time, the resulting date and time are adjusted accordingly.

27.7 Pure numbers in date strings

The precise interpretation of a pure decimal number depends on the context in the date string.

If the decimal number is of the form yyyymmdd and no other calendar date item (see Section 27.2 [Calendar date items], page 142) appears before it in the date string, then yyyy is read as the year, mm as the month number and dd as the day of the month, for the specified calendar date.

If the decimal number is of the form hhmm and no other time of day item appears before it in the date string, then hh is read as the hour of the day and mm as the minute of the hour, for the specified time of the day. mm can also be omitted.

If both a calendar date and a time of day appear to the left of a number in the date string, but no relative item, then the number overrides the year.

27.8 Authors of getdate

getdate was originally implemented by Steven M. Bellovin (smb@research.att.com) while at the University of North Carolina at Chapel Hill. The code was later tweaked by a couple of people on Usenet, then completely overhauled by Rich $alz (rsalz@bbn.com) and Jim Berets (jberets@bbn.com) in August, 1990. Various revisions for the gnu system were made by David MacKenzie, Jim Meyering, Paul Eggert and others.

This chapter was originally produced by François Pinard (pinard@iro.umontreal.ca) from the ‘getdate.y’ source code, and then edited by K. Berry (kb@cs.umb.edu).

28 Opening the Software Toolbox

This chapter originally appeared in Linux Journal, volume 1, number 2, in the What’s GNU? column. It was written by Arnold Robbins.

Toolbox Introduction

This month’s column is only peripherally related to the GNU Project, in that it describes a number of the GNU tools on your GNU/Linux system and how they might be used. What it’s really about is the “Software Tools” philosophy of program development and usage.

The software tools philosophy was an important and integral concept in the initial design and development of Unix (of which Linux and GNU are essentially clones). Unfortunately, in the modern day press of Internetworking and flashy GUIs, it seems to have fallen by the wayside. This is a shame, since it provides a powerful mental model for solving many kinds of problems.

Many people carry a Swiss Army knife around in their pants pockets (or purse). A Swiss Army knife is a handy tool to have: it has several knife blades, a screwdriver, tweezers, toothpick, nail file, corkscrew, and perhaps a number of other things on it. For the everyday, small miscellaneous jobs where you need a simple, general purpose tool, it’s just the thing.

On the other hand, an experienced carpenter doesn’t build a house using a Swiss Army knife. Instead, he has a toolbox chock full of specialized tools—a saw, a hammer, a screwdriver, a plane, and so on. And he knows exactly when and where to use each tool; you won’t catch him hammering nails with the handle of his screwdriver.

The Unix developers at Bell Labs were all professional programmers and trained computer scientists. They had found that while a one-size-fits-all program might appeal to a user because there’s only one program to use, in practice such programs are a. difficult to write, b. difficult to maintain and debug, and c. difficult to extend to meet new situations.

Instead, they felt that programs should be specialized tools. In short, each program “should do one thing well.” No more and no less. Such programs are simpler to design, write, and get right—they only do one thing.

Furthermore, they found that with the right machinery for hooking programs together, that the whole was greater than the sum of the parts. By combining several special purpose programs, you could accomplish a specific task that none of the programs was designed for, and accomplish it much more quickly and easily than if you had to write a special purpose program. We will see some (classic) examples of this further on in the column. (An important additional point was that, if necessary, take a detour and build any software tools you may need first, if you don’t already have something appropriate in the toolbox.)

I/O Redirection

Hopefully, you are familiar with the basics of I/O redirection in the shell, in particular the concepts of “standard input,” “standard output,” and “standard error”. Briefly, “standard input” is a data source, where data comes from. A program should not need to either know or care if the data source is a disk file, a keyboard, a magnetic tape, or even a punched card reader. Similarly, “standard output” is a data sink, where data goes to. The program should neither know nor care where this might be. Programs that only read their standard input, do something to the data, and then send it on, are called filters, by analogy to filters in a water pipeline.

With the Unix shell, it’s very easy to set up data pipelines:

     program_to_create_data | filter1 | .... | filterN > final.pretty.data

We start out by creating the raw data; each filter applies some successive transformation to the data, until by the time it comes out of the pipeline, it is in the desired form.

This is fine and good for standard input and standard output. Where does the standard error come in to play? Well, think about filter1 in the pipeline above. What happens if it encounters an error in the data it sees? If it writes an error message to standard output, it will just disappear down the pipeline into filter2’s input, and the user will probably never see it. So programs need a place where they can send error messages so that the user will notice them. This is standard error, and it is usually connected to your console or window, even if you have redirected standard output of your program away from your screen.

For filter programs to work together, the format of the data has to be agreed upon. The most straightforward and easiest format to use is simply lines of text. Unix data files are generally just streams of bytes, with lines delimited by the ascii lf (Line Feed) character, conventionally called a “newline” in the Unix literature. (This is '\n' if you’re a C programmer.) This is the format used by all the traditional filtering programs. (Many earlier operating systems had elaborate facilities and special purpose programs for managing binary data. Unix has always shied away from such things, under the philosophy that it’s easiest to simply be able to view and edit your data with a text editor.)

OK, enough introduction. Let’s take a look at some of the tools, and then we’ll see how to hook them together in interesting ways. In the following discussion, we will only present those command line options that interest us. As you should always do, double check your system documentation for the full story.

The who Command

The first program is the who command. By itself, it generates a list of the users who are currently logged in. Although I’m writing this on a single-user system, we’ll pretend that several people are logged in:

     $ who
     ⊣ arnold   console Jan 22 19:57
     ⊣ miriam   ttyp0   Jan 23 14:19(:0.0)
     ⊣ bill     ttyp1   Jan 21 09:32(:0.0)
     ⊣ arnold   ttyp2   Jan 23 20:48(:0.0)

Here, the ‘$’ is the usual shell prompt, at which I typed ‘who’. There are three people logged in, and I am logged in twice. On traditional Unix systems, user names are never more than eight characters long. This little bit of trivia will be useful later. The output of who is nice, but the data is not all that exciting.

The cut Command

The next program we’ll look at is the cut command. This program cuts out columns or fields of input data. For example, we can tell it to print just the login name and full name from the ‘/etc/passwd’ file. The ‘/etc/passwd’ file has seven fields, separated by colons:

     arnold:xyzzy:2076:10:Arnold D. Robbins:/home/arnold:/bin/bash

To get the first and fifth fields, we would use cut like this:

     $ cut -d: -f1,5 /etc/passwd
     ⊣ root:Operator
     ...
     ⊣ arnold:Arnold D. Robbins
     ⊣ miriam:Miriam A. Robbins
     ...

With the ‘-c’ option, cut will cut out specific characters (i.e., columns) in the input lines. This command looks like it might be useful for data filtering.

The sort Command

Next we’ll look at the sort command. This is one of the most powerful commands on a Unix-style system; one that you will often find yourself using when setting up fancy data plumbing.

The sort command reads and sorts each file named on the command line. It then merges the sorted data and writes it to standard output. It will read standard input if no files are given on the command line (thus making it into a filter). The sort is based on the character collating sequence or based on user-supplied ordering criteria.

The uniq Command

Finally (at least for now), we’ll look at the uniq program. When sorting data, you will often end up with duplicate lines, lines that are identical. Usually, all you need is one instance of each line. This is where uniq comes in. The uniq program reads its standard input, which it expects to be sorted. It only prints out one copy of each duplicated line. It does have several options. Later on, we’ll use the ‘-c’ option, which prints each unique line, preceded by a count of the number of times that line occurred in the input.

Putting the Tools Together

Now, let’s suppose this is a large ISP server system with dozens of users logged in. The management wants the system administrator to write a program that will generate a sorted list of logged in users. Furthermore, even if a user is logged in multiple times, his or her name should only show up in the output once.

The administrator could sit down with the system documentation and write a C program that did this. It would take perhaps a couple of hundred lines of code and about two hours to write it, test it, and debug it. However, knowing the software toolbox, the administrator can instead start out by generating just a list of logged on users:

     $ who | cut -c1-8
     ⊣ arnold
     ⊣ miriam
     ⊣ bill
     ⊣ arnold

Next, sort the list:

     $ who | cut -c1-8 | sort
     ⊣ arnold
     ⊣ arnold
     ⊣ bill
     ⊣ miriam

Finally, run the sorted list through uniq, to weed out duplicates:

     $ who | cut -c1-8 | sort | uniq
     ⊣ arnold
     ⊣ bill
     ⊣ miriam

The sort command actually has a ‘-u’ option that does what uniq does. However, uniq has other uses for which one cannot substitute ‘sort -u’.

The administrator puts this pipeline into a shell script, and makes it available for all the users on the system (‘#’ is the system administrator, or root, prompt):

     # cat > /usr/local/bin/listusers
     who | cut -c1-8 | sort | uniq
     ^D
     # chmod +x /usr/local/bin/listusers

There are four major points to note here. First, with just four programs, on one command line, the administrator was able to save about two hours worth of work. Furthermore, the shell pipeline is just about as efficient as the C program would be, and it is much more efficient in terms of programmer time. People time is much more expensive than computer time, and in our modern “there’s never enough time to do everything” society, saving two hours of programmer time is no mean feat.

Second, it is also important to emphasize that with the combination of the tools, it is possible to do a special purpose job never imagined by the authors of the individual programs.

Third, it is also valuable to build up your pipeline in stages, as we did here. This allows you to view the data at each stage in the pipeline, which helps you acquire the confidence that you are indeed using these tools correctly.

Finally, by bundling the pipeline in a shell script, other users can use your command, without having to remember the fancy plumbing you set up for them. In terms of how you run them, shell scripts and compiled programs are indistinguishable.

After the previous warm-up exercise, we’ll look at two additional, more complicated pipelines. For them, we need to introduce two more tools.

The first is the tr command, which stands for “transliterate.” The tr command works on a character-by-character basis, changing characters. Normally it is used for things like mapping upper case to lower case:

$ echo ThIs ExAmPlE HaS MIXED case! | tr '[A-Z]' '[a-z]'
⊣ this example has mixed case!

There are several options of interest:

-c work on the complement of the listed characters, i.e., operations apply to characters not in the given set

-d delete characters in the first set from the output

-s squeeze repeated characters in the output into just one character.

We will be using all three options in a moment. The other command we’ll look at is comm. The comm command takes two sorted input

files as input data, and prints out the files’ lines in three columns. The output columns are the data lines unique to the first file, the data lines unique to the second file, and the data lines that are common to both. The ‘-1’, ‘-2’, and ‘-3’ command line options omit the respective columns. (This is non-intuitive and takes a little getting used to.) For example:

$ cat f1
⊣ 11111
⊣ 22222
⊣ 33333
⊣ 44444
$ cat f2
⊣ 00000
⊣ 22222
⊣ 33333
⊣ 55555
$ comm f1 f2
⊣  00000
⊣ 11111
⊣  22222
⊣  33333
⊣ 44444
⊣  55555

The single dash as a filename tells comm to read standard input instead of a regular file. Now we’re ready to build a fancy pipeline. The first application is a word frequency

counter. This helps an author determine if he or she is over-using certain words.

The first step is to change the case of all the letters in our input file to one case. “The” and “the” are the same word when doing counting.

$ tr '[A-Z]' '[a-z]' < whats.gnu | ...

The next step is to get rid of punctuation. Quoted words and unquoted words should be treated identically; it’s easiest to just get the punctuation out of the way.

$ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' | ...

The second tr command operates on the complement of the listed characters, which are all the letters, the digits, the underscore, and the blank. The ‘\012’ represents the newline character; it has to be left alone. (The ascii tab character should also be included for good measure in a production script.)

At this point, we have data consisting of words separated by blank space. The words only contain alphanumeric characters (and the underscore). The next step is break the data apart so that we have one word per line. This makes the counting operation much easier, as we will see shortly.

     $ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' |
     > tr -s '[ ]' '\012' | ...

This command turns blanks into newlines. The ‘-s’ option squeezes multiple newline characters in the output into just one. This helps us avoid blank lines. (The ‘>’ is the shell’s “secondary prompt.” This is what the shell prints when it notices you haven’t finished typing in all of a command.)

We now have data consisting of one word per line, no punctuation, all one case. We’re ready to count each word:

     $ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' |
     > tr -s '[ ]' '\012' | sort | uniq -c | ...

At this point, the data might look something like this:

       60 a
        2 able
        6 about
        1 above
        2 accomplish
        1 acquire
        1 actually
        2 additional

The output is sorted by word, not by count! What we want is the most frequently used words first. Fortunately, this is easy to accomplish, with the help of two more sort options:

-n do a numeric sort, not a textual one

-r reverse the order of the sort

The final pipeline looks like this:

     $ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' |
     > tr -s '[ ]' '\012' | sort | uniq -c | sort -nr
     ⊣ 156 the
     ⊣ 60 a
     ⊣ 58 to
     ⊣ 51 of
     ⊣ 51 and
     ...

Whew! That’s a lot to digest. Yet, the same principles apply. With six commands, on two lines (really one long one split for convenience), we’ve created a program that does something interesting and useful, in much less time than we could have written a C program to do the same thing.

A minor modification to the above pipeline can give us a simple spelling checker! To determine if you’ve spelled a word correctly, all you have to do is look it up in a dictionary. If it is not there, then chances are that your spelling is incorrect. So, we need a dictionary. The conventional location for a dictionary is ‘/usr/dict/words’. On my GNU/Linux system,^{1} this is a is a sorted, 45,402 word dictionary.

Now, how to compare our file with the dictionary? As before, we generate a sorted list of words, one per line:

     $ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' |
     > tr -s '[ ]' '\012' | sort -u | ...

Now, all we need is a list of words that are not in the dictionary. Here is where the comm command comes in.

     $ tr '[A-Z]' '[a-z]' < whats.gnu | tr -cd '[A-Za-z0-9_\012]' |
     > tr -s '[ ]' '\012' | sort -u |
     > comm -23 - /usr/dict/words

The ‘-2’ and ‘-3’ options eliminate lines that are only in the dictionary (the second file), and lines that are in both files. Lines only in the first file (standard input, our stream of words), are words that are not in the dictionary. These are likely candidates for spelling errors. This pipeline was the first cut at a production spelling checker on Unix.

There are some other tools that deserve brief mention.

grep search files for text that matches a regular expression

wc count lines, words, characters

tee a T-fitting for data pipes, copies data to files and to standard output

sed the stream editor, an advanced tool

awk a data manipulation language, another advanced tool

The software tools philosophy also espoused the following bit of advice: “Let someone else do the hard part.” This means, take something that gives you most of what you need, and then massage it the rest of the way until it’s in the form that you want.

To summarize:

1. Each program should do one thing well. No more, no less.

2. Combining programs with appropriate plumbing leads to results where the whole is greater than the sum of the parts. It also leads to novel uses of programs that the authors might never have imagined.

3. Programs should never print extraneous header or trailer data, since these could get sent on down a pipeline. (A point we didn’t mention earlier.)

4. Let someone else do the hard part.

5. Know your toolbox! Use each program appropriately. If you don’t have an appropriate tool, build one.

As of this writing, all the programs we’ve discussed are available via anonymous ftp from:

1 Redhat Linux 6.1, for the November 2000 revision of this article.

ftp://gnudist.gnu.org/textutils/textutils-1.22.tar.gz. (There may be more re-

cent versions available now.)

None of what I have presented in this column is new. The Software Tools philosophy was first introduced in the book Software Tools, by Brian Kernighan and P.J. Plauger (Addison-Wesley, ISBN 0-201-03669-X). This book showed how to write and use software tools. It was written in 1976, using a preprocessor for FORTRAN named ratfor (RATional FORtran). At the time, C was not as ubiquitous as it is now; FORTRAN was. The last chapter presented a ratfor to FORTRAN processor, written in ratfor. ratfor looks an awful lot like C; if you know C, you won’t have any problem following the code.

In 1981, the book was updated and made available as Software Tools in Pascal (Addison-Wesley, ISBN 0-201-10342-7). The first book is still in print; the second, alas, is not. Both books are well worth reading if you’re a programmer. They certainly made a major change in how I view programming.

Initially, the programs in both books were available (on 9-track tape) from Addison-Wesley. Unfortunately, this is no longer the case, although the ratfor versions are available from Brian Kernighan’s home page (http://cm.bell-labs.come/who/bwk), and you might be able to find copies of the Pascal versions floating around the Internet. For a number of years, there was an active Software Tools Users Group, whose members had ported the original ratfor programs to essentially every computer system with a FORTRAN compiler. The popularity of the group waned in the middle 1980s as Unix began to spread beyond universities.

With the current proliferation of GNU code and other clones of Unix programs, these programs now receive little attention; modern C versions are much more efficient and do more than these programs do. Nevertheless, as exposition of good programming style, and evangelism for a still-valuable philosophy, these books are unparalleled, and I recommend them highly.

Acknowledgment: I would like to express my gratitude to Brian Kernighan of Bell Labs, the original Software Toolsmith, for reviewing this column.

Appendix A GNU Free Documentation License

Version 1.1, March 2000

Copyright (C) 2000 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307 USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE

The purpose of this License is to make a manual, textbook, or other written document “free” in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others.

This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software.

We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS

This License applies to any manual or other work that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”.

A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language.

A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (For example, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, whose contents can be viewed and edited directly and straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup has been designed to thwart or discourage subsequent modification by readers is not Transparent. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTeX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML designed for human modification. Opaque formats include PostScript, PDF, proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

2. VERBATIM COPYING

You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3.

You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY

If you publish printed copies of the Document numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a publicly-accessible computer-network location containing a complete Transparent copy of the Document, free of added material, which the general network-using public has access to download anonymously at no charge using public-standard network protocols. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission. B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has less than five). C. State on the Title page the name of the publisher of the Modified Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below. G. Preserve in that license notice the full lists of Invariant Sections and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section entitled “History”, and its title, and add to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence. J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission. K. In any section entitled “Acknowledgements” or “Dedications”, preserve the section’s title, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein. L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles. M. Delete any section entitled “Endorsements”. Such a section may not be included in the Modified Version. N. Do not retitle any existing section as “Endorsements” or to conflict in title with any Invariant Section.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties–for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections entitled “History” in the various original documents, forming one section entitled “History”; likewise combine any sections entitled “Acknowledgements”, and any sections entitled “Dedications”. You must delete all sections entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, does not as a whole count as a Modified Version of the Document, provided no compilation copyright is claimed for the compilation. Such a compilation is called an “aggregate”, and this License does not apply to the other self-contained works thus compiled with the Document, on account of their being thus compiled, if they are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one quarter of the entire aggregate, the Document’s Cover Texts may be placed on covers that surround only the Document within the aggregate. Otherwise they must appear on covers around the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License provided that you also include the original English version of this License. In case of a disagreement between the translation and the original English version of this License, the original English version will prevail.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See http://www.gnu.org/copyleft/. Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

   Copyright (C)  year  your name.
   Permission is granted to copy, distribute and/or modify this document
   under the terms of the GNU Free Documentation License, Version 1.1
   or any later version published by the Free Software Foundation;
   with the Invariant Sections being list their titles, with the
   Front-Cover Texts being list, and with the Back-Cover Texts being list.
   A copy of the license is included in the section entitled ``GNU
   Free Documentation License''.

If you have no Invariant Sections, write “with no Invariant Sections” instead of saying which ones are invariant. If you have no Front-Cover Texts, write “no Front-Cover Texts” instead of “Front-Cover Texts being list”; likewise for Back-Cover Texts.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Index 160

Index

!

! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 != . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

%

% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

&

& . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 +page_range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

-

- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79, 103, 126, 128

- and Unix rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 ‘-’, removing files beginning with . . . . . . . . . . . . . 77 -- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --across. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --address-radix . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --adjustment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 --all. . . . . . . . . . . . . . . . . . . . 55, 56, 91, 93, 107, 123 --all-repeated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 --almost-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 --append . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 --archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 --author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 --backup . . . . . . . . . . . . . . . . . . . . . . . 2, 68, 74, 76, 82 --before. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --binary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 29 --block-size . . . . . . . . . . . . . . . . . . . . . . . . . 5, 91, 93 --block-size=size . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --body-numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --bourne-shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --buffer-size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --bytes . . . . . . . . . . . . . . . . 20, 21, 22, 24, 27, 46, 93 --c-shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --changes. . . . . . . . . . . . . . . . . . . . . . . . . . . . 86, 88, 89 --characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 --chars. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 --check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 --check-chars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 --classify . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --color. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 --columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 --count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35, 115 --count-links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 --crown-margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --csh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90, 120 --delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 --delimiters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --dereference . . . . . . . . . . . . . 57, 69, 87, 88, 94, 95 --dereference-args . . . . . . . . . . . . . . . . . . . . . . . . . 93

--dereference-command-line . . . . . . . . . . . . . . . . 57

--dictionary-order . . . . . . . . . . . . . . . . . . . . . . . . . 31 --digits. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 --directory . . . . . . . . . . . . . . . . . . . . . . 56, 74, 77, 82 --dired. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 --double-space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 --elide-empty-files. . . . . . . . . . . . . . . . . . . . . . . . 26 --escape. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 --exact. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 --exclude-from=FILE . . . . . . . . . . . . . . . . . . . . . . . 94 --exclude-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 --exclude=PATTERN . . . . . . . . . . . . . . . . . . . . . . . . . . 94 --expand-tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --fast. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 --field-separator . . . . . . . . . . . . . . . . . . . . . . . . . . 33 --fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 --file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107, 120 --file-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 --first-line-number. . . . . . . . . . . . . . . . . . . . . . . . 18 --follow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 --footer-numbering . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --force . . . . . . . . . . . . . . . . . . . . . . . 69, 76, 77, 79, 82 --form-feed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --format . . . . . . . . . . . . . . . . . . . . . . 11, 59, 62, 63, 95 --format=format . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 --from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 --full-time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

--general-numeric-sort . . . . . . . . . . . . . . . . . . . . 31

--group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 114 --groups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 --hardware-platform. . . . . . . . . . . . . . . . . . . . . . . 123 --header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --header-numbering . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --heading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --help . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --hide-control-chars. . . . . . . . . . . . . . . . . . . . . . . 66 --human-readable . . . . . . . . . . . . . . . . . . 5, 59, 91, 94 --idle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --ignore-backups . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 --ignore-case . . . . . . . . . . . . . . . . . . . . . . . 31, 35, 48 --ignore-environment. . . . . . . . . . . . . . . . . . . . . . 126

--ignore-fail-on-non-empty . . . . . . . . . . . . . . . . 85

--ignore-interrupts. . . . . . . . . . . . . . . . . . . . . . . 105

--ignore-leading-blanks . . . . . . . . . . . . . . . . . . . 31

--ignore-nonprinting. . . . . . . . . . . . . . . . . . . . . . . 31 --ignore=pattern . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 --indent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 --indicator-style . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --initial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 --inode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 --inodes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --interactive . . . . . . . . . . . . . . . . . . . . 69, 76, 77, 82 --iso-8601[=timespec] . . . . . . . . . . . . . . . . . . . . 120 --iterations=NUMBER . . . . . . . . . . . . . . . . . . . . . . . 79 --join-blank-lines . . . . . . . . . . . . . . . . . . . . . . . . . 10 --join-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --keep-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 --kernel-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --kernel-release . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --kernel-version . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --key . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 --length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --line-bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 --lines . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 23, 24, 27 --link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 --literal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 --local. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --login. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 --lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --machine . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --max-depth=DEPTH . . . . . . . . . . . . . . . . . . . . . . . . . . 94 --max-line-length . . . . . . . . . . . . . . . . . . . . . . . . . . 27 --max-unchanged-stats . . . . . . . . . . . . . . . . . . . . . 23 --merge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 30 --mesg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --message . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74, 83, 84 --month-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 --name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 --no-create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 --no-dereference. . . . . . . . . . . . . . . . . 69, 82, 87, 88 --no-file-warnings . . . . . . . . . . . . . . . . . . . . . . . . . 18 --no-group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 --no-renumber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --no-sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --nodename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --number-format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --number-lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 --number-nonblank . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --number-separator . . . . . . . . . . . . . . . . . . . . . . . . . 10 --number-width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 --numeric-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --numeric-uid-gid . . . . . . . . . . . . . . . . . . . . . . . . . . 60 --omit-header . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --omit-pagination . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --one-file-system . . . . . . . . . . . . . . . . . . . . . . 72, 94 --only-delimited . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 --operating-system. . . . . . . . . . . . . . . . . . . . . . . . 123 --output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 --output-delimiter . . . . . . . . . . . . . . . . . . . . . . . . . 46 --output-duplicates. . . . . . . . . . . . . . . . . . . . . . . . 12 --output-tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 --owner. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 --page-increment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --page_width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --pages=page_range . . . . . . . . . . . . . . . . . . . . . . . . . 16 --parents. . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 83, 85 --pid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 --portability. . . . . . . . . . . . . . . . . . . . . . . . . . 92, 106 --prefix. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 --preserve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

                                                --preserve-environment . . . . . . . . . . . . . . . . . . . 128

--preserve-timestamps . . . . . . . . . . . . . . . . . . . . . 74 --print-database . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --print-type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --processor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 --quiet . . . . . . . . . . . . . . . 21, 23, 26, 86, 88, 89, 113 --quote-name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 --quoting-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 --read-bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --real. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 --recursive . . . . . . . . . . . . . . . 57, 70, 77, 87, 88, 89 --reference. . . . . . . . . . . . . . . . . . 87, 88, 89, 90, 121 --regex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --remove. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 --remove-destination. . . . . . . . . . . . . . . . . . . . . . . 71 --repeated . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --reply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 76 --retry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 --reverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 61 --rfc-822 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 --save. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 --section-delimiter . . . . . . . . . . . . . . . . . . . . . . . . . 9 --sep-string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 --separate-dirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 --separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 18 --serial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 --set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 --sh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 --shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 --show-all . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --show-control-chars . . . . . . . . . . . . . . . . . . . 16, 66 --show-ends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --show-nonprinting . . . . . . . . . . . . . . . . . . . . . . 8, 19 --show-tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --si . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 60, 91, 94 --silent . . . . . . . . . . . . . . 21, 23, 26, 86, 88, 89, 113 --size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 --size=BYTES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 --skip-bytes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --skip-chars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --skip-fields . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 --sleep-interval . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 --sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 --spaces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 --sparse=when . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --split-only . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --squeeze-blank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

                                                --starting-line-number . . . . . . . . . . . . . . . . . . . . 10

--status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 --strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 --strip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

--strip-trailing-slashes . . . . . . . . . . . . . . . 71, 76

--suffix. . . . . . . . . . . . . . . . . . . . 3, 25, 71, 75, 76, 82 --suffix-length . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 --summarize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 --symbolic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 --symbolic-link . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 --sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --sysv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 --tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54, 55 --tabsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 --tagged-paragraph . . . . . . . . . . . . . . . . . . . . . . . . . 14 --target-directory . . . . . . . . . . . . 5, 71, 75, 76, 82 --temporary-directory . . . . . . . . . . . . . . . . . . . . . 33 --terse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 --text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 --time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61, 90 --time-style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 --total. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 --traditional . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 --uniform-spacing . . . . . . . . . . . . . . . . . . . . . . . . . . 14 --unique . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33, 36 --universal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 --update. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 --user. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 --utc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 --verbose … 21, 23, 24, 71, 75, 76, 77, 79, 83, 85,

87, 88, 89 --version. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --version-control . . . . . . . . . . . . . 3, 72, 75, 76, 83 --warn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 --width . . . . . . . . . . . . . . . . . . . . . . . 12, 15, 19, 20, 64 --words. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 --writable . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 --zero . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 --zero-terminated . . . . . . . . . . . . . . . . . . . . . . . . . . 33 -1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36, 48, 62 -2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36, 48 -3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 -a . . 12, 16, 24, 48, 55, 56, 68, 90, 91, 93, 102, 105,

107, 123 -A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 56 -b. . 2, 7, 8, 9, 12, 20, 24, 25, 29, 31, 46, 66, 67, 68,

74, 76, 82, 93, 100 -B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 56, 91, 93 -c …. 12, 14, 16, 21, 22, 27, 30, 35, 46, 61, 67, 74,

86, 88, 89, 90, 93, 95, 100, 127 -C. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24, 62 -column . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 -d. . 9, 12, 16, 31, 35, 46, 47, 56, 69, 74, 77, 82, 90,

100, 120 -D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36, 57, 93 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 17, 48, 97, 100 -E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -ef . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -eq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -f. . 9, 12, 17, 22, 25, 31, 35, 46, 61, 69, 76, 77, 79,

82, 86, 88, 89, 90, 95, 100, 120, 128 -F . . . . . . . . . . . . . . . . . . . . . . . . . . . 17, 23, 63, 82, 107 -f format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 -g . . . . . . . . . . . . . . . . . . . . . . 31, 59, 74, 100, 107, 114 -G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59, 100, 114 -ge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -gt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -h . . . . . . . . . . . . 5, 9, 12, 17, 59, 87, 88, 91, 94, 100 -H . . . . . . . . . . . . . . . . . . . . . . . . . . . 57, 69, 91, 94, 116 -i. . 9, 12, 17, 31, 35, 48, 54, 59, 69, 76, 77, 82, 92,

105, 116, 123, 126 -I . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 -I timespec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 -j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 -J . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 -j1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 -j2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 -k. . . . . . . . . . . . . . . . . . . . . . 5, 26, 33, 63, 92, 94, 100 -l . . . . . . 10, 13, 17, 24, 27, 59, 69, 92, 94, 116, 128 -L . . . . . . . . . . . . . . . . . . . . . . . . 27, 57, 69, 94, 95, 100 -le . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -lt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -m . . . . . 17, 27, 30, 63, 74, 83, 84, 90, 115, 123, 128 -M . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 -n …. 7, 10, 18, 21, 23, 26, 32, 46, 60, 82, 97, 101,

114, 123, 126 -N . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 18, 66 -n NUMBER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 -ne . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 -nt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 -o . . . . . . . . . . . . . . . . . . . 13, 18, 32, 60, 74, 102, 123 -O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 -ot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 -p . . . . . . . 10, 67, 70, 74, 83, 85, 100, 106, 123, 128 -P. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69, 92 -q . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 23, 26, 66, 115 -Q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 -r. . 8, 18, 28, 32, 61, 70, 77, 90, 100, 114, 121, 123 -R. . . . . . . . . . . . . . . . . . . . . 57, 70, 77, 87, 88, 89, 121 -s… 7, 8, 10, 11, 14, 18, 20, 26, 28, 35, 46, 47, 60, 71, 74, 82, 94, 100, 113, 115, 121, 123, 128 -S . . . . . . . . . . . 3, 19, 32, 61, 71, 75, 76, 82, 94, 100 -s BYTES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 ‘-su’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 -t. . . . . 8, 11, 14, 19, 29, 33, 54, 55, 61, 92, 95, 100 -T . . . . . . . . . . . . . . . . . . . . . . . . . 8, 19, 33, 63, 92, 116 -u …. 8, 14, 33, 36, 61, 76, 79, 100, 114, 116, 121,

126 -U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -unset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 -v. . 8, 10, 12, 19, 21, 23, 61, 71, 75, 76, 77, 79, 83,

85, 87, 88, 89, 123 -V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72, 75, 76, 83 -w . . . . . . 10, 12, 15, 19, 20, 27, 29, 36, 64, 100, 116 -W . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 -width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 -x. . . . . . . . . . . . . . . . . . . . . 13, 63, 72, 79, 93, 94, 100 -X . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 -X FILE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 -z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 33, 79, 101

.

.cshrc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128

/

/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 /bin/sh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 /etc/passwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 /etc/shells . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 /etc/utmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114, 115 /etc/wtmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

<

< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 <= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

=

=. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101, 103 == . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

>

> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 >= . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

_POSIX2_VERSION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

\

\( regexp operator. . . . . . . . . . . . . . . . . . . . . . . . . . 102 \+ regexp operator. . . . . . . . . . . . . . . . . . . . . . . . . . 102 \? regexp operator. . . . . . . . . . . . . . . . . . . . . . . . . . 102 \| regexp operator. . . . . . . . . . . . . . . . . . . . . . . . . . 102 \0ooo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 \c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 \uhhhh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 \Uhhhhhhhh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 \xhh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

|

| . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103

1

128-bit checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 16-bit checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

4

4.2 filesystem type . . . . . . . . . . . . . . . . . . . . . . . . . . 93

A

abbreviations for months . . . . . . . . . . . . . . . . . . . . 143 access permission tests . . . . . . . . . . . . . . . . . . . . . . 100 access permissions, changing . . . . . . . . . . . . . . . . . . 88 access time, changing . . . . . . . . . . . . . . . . . . . . . . . 90 access time, printing or sorting files by . . . . . . . 61 across columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 across, listing files. . . . . . . . . . . . . . . . . . . . . . . . . . . 63 adding permissions. . . . . . . . . . . . . . . . . . . . . . . . . . 136 addition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 ago in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . 145 all duplicate lines, outputting . . . . . . . . . . . . . . . . . 36 alnum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 alternate ebcdic, converting to . . . . . . . . . . . . . . 73 always color option . . . . . . . . . . . . . . . . . . . . . . . . . . 62 am i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 am in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 and operator . . . . . . . . . . . . . . . . . . . . . . . . . . . 102, 103 appropriate privileges . . . . . . . . . . . 74, 119, 123, 126 arbitrary date strings, parsing . . . . . . . . . . . . . . . 120 arbitrary text, displaying . . . . . . . . . . . . . . . . . . . . . 97 arithmetic tests. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 ASCII dump of files. . . . . . . . . . . . . . . . . . . . . . . . . . 10 ascii, converting to . . . . . . . . . . . . . . . . . . . . . . . . . 73 atime, changing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 atime, printing or sorting files by . . . . . . . . . . . . . 61 attributes, file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 authors of getdate. . . . . . . . . . . . . . . . . . . . . . . . . . 145 auto color option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 automounter filesystems . . . . . . . . . . . . . . . . . . . . . . 91

B

b for block special file . . . . . . . . . . . . . . . . . . . . . . . . 84 background jobs, stopping at terminal write . . 110 backslash escapes. . . . . . . . . . . . . . . . . . . . . . . . . 50, 97 backslash sequences for file names . . . . . . . . . . . . . 66 backup files, ignoring. . . . . . . . . . . . . . . . . . . . . . . . . 56 backup options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 backup suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 backups, making . . . . . . . . . . . . . . . . 2, 68, 74, 76, 82 backups, making only . . . . . . . . . . . . . . . . . . . . . . . . 68 basename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 baud rate, setting. . . . . . . . . . . . . . . . . . . . . . . . . . . 112 beeping at input buffer full . . . . . . . . . . . . . . . . . . 109 beginning of time . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 beginning of time, for posix . . . . . . . . . . . . . . . . . 141 Bellovin, Steven M. . . . . . . . . . . . . . . . . . . . . . . . . . 145 Berets, Jim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 Berry, K. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 145 binary and text I/O in cat. . . . . . . . . . . . . . . . . . . . . 7 binary input files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 blank . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 blank lines, numbering . . . . . . . . . . . . . . . . . . . . . . . 10 blanks, ignoring leading . . . . . . . . . . . . . . . . . . . . . . 31 block (space-padding) . . . . . . . . . . . . . . . . . . . . . . . 73 block size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 72 block size of conversion. . . . . . . . . . . . . . . . . . . . . . . 72 block size of input . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 block size of output . . . . . . . . . . . . . . . . . . . . . . . . . . 72 block special check. . . . . . . . . . . . . . . . . . . . . . . . . . 100 block special files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 block special files, creating . . . . . . . . . . . . . . . . . . . 84 BLOCK_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 body, numbering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 Bourne shell syntax for color setup . . . . . . . . . . . . 67 breaks, cause interrupts . . . . . . . . . . . . . . . . . . . . . 108 breaks, ignoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 brkint. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 bs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 BSD sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 BSD tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 BSD touch compatibility . . . . . . . . . . . . . . . . . . . . . 90 bsn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 built-in shell commands, conflicts with … 99, 107,

126 byte count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 byte-swapping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

C

c for character special file . . . . . . . . . . . . . . . . . . . . 84 C shell syntax for color setup . . . . . . . . . . . . . . . . . 67 C-s/C-q flow control . . . . . . . . . . . . . . . . . . . . . . 108 calendar date item . . . . . . . . . . . . . . . . . . . . . . . . . . 142 case folding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 case translation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 case, ignored in dates . . . . . . . . . . . . . . . . . . . . . . . 142 cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 cbreak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 cbs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 CD-ROM filesystem type . . . . . . . . . . . . . . . . . . . . . 93 cdfs filesystem type . . . . . . . . . . . . . . . . . . . . . . . . . 93 change or print terminal settings . . . . . . . . . . . . . 107 changed files, verbosely describing. . . . . . . . . . . . . 88 changed owners, verbosely describing . . . . . . . . . . 86 changing access permissions . . . . . . . . . . . . . . . . . . 88 changing file attributes . . . . . . . . . . . . . . . . . . . . . . . 86 changing file ownership. . . . . . . . . . . . . . . . . . . . . . . 86 changing file timestamps . . . . . . . . . . . . . . . . . . . . . 89 changing group ownership. . . . . . . . . . . . . . . . . 86, 87 changing special permissions . . . . . . . . . . . . . . . . . 137 character classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 character count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 character size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 character special check . . . . . . . . . . . . . . . . . . . . . . 100 character special files. . . . . . . . . . . . . . . . . . . . . . . . . 84 character special files, creating . . . . . . . . . . . . . . . . 84 characters, special . . . . . . . . . . . . . . . . . . . . . . . . . . 111 check file types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 checking for sortedness . . . . . . . . . . . . . . . . . . . . . . . 30 checksum, 128-bit. . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 checksum, 16-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 chgrp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 chroot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 cksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 clocal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 cntrl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 color database, printing . . . . . . . . . . . . . . . . . . . . . . 67 color setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 color, distinguishing file types with . . . . . . . . . . . . 62 cols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 columns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 COLUMNS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64, 112 combination settings . . . . . . . . . . . . . . . . . . . . . . . . 110 comm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 commands for controlling processes . . . . . . . . . . 129 commands for delaying . . . . . . . . . . . . . . . . . . . . . . 131 commands for exit status . . . . . . . . . . . . . . . . . . . . . 99 commands for file name manipulation . . . . . . . . 106 commands for invoking other commands. . . . . . 125 commands for printing text . . . . . . . . . . . . . . . . . . . 97 commands for printing the working context … 107 commands for printing user information . . . . . . 114 commands for redirection. . . . . . . . . . . . . . . . . . . . 105 commands for system context. . . . . . . . . . . . . . . . 117 commas, outputting between files . . . . . . . . . . . . . . 63 comments, in dates . . . . . . . . . . . . . . . . . . . . . . . . . 142 common field, joining on . . . . . . . . . . . . . . . . . . . . . 47 common lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 compare values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 comparing sorted files . . . . . . . . . . . . . . . . . . . . . . . . 36 comparison operators . . . . . . . . . . . . . . . . . . . . . . . 103 concatenate and write files. . . . . . . . . . . . . . . . . . . . . 7 conditional executability . . . . . . . . . . . . . . . . . . . . 138 conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 conflicts with shell built-ins . . . . . . . . . . 99, 107, 126 connectives, logical . . . . . . . . . . . . . . . . . . . . . 101, 103 context splitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 context, system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 control characters, using ‘^c’ . . . . . . . . . . . . . . . . 110 control settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 conv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 conversion block size . . . . . . . . . . . . . . . . . . . . . . . . . 72 converting tabs to spaces . . . . . . . . . . . . . . . . . . . . . 54 converting while copying a file . . . . . . . . . . . . . . . . 72 cooked. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 Coordinated Universal Time . . . . . . . . . . . . . . . . . 121 copying directories recursively . . . . . . . . . . . . . 69, 70 copying existing permissions . . . . . . . . . . . . . . . . . 137 copying files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 copying files and directories. . . . . . . . . . . . . . . . . . . 68 copying files and setting attributes . . . . . . . . . . . . 73 count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 cp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 crashes and corruption . . . . . . . . . . . . . . . . . . . . . . . 96 CRC checksum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 cread . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 creating directories. . . . . . . . . . . . . . . . . . . . . . . . . . . 83 creating FIFOs (named pipes) . . . . . . . . . . . . . . . . 84 creating links (hard only). . . . . . . . . . . . . . . . . . . . . 81 creating links (hard or soft). . . . . . . . . . . . . . . . . . . 81 crn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 crown margin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 crt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 crterase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 crtkill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 crtscts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 csh syntax for color setup . . . . . . . . . . . . . . . . . . . . 67 csn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 csplit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 cstopb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 ctime, printing or sorting by. . . . . . . . . . . . . . . . . . 61 ctlecho. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 current working directory, printing . . . . . . . . . . . 107 cut . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 cyclic redundancy check . . . . . . . . . . . . . . . . . . . . . . 28

D

data, erasing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 database for color setup, printing . . . . . . . . . . . . . 67 date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 date directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 date format, iso 8601 . . . . . . . . . . . . . . . . . . . . . . . 143 date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . 141 date options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 date strings, parsing . . . . . . . . . . . . . . . . . . . . . . . . 120 day in date strings . . . . . . . . . . . . . . . . . . . . . 144, 145 day of week item . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 dd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 dec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 decctlq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 delay for a specified time . . . . . . . . . . . . . . . . . . . . 131 delaying commands . . . . . . . . . . . . . . . . . . . . . . . . . 131 deleting characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 dereferencing symbolic links . . . . . . . . . . . . . . . . . . 81 descriptor follow option . . . . . . . . . . . . . . . . . . . . . 22 destination directory . . . . . . . . . . . . 5, 71, 75, 76, 82 destinations, multiple output . . . . . . . . . . . . . . . . 105 device file, disk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 df . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 DF_BLOCK_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 dictionary order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 differing lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 digit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 dircolors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 directives, date . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 directives, literal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 directives, time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 directories, copying . . . . . . . . . . . . . . . . . . . . . . . . . . 68 directories, copying recursively . . . . . . . . . . . . 69, 70 directories, creating . . . . . . . . . . . . . . . . . . . . . . . . . . 83 directories, creating with given attributes . . . . . . 74 directories, removing (recursively) . . . . . . . . . . . . . 77 directories, removing empty . . . . . . . . . . . . . . . . . . 85 directories, removing with unlink . . . . . . . . . . . . . 77 directory check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 directory components, printing. . . . . . . . . . . . . . . 106 directory deletion, ignoring failures . . . . . . . . . . . . 85 directory deletion, reporting . . . . . . . . . . . . . . . . . . 85 directory listing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 directory listing, brief . . . . . . . . . . . . . . . . . . . . . . . . 67 directory listing, recursive . . . . . . . . . . . . . . . . . . . . 57 directory listing, verbose . . . . . . . . . . . . . . . . . . . . . 67 directory order, listing by . . . . . . . . . . . . . . . . . . . . 61 directory, stripping from file names. . . . . . . . . . . 106 dired Emacs mode support . . . . . . . . . . . . . . . . . . . 57 dirname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 disabling special characters . . . . . . . . . . . . . . . . . . 111 disk allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 disk device file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 disk usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 disk usage by filesystem . . . . . . . . . . . . . . . . . . . . . . 91 disk usage for files . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 diskette filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 displacement of dates . . . . . . . . . . . . . . . . . . . . . . . 144 displaying text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 division . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 do nothing, successfully . . . . . . . . . . . . . . . . . . . . . . 99 do nothing, unsuccessfully . . . . . . . . . . . . . . . . . . . . 99 DOS filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 double spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 down columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 dsusp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 du . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 DU_BLOCK_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 duplicate lines, outputting . . . . . . . . . . . . . . . . . . . . 35

E

ebcdic, converting to . . . . . . . . . . . . . . . . . . . . . . . . 73 echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97, 110 echoctl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 echoe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 echok . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 echoke. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 echonl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 echoprt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 effective uid and gid, printing . . . . . . . . . . . . . . . . 114 effective UID, printing . . . . . . . . . . . . . . . . . . . . . . 114 efs filesystem type . . . . . . . . . . . . . . . . . . . . . . . . . . 93 Eggert, Paul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 eight-bit characters . . . . . . . . . . . . . . . . . . . . . 108, 111 eight-bit input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 ek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 empty files, creating . . . . . . . . . . . . . . . . . . . . . . . . . 89 empty lines, numbering . . . . . . . . . . . . . . . . . . . . . . 10 entire files, output of . . . . . . . . . . . . . . . . . . . . . . . . . . 7 env . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 environment variables, printing . . . . . . . . . . . . . . 113 environment, preserving . . . . . . . . . . . . . . . . . . . . . 128 environment, printing . . . . . . . . . . . . . . . . . . . . . . . 126 environment, running a program in a modified

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 eof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 eol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 eol2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 epoch, for posix . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 epoch, seconds since . . . . . . . . . . . . . . . . . . . . . . . . 117 equal string check. . . . . . . . . . . . . . . . . . . . . . . . . . . 101 equivalence classes . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 erase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 erasing data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 error messages, omitting . . . . . . . . . . . . . . . 86, 88, 89 evaluation of expressions . . . . . . . . . . . . . . . . . . . . 102 even parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 evenp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 exabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . . 4 examples of date . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 examples of expr . . . . . . . . . . . . . . . . . . . . . . . . . . . 104 exbibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 excluding files from du . . . . . . . . . . . . . . . . . . . . . . . 94 executable file check . . . . . . . . . . . . . . . . . . . . . . . . 100 executables and file type, marking. . . . . . . . . . . . . 63 execute permission . . . . . . . . . . . . . . . . . . . . . . . . . . 135 execute permission, symbolic . . . . . . . . . . . . . . . . 136 existence-of-file check . . . . . . . . . . . . . . . . . . . . . . . 100 existing backup method. . . . . . . . . . . . . . . . . . . . . . 3 exit status commands . . . . . . . . . . . . . . . . . . . . . . . . 99 exit status of expr . . . . . . . . . . . . . . . . . . . . . . . . . . 102 exit status of false . . . . . . . . . . . . . . . . . . . . . . . . . . 99 exit status of nohup . . . . . . . . . . . . . . . . . . . . . . . . . 127 exit status of pathchk . . . . . . . . . . . . . . . . . . . . . . . 106 exit status of printenv. . . . . . . . . . . . . . . . . . . . . . 113 exit status of true . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 exit status of tty . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 expand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 expr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 expression evaluation . . . . . . . . . . . . . . . . . . . . 99, 102 expressions, numeric . . . . . . . . . . . . . . . . . . . . . . . . 103 expressions, string . . . . . . . . . . . . . . . . . . . . . . . . . . 102 extension, sorting files by. . . . . . . . . . . . . . . . . . . . 61

F

factor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 failure exit status . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 false . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 fascism . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 ffn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 field separator character . . . . . . . . . . . . . . . . . . . . . . 33 fields, padding numeric. . . . . . . . . . . . . . . . . . . . . . 119 FIFOs, creating. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 file attributes, changing . . . . . . . . . . . . . . . . . . . . . . 86 file characteristic tests . . . . . . . . . . . . . . . . . . . . . . 100 file contents, dumping unambiguously . . . . . . . . . 10 file information, preserving . . . . . . . . . . . . . . . . . . . 70 file name manipulation . . . . . . . . . . . . . . . . . . . . . . 106 file name pattern expansion, disabled. . . . . . . . . 128 file names, checking validity and portability… 106 file names, stripping directory and suffix . . . . . . 106 file offset radix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 file ownership, changing . . . . . . . . . . . . . . . . . . . . . . 86 file permissions, numeric . . . . . . . . . . . . . . . . . . . . 139 file sizes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 file space usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 file status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 file timestamps, changing . . . . . . . . . . . . . . . . . . . . . 89 file type and executables, marking. . . . . . . . . . . . . 63 file type tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 file type, marking . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 file types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 file types, special . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 file utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 files beginning with ‘-’, removing . . . . . . . . . . . . . 77 files, copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 filesystem disk usage . . . . . . . . . . . . . . . . . . . . . . . . . 91 filesystem sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 filesystem space, retrieving current data more

slowly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 filesystem space, retrieving old data more quickly

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 filesystem status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 filesystem types, limiting output to certain. . . . . 92 filesystem types, printing . . . . . . . . . . . . . . . . . . . . . 92 filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 filesystems and hard links . . . . . . . . . . . . . . . . . . . . 81 filesystems, omitting copying to different . . . . . . 72 fingerprint, 128-bit . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 first in date strings . . . . . . . . . . . . . . . . . . . . . . . . 142 first part of files, outputting . . . . . . . . . . . . . . . . . . 21 flow control, hardware . . . . . . . . . . . . . . . . . . . . . . 108 flow control, software . . . . . . . . . . . . . . . . . . . . . . . 108 flushing, disabling . . . . . . . . . . . . . . . . . . . . . . . . . . 110 fmt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 fold . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 folding long input lines . . . . . . . . . . . . . . . . . . . . . . . 20 footers, numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 force deletion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 formatting file contents. . . . . . . . . . . . . . . . . . . . . . . 14 formatting of numbers in seq . . . . . . . . . . . . . . . . 132 formatting times . . . . . . . . . . . . . . . . . . . . . . . . 16, 117 fortnight in date strings . . . . . . . . . . . . . . . . . . . 144 fsck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

G

general date syntax . . . . . . . . . . . . . . . . . . . . . . . . . 141 general numeric sort . . . . . . . . . . . . . . . . . . . . . . . . . 31 getdate. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 gibibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 gigabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 giving away permissions . . . . . . . . . . . . . . . . . . . . . 139 globbing, disabled . . . . . . . . . . . . . . . . . . . . . . . . . . 128 GMT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 grand total of disk space . . . . . . . . . . . . . . . . . . . . . 93 graph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 Greenwich Mean Time . . . . . . . . . . . . . . . . . . . . . . 121 group owner, default . . . . . . . . . . . . . . . . . . . . . . . . 135 group ownership of installed files, setting . . . . . . 74 group ownership, changing . . . . . . . . . . . . . . . . 86, 87 group wheel, not supported. . . . . . . . . . . . . . . . . . 128 group, permissions for. . . . . . . . . . . . . . . . . . . . . . . 136 groups. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 growing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

H

hangups, immunity to . . . . . . . . . . . . . . . . . . . . . . . 126 hard link check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 hard link, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 hard links to directories . . . . . . . . . . . . . . . . . . . . . . 82 hard links, counting in du . . . . . . . . . . . . . . . . . . . . 94 hard links, creating . . . . . . . . . . . . . . . . . . . . . . . . . . 81 hard links, preserving . . . . . . . . . . . . . . . . . . . . . . . . 69 hardware class. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 hardware flow control . . . . . . . . . . . . . . . . . . . . . . . 108 hardware platform . . . . . . . . . . . . . . . . . . . . . . . . . . 123 hardware type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 hat notation for control characters . . . . . . . . . . . 110 head . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 headers, numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 help, online. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 hex dump of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 High Sierra filesystem . . . . . . . . . . . . . . . . . . . . . . . . 93 holes, copying files with . . . . . . . . . . . . . . . . . . . . . . 71 HOME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 horizontal, listing files . . . . . . . . . . . . . . . . . . . . . . 63 host processor type . . . . . . . . . . . . . . . . . . . . . . . . . 123 hostid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 hour in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 144 hsfs filesystem type . . . . . . . . . . . . . . . . . . . . . . . . . 93 human-readable output . . . . . . . . . . . . . 3, 59, 91, 94 hup[cl]. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 hurd, author, printing . . . . . . . . . . . . . . . . . . . . . . . . 57

I

ibs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 icanon. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 icrnl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 idle time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 iexten. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ignbrk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 igncr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 ignore filesystems . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 ignoring case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 ignpar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 imaxbel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 immunity to hangups . . . . . . . . . . . . . . . . . . . . . . . 126 implementation, hardware . . . . . . . . . . . . . . . . . . . 123 indenting lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 information, about current users . . . . . . . . . . . . . 115 initial part of files, outputting . . . . . . . . . . . . . . . . 21 initial tabs, converting . . . . . . . . . . . . . . . . . . . . . . . 54 inlcr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 inode number, printing . . . . . . . . . . . . . . . . . . . . . . . 59 inode usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 92 inode, and hard links. . . . . . . . . . . . . . . . . . . . . . . . . 81 inodes, written buffered . . . . . . . . . . . . . . . . . . . . . . 96 inpck . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 input block size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 input settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 input tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 install. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 interactivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70, 76 intr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 invocation of commands, modified. . . . . . . . . . . . 125 isig . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 iso 8601 date format . . . . . . . . . . . . . . . . . . . . . . . . 143 ispeed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 istrip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 items in date strings . . . . . . . . . . . . . . . . . . . . . . . . 141 iterations, selecting the number of. . . . . . . . . . . . . 79 iuclc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 ixany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 ixoff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 ixon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108

J

join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

K

kernel name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 kernel release . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 kernel version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 kibibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 kibibytes for file sizes . . . . . . . . . . . . . . . . . . . . . . . . 94 kibibytes for filesystem sizes . . . . . . . . . . . . . . . . . . 92 kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112, 129 kilobyte, definition of. . . . . . . . . . . . . . . . . . . . . . . . . . 4 Knuth, Donald E.. . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

L

language, in dates . . . . . . . . . . . . . . . . . . . . . . . . . . 142 last day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 144 last in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 142 last part of files, outputting . . . . . . . . . . . . . . . . . . 21 LC_ALL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 56 LC_COLLATE . . . . . . . . . . . . . . . . . . . 30, 35, 36, 47, 103 LC_CTYPE. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 LC_MESSAGES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 LC_NUMERIC. . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 31, 32 LC_TIME . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 31, 64, 65 lcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 LCASE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 lcase, converting to . . . . . . . . . . . . . . . . . . . . . . . . . 73 lchown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87, 88 leading directories, creating missing . . . . . . . . . . . 74 leading directory components, stripping . . . . . . 106 left margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 length. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 limiting output of du . . . . . . . . . . . . . . . . . . . . . . . . . 94 line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 line count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 line numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 line settings of terminal . . . . . . . . . . . . . . . . . . . . . 107 line-breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 line-by-line comparison . . . . . . . . . . . . . . . . . . . . . . . 36 LINES . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 links, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Linux filesystem types. . . . . . . . . . . . . . . . . . . . . . . . 93 literal directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 litout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 ln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 ln format for nl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 lnext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 local filesystem types. . . . . . . . . . . . . . . . . . . . . . . . . 93 local settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 logging out and continuing to run . . . . . . . . . . . . 126 logical and operator . . . . . . . . . . . . . . . . . . . . 102, 103 logical connectives. . . . . . . . . . . . . . . . . . . . . . 101, 103 logical or operator. . . . . . . . . . . . . . . . . . . . . . 102, 103 logical pages, numbering on. . . . . . . . . . . . . . . . . . . . 8 login name, printing . . . . . . . . . . . . . . . . . . . . . . . . 114 login sessions, printing users with . . . . . . . . . . . . 115 login shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 login shell, creating . . . . . . . . . . . . . . . . . . . . . . . . . 128 login time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 logname. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 LOGNAME. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 long ls format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 lower . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 lowercase, translating to output . . . . . . . . . . . . . . 109 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 LS_BLOCK_SIZE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 LS_COLORS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

M

machine type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 machine-readable stty output . . . . . . . . . . . . . . . 107 MacKenzie, D.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 MacKenzie, David . . . . . . . . . . . . . . . . . . . . . . . . . . 145 Makefiles, installing programs in . . . . . . . . . . . . . . 73 manipulating files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 manipulation of file names. . . . . . . . . . . . . . . . . . . 106 match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 matching patterns . . . . . . . . . . . . . . . . . . . . . . . . . . 102 md5sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 mebibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . 4 megabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . 4 merging files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 merging files in parallel. . . . . . . . . . . . . . . . . . . . . . . 15 merging sorted files . . . . . . . . . . . . . . . . . . . . . . . . . . 30 message status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 message-digest, 128-bit . . . . . . . . . . . . . . . . . . . . . . . 28 Meyering, J. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Meyering, Jim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 midnight in date strings . . . . . . . . . . . . . . . . . . . . 143 min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 minute in date strings. . . . . . . . . . . . . . . . . . . . . . . 144 minutes, time zone correction by . . . . . . . . . . . . . 143 MIT AI lab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 mkfifo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 mknod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 modem control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 modes and umask. . . . . . . . . . . . . . . . . . . . . . . . . . . 139 modes of created directories, setting . . . . . . . . . . . 83 modes of created FIFOs, setting . . . . . . . . . . . . . . 84 modification time, sorting files by . . . . . . . . . . . 61 modified command invocation . . . . . . . . . . . . . . . 125 modified environment, running a program in a

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 modify time, changing . . . . . . . . . . . . . . . . . . . . . . . 90 modifying scheduling priority . . . . . . . . . . . . . . . . 126 month in date strings . . . . . . . . . . . . . . . . . . . . . . . . 144 month names in date strings . . . . . . . . . . . . . . . . . 143 months, sorting by . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 months, written-out. . . . . . . . . . . . . . . . . . . . . . . . . 142 MS-DOS filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . 93 mtime, changing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 multicolumn output, generating . . . . . . . . . . . . . . . 15 multiple changes to permissions . . . . . . . . . . . . . . 138 multiplication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 multipliers after numbers . . . . . . . . . . . . . . . . . . . . . 72 mv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

N

name follow option . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 name of kernel. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 named pipe check. . . . . . . . . . . . . . . . . . . . . . . . . . . 100 named pipes, creating . . . . . . . . . . . . . . . . . . . . . . . . 84 network node name . . . . . . . . . . . . . . . . . . . . . . . . . 123 newer files, moving only . . . . . . . . . . . . . . . . . . . . . . 76 newer-than file check . . . . . . . . . . . . . . . . . . . . . . . . 100 newline echoing after kill. . . . . . . . . . . . . . . . . . . 110 newline, echoing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 newline, translating to crlf. . . . . . . . . . . . . . . . . . . 109 newline, translating to return . . . . . . . . . . . . . . . . 108 next day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120, 144 next in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 142 NFS filesystem type. . . . . . . . . . . . . . . . . . . . . . . . . . 92 NFS mounts from BSD to HP-UX . . . . . . . . . 60, 94 nice . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 nl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 110 nln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 no-op . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 node name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 noerror. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 noflsh. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 nohup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 nohup.out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 non-directories, copying as special files . . . . . 69, 70 non-directory suffix, stripping. . . . . . . . . . . . . . . . 106 none backup method . . . . . . . . . . . . . . . . . . . . . . . . . . 2 none color option . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 none, sorting option for ls. . . . . . . . . . . . . . . . . . . . 61 nonempty file check . . . . . . . . . . . . . . . . . . . . . . . . . 100 nonprinting characters, ignoring. . . . . . . . . . . . . . . 31 nonzero-length string check . . . . . . . . . . . . . . . . . . 101 noon in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 143 not-equal string check . . . . . . . . . . . . . . . . . . . . . . . 101 notrunc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 now in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . 145 numbered backup method. . . . . . . . . . . . . . . . . . . . . . 3 numbering lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 numbers, written-out. . . . . . . . . . . . . . . . . . . . . . . . 142 numeric expressions . . . . . . . . . . . . . . . . . . . . . . . . . 103 numeric field padding . . . . . . . . . . . . . . . . . . . . . . . 119 numeric modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 numeric operations . . . . . . . . . . . . . . . . . . . . . . . . . 132 numeric sequences . . . . . . . . . . . . . . . . . . . . . . . . . . 132 numeric sort. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 numeric tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 numeric uid and gid. . . . . . . . . . . . . . . . . . . . . . . . . . 60

O

obs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ocrnl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 octal dump of files . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 octal numbers for file modes . . . . . . . . . . . . . . . . . 139 od . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 odd parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 oddp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ofdel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 ofill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 olcuc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 older-than file check . . . . . . . . . . . . . . . . . . . . . . . . 100 one filesystem, restricting du to . . . . . . . . . . . . . . . 94 one-line output format . . . . . . . . . . . . . . . . . . . . . . . 92 onlcr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 onlret. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 onocr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 operating on characters . . . . . . . . . . . . . . . . . . . . . . 50 operating on sorted files . . . . . . . . . . . . . . . . . . . . . . 30 operating system name . . . . . . . . . . . . . . . . . . . . . . 123 opost . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 option delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 options for date . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 or operator . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102, 103 ordinal numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 ospeed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 other permissions . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 output block size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 output file name prefix. . . . . . . . . . . . . . . . . . . . 24, 25 output file name suffix . . . . . . . . . . . . . . . . . . . . . . . 25 output format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 output format, portable . . . . . . . . . . . . . . . . . . . . . . 92 output of entire files . . . . . . . . . . . . . . . . . . . . . . . . . . 7 output of parts of files . . . . . . . . . . . . . . . . . . . . . . . 21 output settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 output tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 overwriting of input, allowed. . . . . . . . . . . . . . . . . . 32 owned by effective gid check . . . . . . . . . . . . . . . . . 100 owned by effective uid check . . . . . . . . . . . . . . . . . 100 owner of file, permissions for. . . . . . . . . . . . . . . . . 136 owner, default . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 ownership of installed files, setting . . . . . . . . . . . . 74

P

p for FIFO file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 pad character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 pad instead of timing for delaying . . . . . . . . . . . . 109 padding of numeric fields . . . . . . . . . . . . . . . . . . . . 119 paragraphs, reformatting . . . . . . . . . . . . . . . . . . . . . 14 parenb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 parent directories and cp . . . . . . . . . . . . . . . . . . . . . 70 parent directories, creating . . . . . . . . . . . . . . . . . . . 83 parent directories, creating missing . . . . . . . . . . . . 74 parent directories, removing . . . . . . . . . . . . . . . . . . 85 parentheses for grouping . . . . . . . . . . . . . . . . . . . . 102 parity. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 parity errors, marking . . . . . . . . . . . . . . . . . . . . . . . 108 parity, ignoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 parmrk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 parodd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 parsing date strings . . . . . . . . . . . . . . . . . . . . . . . . . 120 parts of files, output of . . . . . . . . . . . . . . . . . . . . . . . 21 pass8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 passwd entry, and su shell . . . . . . . . . . . . . . . . . . . 127 paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 Paterson, R. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 PATH . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125, 128 pathchk. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 pattern matching . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 PC filesystem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 pcfs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 pebibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 permission tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 permissions of installed files, setting . . . . . . . . . . . 74 permissions, changing access . . . . . . . . . . . . . . . . . . 88 permissions, copying existing . . . . . . . . . . . . . . . . 137 permissions, for changing file timestamps . . . . . . 89 permissions, output by ls . . . . . . . . . . . . . . . . . . . . 59 petabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 phone directory order . . . . . . . . . . . . . . . . . . . . . . . . 31 pieces, splitting a file into . . . . . . . . . . . . . . . . . . . . 24 Pinard, F. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1, 145 pipe fitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 Plass, Michael F. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 platform, hardware . . . . . . . . . . . . . . . . . . . . . . . . . 123 pm in date strings . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 portable file names, checking for . . . . . . . . . . . . . 106 portable output format. . . . . . . . . . . . . . . . . . . . . . . 92 posix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 posix output format . . . . . . . . . . . . . . . . . . . . . . . . . 92 POSIXLY_CORRECT . . . . . . . . . . . . . 2, 6, 16, 27, 32, 54 POSIXLY_CORRECT, and block size . . . . . . . . . . . . . . . 3 pr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 prime factors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 print name of current directory . . . . . . . . . . . . . . 107 print system information . . . . . . . . . . . . . . . . . . . . 122 print terminal file name . . . . . . . . . . . . . . . . . . . . . 113 printenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 printf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 printing all or some environment variables …. 113 printing color database . . . . . . . . . . . . . . . . . . . . . . . 67 printing current user information . . . . . . . . . . . . 115 printing current usernames . . . . . . . . . . . . . . . . . . 115 printing groups a user is in . . . . . . . . . . . . . . . . . . 115 printing real and effective uid and gid . . . . . . . . 114 printing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 printing text, commands for . . . . . . . . . . . . . . . . . . 97 printing the current time . . . . . . . . . . . . . . . . . . . . 117 printing the effective UID . . . . . . . . . . . . . . . . . . . 114 printing the host identifier. . . . . . . . . . . . . . . . . . . 123 printing the hostname. . . . . . . . . . . . . . . . . . . . . . . 123 printing user’s login name . . . . . . . . . . . . . . . . . . . 114 printing, preparing files for . . . . . . . . . . . . . . . . . . . 15 priority, modifying . . . . . . . . . . . . . . . . . . . . . . . . . . 126 processes, commands for controlling . . . . . . . . . . 129 prompting, and ln . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 prompting, and mv . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 prompting, and rm . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 prompts, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 prompts, omitting . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 prterase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 ptx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 punct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 pure numbers in date strings . . . . . . . . . . . . . . . . 145 pwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107

Q

quit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 quoting style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

R

radix for file offsets . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 raw . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 read errors, ignoring . . . . . . . . . . . . . . . . . . . . . . . . . 73 read from stdin and write to stdout and files . . 105 read permission. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 read permission, symbolic . . . . . . . . . . . . . . . . . . . 136 read system call, and holes . . . . . . . . . . . . . . . . . . . 71 readable file check . . . . . . . . . . . . . . . . . . . . . . . . . . 100 real uid and gid, printing . . . . . . . . . . . . . . . . . . . . 114 recursive directory listing . . . . . . . . . . . . . . . . . . . . . 57 recursively changing access permissions . . . . . . . . 89 recursively changing file ownership . . . . . . . . . . . . 87 recursively changing group ownership. . . . . . . . . . 88 recursively copying directories . . . . . . . . . . . . . 69, 70 redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 reformatting paragraph text . . . . . . . . . . . . . . . . . . 14 regular expression matching . . . . . . . . . . . . . . . . . 102 regular file check. . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 relations, numeric or string . . . . . . . . . . . . . . . . . . 103 relative items in date strings. . . . . . . . . . . . . . . . . 144 release of kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 remainder . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 remote hostname . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 removing empty directories . . . . . . . . . . . . . . . . . . . 85 removing files after shredding . . . . . . . . . . . . . . . . . 79 removing files or directories. . . . . . . . . . . . . . . . . . . 76 removing files or directories (via the unlink syscall) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 removing permissions . . . . . . . . . . . . . . . . . . . . . . . 136 repeated characters . . . . . . . . . . . . . . . . . . . . . . . . . . 51 repeated output of a string . . . . . . . . . . . . . . . . . . . 98 restricted deletion flag . . . . . . . . . . . . . . . . . . . . . . 135 restricted shell. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 return, ignoring. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 return, translating to newline . . . . . . . . . . . 108, 109 reverse sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . 32, 61 reversing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 rmdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 rn format for nl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 root as default owner . . . . . . . . . . . . . . . . . . . . . . . . 74 root directory, running a program in a specified

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 root, becoming . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 rows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 rprnt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 RTS/CTS flow control . . . . . . . . . . . . . . . . . . . . . . 108 running a program in a modified environment

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 running a program in a specified root directory

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 rz format for nl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

S

Salz, Rich . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145 same file check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 sane . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 scheduling priority, modifying. . . . . . . . . . . . . . . . 126 screen columns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 seconds since the epoch . . . . . . . . . . . . . . . . . . . . . 117 section delimiters of pages . . . . . . . . . . . . . . . . . . . . . 9 seek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 self-backups . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 send a signal to processes. . . . . . . . . . . . . . . . . . . . 129 sentences and line-breaking . . . . . . . . . . . . . . . . . . . 14 separator for numbers in seq . . . . . . . . . . . . . . . . 132 seq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 sequence of numbers . . . . . . . . . . . . . . . . . . . . . . . . 132 set-group-id check . . . . . . . . . . . . . . . . . . . . . . . . . . 100 set-user-id check . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 setgid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 setting permissions. . . . . . . . . . . . . . . . . . . . . . . . . . 136 setting the hostname . . . . . . . . . . . . . . . . . . . . . . . . 123 setting the time . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 setuid. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 setup for color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 sh syntax for color setup . . . . . . . . . . . . . . . . . . . . . 67 SHELL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 SHELL environment variable, and color . . . . . . . . . 67 shell utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 shred . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 SI output . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 60, 91, 94 simple backup method . . . . . . . . . . . . . . . . . . . . . . . . 3 SIMPLE_BACKUP_SUFFIX. . . . . . . . . . . . . . . . . . . . . . . . 3 single-column output of files. . . . . . . . . . . . . . . . . 62 size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 size for main memory sorting . . . . . . . . . . . . . . . . . 32 size of file to shred . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 size of files, reporting . . . . . . . . . . . . . . . . . . . . . . . . 60 size of files, sorting files by . . . . . . . . . . . . . . . . 61 skip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 sleep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 socket check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 software flow control . . . . . . . . . . . . . . . . . . . . . . . . 108 sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 sort field . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 sort zero-terminated lines. . . . . . . . . . . . . . . . . . . . . 33 sorted files, operations on . . . . . . . . . . . . . . . . . . . . 30 sorting files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 sorting ls output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 sparse files, copying . . . . . . . . . . . . . . . . . . . . . . . . . . 71 special characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 special file types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 special files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 special settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 specifying sets of characters . . . . . . . . . . . . . . . . . . 50 speed . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 splitting a file into pieces . . . . . . . . . . . . . . . . . . . . . 24 splitting a file into pieces by context. . . . . . . . . . . 25 squeezing blank lines . . . . . . . . . . . . . . . . . . . . . . . . . . 7 squeezing repeat characters . . . . . . . . . . . . . . . . . . . 52 Stallman, R.. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 standard input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 standard output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 start . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 status time, printing or sorting by . . . . . . . . . . . 61 sticky . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 sticky bit check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 stop bits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 strftime and date . . . . . . . . . . . . . . . . . . . . . . . . . 117 string constants, outputting . . . . . . . . . . . . . . . . . . 11 string expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 string tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 strip directory and suffix from file names . . . . . 106 stripping non-directory suffix . . . . . . . . . . . . . . . . 106 stripping symbol table information . . . . . . . . . . . . 74 stripping trailing slashes . . . . . . . . . . . . . . . . . . 71, 76 stty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 su . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 substitute user and group ids . . . . . . . . . . . . . . . . 127 substr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 subtracting permissions . . . . . . . . . . . . . . . . . . . . . 136 subtraction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 successful exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 suffix, stripping from file names . . . . . . . . . . . . . . 106 sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 summarizing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 super-user, becoming. . . . . . . . . . . . . . . . . . . . . . . . 127 superblock, writing . . . . . . . . . . . . . . . . . . . . . . . . . . 96 supplementary groups, printing . . . . . . . . . . . . . . 115 susp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 swab (byte-swapping) . . . . . . . . . . . . . . . . . . . . . . . . 73 swap space, saving text image in . . . . . . . . . . . . . 135 swtch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 symbol table information, stripping . . . . . . . . . . . 74 symbolic (soft) links, creating . . . . . . . . . . . . . . . . . 81 symbolic link check . . . . . . . . . . . . . . . . . . . . . . . . . 100 symbolic link, defined . . . . . . . . . . . . . . . . . . . . . . . . 81 symbolic links and pwd . . . . . . . . . . . . . . . . . . . . . . 107 symbolic links, changing group. . . . . . . . . . . . . . . . 88 symbolic links, changing owner. . . . . . . . . 86, 87, 88 symbolic links, copying . . . . . . . . . . . . . . . . . . . . . . . 69 symbolic links, copying with . . . . . . . . . . . . . . . . . . 71 symbolic links, dereferencing. . . . . . . . . . . . . . . . . . 57 symbolic links, dereferencing in du . . . . . . . . . . . . 94 symbolic links, dereferencing in stat . . . . . . . . . . 95 symbolic links, permissions of . . . . . . . . . . . . . . . . . 88 symbolic modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 sync (padding with nulls) . . . . . . . . . . . . . . . . . . . . 73 synchronize disk and memory . . . . . . . . . . . . . . . . . 96 syslog. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 system context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 system information, printing . . . . . . . . . . . . . . . . . 122 system name, printing. . . . . . . . . . . . . . . . . . . . . . . 123 System V sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

T

tabn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 tabs to spaces, converting . . . . . . . . . . . . . . . . . . . . 54 tabstops, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 tac . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 tagged paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 tail . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 tandem. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 target directory . . . . . . . . . . . . . . . . . 5, 71, 75, 76, 82 tebibyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 tee . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 telephone directory order . . . . . . . . . . . . . . . . . . . . . 31 temporary directory. . . . . . . . . . . . . . . . . . . . . . . . . . 33 terabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . . 4 TERM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 terminal check. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 terminal file name, printing. . . . . . . . . . . . . . . . . . 113 terminal lines, currently used . . . . . . . . . . . . . . . . 115 terminal settings. . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 terminal, using color iff. . . . . . . . . . . . . . . . . . . . . . . 62 terse output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 text image, saving in swap space . . . . . . . . . . . . . 135 text input files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 text utilities . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 text, displaying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 text, reformatting. . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 this in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 145 time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90, 112 time directives. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 time formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16, 117 time of day item . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 time setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 time style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 time units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 time zone correction . . . . . . . . . . . . . . . . . . . . . . . . 143 time zone item . . . . . . . . . . . . . . . . . . . . . . . . . 142, 144 time, printing or setting . . . . . . . . . . . . . . . . . . . . . 117 TIME_STYLE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 timestamps of installed files, preserving . . . . . . . . 74 timestamps, changing file . . . . . . . . . . . . . . . . . . . . . 89 TMPDIR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30, 33 today in date strings . . . . . . . . . . . . . . . . . . . . . . . . 145 tomorrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 tomorrow in date strings . . . . . . . . . . . . . . . . . . . . 145 topological sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 tostop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 total counts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 touch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 tr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 trailing slashes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 translating characters . . . . . . . . . . . . . . . . . . . . . . . . 52 true . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 truncating output file, avoiding . . . . . . . . . . . . . . . 73 tsort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 tty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 Twenex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 two-way parity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 type size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

U

u, and disabling special characters . . . . . . . . . . . 111 ucase, converting to . . . . . . . . . . . . . . . . . . . . . . . . . 73 ufs filesystem type . . . . . . . . . . . . . . . . . . . . . . . . . . 93 umask and modes. . . . . . . . . . . . . . . . . . . . . . . . . . . 139 uname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 unblock. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 unexpand. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 uniq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 unique lines, outputting . . . . . . . . . . . . . . . . . . . . . . 36 uniquify files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 uniquifying output . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77, 85 unprintable characters, ignoring. . . . . . . . . . . . . . . 31 unsorted directory listing . . . . . . . . . . . . . . . . . . . . . 61 upper . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 uppercase, translating to lowercase . . . . . . . . . . . 109 use time, changing. . . . . . . . . . . . . . . . . . . . . . . . . . . 90 use time, printing or sorting files by . . . . . . . . . . 61 USER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 user id, switching . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 user information, commands for. . . . . . . . . . . . . . 114 user name, printing . . . . . . . . . . . . . . . . . . . . . . . . . 114 usernames, printing current. . . . . . . . . . . . . . . . . . 115 users . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 UTC . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 utmp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114

V

valid file names, checking for . . . . . . . . . . . . . . . . 106 vdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 verbose ls format . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 verifying MD5 checksums. . . . . . . . . . . . . . . . . . . . . 29 version number, finding . . . . . . . . . . . . . . . . . . . . . . . 2 version of kernel . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 version, sorting option for ls . . . . . . . . . . . . . . . . 61 version-control Emacs variable . . . . . . . . . . . . . . 2 VERSION_CONTROL . . . . . . . . . . . . . . . 2, 68, 74, 76, 82 vertical sorted files in columns . . . . . . . . . . . . . . 62 vtn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109

W

wc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 week in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 144 werase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 wheel group, not supported. . . . . . . . . . . . . . . . . . 128 who . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 who am i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 whoami. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 word count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 working context . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 working directory, printing . . . . . . . . . . . . . . . . . . 107 wrapping long input lines. . . . . . . . . . . . . . . . . . . . . 20 writable file check . . . . . . . . . . . . . . . . . . . . . . . . . . 100 write permission . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 write permission, symbolic. . . . . . . . . . . . . . . . . . . 136 write, allowed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116

X

xcase . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 xdigit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 XON/XOFF flow control . . . . . . . . . . . . . . . . . . . . 108

Y

year in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 144 yes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 yesterday . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 yesterday in date strings . . . . . . . . . . . . . . . . . . . 145 yottabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . 5 Youmans, B. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Z

zero-length string check . . . . . . . . . . . . . . . . . . . . . 101 zettabyte, definition of . . . . . . . . . . . . . . . . . . . . . . . . 4

Short Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 Output of entire files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4 Formatting file contents . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5 Output of parts of files. . . . . . . . . . . . . . . . . . . . . . . . . . . 21 6 Summarizing files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 7 Operating on sorted files . . . . . . . . . . . . . . . . . . . . . . . . . 30 8 Operating on fields within a line . . . . . . . . . . . . . . . . . . . . 46 9 Operating on characters . . . . . . . . . . . . . . . . . . . . . . . . . . 50 10 Directory listing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 11 Basic operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 12 Special file types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 13 Changing file attributes . . . . . . . . . . . . . . . . . . . . . . . . . . 86 14 Disk usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 15 Printing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 16 Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 17 Redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 18 File name manipulation . . . . . . . . . . . . . . . . . . . . . . . . . 106 19 Working context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 20 User information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 21 System context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 22 Modified command invocation . . . . . . . . . . . . . . . . . . . . . 125 23 Process control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 24 Delaying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 25 Numeric operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 26 File permissions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 27 Date input formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 28 Opening the Software Toolbox . . . . . . . . . . . . . . . . . . . . . 146 A GNU Free Documentation License . . . . . . . . . . . . . . . . . . 154 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Common options . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 Backup options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Block size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Target directory. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.4 Trailing slashes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.5 Standards conformance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

3 Output of entire files . . . . . . . . . . . . . . . . . . . . . . . 7

3.1 cat: Concatenate and write files. . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 tac: Concatenate and write files in reverse . . . . . . . . . . . . . . . . 8 3.3 nl: Number lines and write files . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.4 od: Write files in octal or other formats . . . . . . . . . . . . . . . . . . 10

4 Formatting file contents . . . . . . . . . . . . . . . . . . . 14

4.1 fmt: Reformat paragraph text. . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 pr: Paginate or columnate files for printing. . . . . . . . . . . . . . . 15 4.3 fold: Wrap input lines to fit in specified width . . . . . . . . . . . 20

5 Output of parts of files . . . . . . . . . . . . . . . . . . . . 21

5.1 head: Output the first part of files. . . . . . . . . . . . . . . . . . . . . . . 21 5.2 tail: Output the last part of files . . . . . . . . . . . . . . . . . . . . . . . 21 5.3 split: Split a file into fixed-size pieces. . . . . . . . . . . . . . . . . . . 24 5.4 csplit: Split a file into context-determined pieces . . . . . . . . 25

6 Summarizing files . . . . . . . . . . . . . . . . . . . . . . . . . 27

6.1 wc: Print byte, word, and line counts . . . . . . . . . . . . . . . . . . . . 27 6.2 sum: Print checksum and block counts . . . . . . . . . . . . . . . . . . . 28 6.3 cksum: Print CRC checksum and byte counts. . . . . . . . . . . . . 28 6.4 md5sum: Print or check message-digests . . . . . . . . . . . . . . . . . . 28

7 Operating on sorted files . . . . . . . . . . . . . . . . . . 30

7.1 sort: Sort text files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 7.2 uniq: Uniquify files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 7.3 comm: Compare two sorted files line by line . . . . . . . . . . . . . . . 36 7.4 tsort: Topological sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 7.5 tsort: Background. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 7.6 ptx: Produce permuted indexes . . . . . . . . . . . . . . . . . . . . . . . . . 39

7.6.1 General options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.6.2 Charset selection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.6.3 Word selection and input processing . . . . . . . . . . . . 40 7.6.4 Output formatting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 7.6.5 The gnu extensions to ptx . . . . . . . . . . . . . . . . . . . . . 45

8 Operating on fields within a line . . . . . . . . . . . 46

8.1 cut: Print selected parts of lines . . . . . . . . . . . . . . . . . . . . . . . . 46 8.2 paste: Merge lines of files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 8.3 join: Join lines on a common field . . . . . . . . . . . . . . . . . . . . . . 47

9 Operating on characters. . . . . . . . . . . . . . . . . . . 50

9.1 tr: Translate, squeeze, and/or delete characters . . . . . . . . . . 50 9.1.1 Specifying sets of characters . . . . . . . . . . . . . . . . . . . . 50 9.1.2 Translating. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 9.1.3 Squeezing repeats and deleting. . . . . . . . . . . . . . . . . . 52 9.1.4 Warning messages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

9.2 expand: Convert tabs to spaces . . . . . . . . . . . . . . . . . . . . . . . . . 54 9.3 unexpand: Convert spaces to tabs . . . . . . . . . . . . . . . . . . . . . . . 55

10 Directory listing . . . . . . . . . . . . . . . . . . . . . . . . . 56

10.1 ls: List directory contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 10.1.1 Which files are listed. . . . . . . . . . . . . . . . . . . . . . . . . . 56 10.1.2 What information is listed. . . . . . . . . . . . . . . . . . . . . 57 10.1.3 Sorting the output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 10.1.4 More details about version sort . . . . . . . . . . . . . . . . 62 10.1.5 General output formatting . . . . . . . . . . . . . . . . . . . . 62 10.1.6 Formatting file timestamps . . . . . . . . . . . . . . . . . . . . 64 10.1.7 Formatting the file names . . . . . . . . . . . . . . . . . . . . . 65

10.2 dir: Briefly list directory contents . . . . . . . . . . . . . . . . . . . . . . 67 10.3 vdir: Verbosely list directory contents . . . . . . . . . . . . . . . . . . 67 10.4 dircolors: Color setup for ls . . . . . . . . . . . . . . . . . . . . . . . . . 67

11 Basic operations . . . . . . . . . . . . . . . . . . . . . . . . . 68

11.1 cp: Copy files and directories . . . . . . . . . . . . . . . . . . . . . . . . . . 68 11.2 dd: Convert and copy a file . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 11.3 install: Copy files and set attributes . . . . . . . . . . . . . . . . . . 73 11.4 mv: Move (rename) files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 11.5 rm: Remove files or directories. . . . . . . . . . . . . . . . . . . . . . . . . . 76 11.6 shred: Remove files more securely. . . . . . . . . . . . . . . . . . . . . . 77

12 Special file types. . . . . . . . . . . . . . . . . . . . . . . . . 81

12.1 link: Make a hard link via the link syscall . . . . . . . . . . . . . . 81 12.2 ln: Make links between files . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 12.3 mkdir: Make directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 12.4 mkfifo: Make FIFOs (named pipes) . . . . . . . . . . . . . . . . . . . . 84 12.5 mknod: Make block or character special files . . . . . . . . . . . . . 84 12.6 rmdir: Remove empty directories. . . . . . . . . . . . . . . . . . . . . . . 85 12.7 unlink: Remove files via the unlink syscall. . . . . . . . . . . . . . 85

13 Changing file attributes . . . . . . . . . . . . . . . . . . 86

13.1 chown: Change file owner and group . . . . . . . . . . . . . . . . . . . . 86 13.2 chgrp: Change group ownership. . . . . . . . . . . . . . . . . . . . . . . . 87 13.3 chmod: Change access permissions . . . . . . . . . . . . . . . . . . . . . . 88 13.4 touch: Change file timestamps . . . . . . . . . . . . . . . . . . . . . . . . . 89

14 Disk usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91

14.1 df: Report filesystem disk space usage . . . . . . . . . . . . . . . . . . 91 14.2 du: Estimate file space usage . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 14.3 stat: Report file or filesystem status . . . . . . . . . . . . . . . . . . . 95 14.4 sync: Synchronize data on disk with memory . . . . . . . . . . . 96

15 Printing text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97

15.1 echo: Print a line of text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 15.2 printf: Format and print data. . . . . . . . . . . . . . . . . . . . . . . . . 97 15.3 yes: Print a string until interrupted . . . . . . . . . . . . . . . . . . . . 98

16 Conditions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

16.1 false: Do nothing, unsuccessfully . . . . . . . . . . . . . . . . . . . . . . 99 16.2 true: Do nothing, successfully . . . . . . . . . . . . . . . . . . . . . . . . . 99 16.3 test: Check file types and compare values . . . . . . . . . . . . . . 99

16.3.1 File type tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 16.3.2 Access permission tests . . . . . . . . . . . . . . . . . . . . . . 100 16.3.3 File characteristic tests. . . . . . . . . . . . . . . . . . . . . . . 100 16.3.4 String tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 16.3.5 Numeric tests . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 16.3.6 Connectives for test. . . . . . . . . . . . . . . . . . . . . . . . . 101

16.4 expr: Evaluate expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 16.4.1 String expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 16.4.2 Numeric expressions . . . . . . . . . . . . . . . . . . . . . . . . . 103 16.4.3 Relations for expr . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 16.4.4 Examples of using expr . . . . . . . . . . . . . . . . . . . . . . 104

17 Redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105

17.1 tee: Redirect output to multiple files . . . . . . . . . . . . . . . . . . 105

18 File name manipulation . . . . . . . . . . . . . . . . . 106

18.1 basename: Strip directory and suffix from a file name … 106 18.2 dirname: Strip non-directory suffix from a file name …. 106 18.3 pathchk: Check file name portability . . . . . . . . . . . . . . . . . . 106

19 Working context. . . . . . . . . . . . . . . . . . . . . . . . 107

19.1 pwd: Print working directory . . . . . . . . . . . . . . . . . . . . . . . . . . 107 19.2 stty: Print or change terminal characteristics . . . . . . . . . . 107

19.2.1 Control settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 19.2.2 Input settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 19.2.3 Output settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 19.2.4 Local settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 19.2.5 Combination settings . . . . . . . . . . . . . . . . . . . . . . . . 110 19.2.6 Special characters . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 19.2.7 Special settings. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

19.3 printenv: Print all or some environment variables. . . . . . 113 19.4 tty: Print file name of terminal on standard input. . . . . . 113

20 User information . . . . . . . . . . . . . . . . . . . . . . . 114

20.1 id: Print real and effective uid and gid . . . . . . . . . . . . . . . . 114 20.2 logname: Print current login name . . . . . . . . . . . . . . . . . . . . 114 20.3 whoami: Print effective user id . . . . . . . . . . . . . . . . . . . . . . . . 114 20.4 groups: Print group names a user is in . . . . . . . . . . . . . . . . 115 20.5 users: Print login names of users currently logged in … 115 20.6 who: Print who is currently logged in . . . . . . . . . . . . . . . . . . 115

21 System context . . . . . . . . . . . . . . . . . . . . . . . . . 117

21.1 date: Print or set system date and time . . . . . . . . . . . . . . . 117 21.1.1 Time directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 21.1.2 Date directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 21.1.3 Literal directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 21.1.4 Padding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 21.1.5 Setting the time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 21.1.6 Options for date . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 21.1.7 Examples of date . . . . . . . . . . . . . . . . . . . . . . . . . . . 121

21.2 uname: Print system information . . . . . . . . . . . . . . . . . . . . . . 122 21.3 hostname: Print or set system name . . . . . . . . . . . . . . . . . . . 123 21.4 hostid: Print numeric host identifier. . . . . . . . . . . . . . . . . . 123

22 Modified command invocation . . . . . . . . . . . 125

22.1 chroot: Run a command with a different root directory

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 125 22.2 env: Run a command in a modified environment . . . . . . . 125 22.3 nice: Run a command with modified scheduling priority

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 22.4 nohup: Run a command immune to hangups . . . . . . . . . . . 126 22.5 su: Run a command with substitute user and group id . . 127

23 Process control . . . . . . . . . . . . . . . . . . . . . . . . . 129

23.1 kill: Send a signal to processes. . . . . . . . . . . . . . . . . . . . . . . 129

24 Delaying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

24.1 sleep: Delay for a specified time . . . . . . . . . . . . . . . . . . . . . . 131

25 Numeric operations. . . . . . . . . . . . . . . . . . . . . 132

25.1 factor: Print prime factors. . . . . . . . . . . . . . . . . . . . . . . . . . . 132 25.2 seq: Print numeric sequences . . . . . . . . . . . . . . . . . . . . . . . . . 132

26 File permissions . . . . . . . . . . . . . . . . . . . . . . . . 135

26.1 Structure of File Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . 135 26.2 Symbolic Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

26.2.1 Setting Permissions . . . . . . . . . . . . . . . . . . . . . . . . . . 136 26.2.2 Copying Existing Permissions . . . . . . . . . . . . . . . . 137 26.2.3 Changing Special Permissions . . . . . . . . . . . . . . . . 137 26.2.4 Conditional Executability . . . . . . . . . . . . . . . . . . . . 138 26.2.5 Making Multiple Changes . . . . . . . . . . . . . . . . . . . . 138 26.2.6 The Umask and Protection . . . . . . . . . . . . . . . . . . . 139

26.3 Numeric Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139

27 Date input formats . . . . . . . . . . . . . . . . . . . . . 141

27.1 General date syntax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 27.2 Calendar date items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 27.3 Time of day items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 27.4 Time zone items . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 27.5 Day of week items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 27.6 Relative items in date strings . . . . . . . . . . . . . . . . . . . . . . . . . 144 27.7 Pure numbers in date strings. . . . . . . . . . . . . . . . . . . . . . . . . . 145 27.8 Authors of getdate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 145

28 Opening the Software Toolbox. . . . . . . . . . . 146

Toolbox Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 I/O Redirection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 146 The who Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 The cut Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 The sort Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 The uniq Command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 Putting the Tools Together . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148

Appendix A GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154

ADDENDUM: How to use this License for your documents. … 159

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160